A Complete Bibliography of Publications in SIAM Journal on Mathematical Analysis for 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 May 2024 Version 2.03

Title word cross-reference

(N + 1) [BCV06]. 1 [Bos03, FH01, LY05]. 1/|x| [RW00b]. 1 ≤ p < ∞ [BFP08]. 2 [CG00, HL02a, HL02b, Kim09, Kim13]. 2 × 2 [CHS08, HL02a, HL02b]. 3 [CL04, CF01, Mas01, MPP09, ST01]. _{3}F_{2} [BPR00b]. BV [DP09]. C^{1} [KK04a]. C^{∞} [HS08]. d [LL00]. D^{2} [AHM09]. δ [CL03]. div (\frac{Du|}{D}u|) = u [BCN05]. Γ [ABV06, BN08, Hor08, CMS04, GM05, Ped04]. H^{s}(∂Ω) [Auc06]. H_{2}^{1} [Mik09]. ∞ [CDJ08]. k [KS03b]. κ → +∞ [BCM00]. L^{1} [Ha01, HY03]. L^{2} [HZ09, HK06, VZ07, BCN02]. L^{∞} [AG02, Bia01, BN08]. L^{p} [CCD02, IK01]. L_{p} [BP07]. L_{q}((0, T), L_{p}) [Kry01]. R^{3} [BV08, Han01]. R^{n} [NW04, DKS02]. Z [CT09]. N [Alì03, BCV06, KT01]. P [CQW03, DP09, AMRT09, BFP08, CR03, CG08, IL05, Lie05, LL03]. q [AL06]. R^{1} [WW02]. R^{2} [BCN05]. S^{2} [AHM09]. T [MP08, Kri04]. u_{xxxx} + u_{yyyy} = f [KS06a]. ε [Sug09].

-Body [CL04, BCV06]. -Convergence [ABV06, BN08, Hor08, CMS04, Ped04]. -Critical [HK06, VZ07]. -Dimensional [HL02a, HL02b, Alì03, BCV06]. -Harmonic [BFP08]. -Junctions [MP08]. -Laplace [IL05]. -Laplacian [AMRT09, CQW03]. -Limit [GM05]. -Manifolds [CG00]. -Monotone [KS03b]. -Racah [AL06]. -Regularity [Sug09]. -Shocks [CL03]. -Solutions [Mik09]. -Spaces [Kry01]. -System [CG08, DP09]. -Theory [BP07].

-Wasserstein [CDJ08]. -Wave [LL00, LL03]. -Waves [KT01].

1D [Bos05].

2D [Wu05].

34 [Pla09b].

Absence [ET09]. Absolutely [FLT09]. Absorption [CQW03]. Accelerated [GK05]. Acceleration [MO09, SS09a]. Acoustic [HH01, HLTT08, Mei08]. Acoustical [BV08]. Acoustics [KP08]. Across [O’D08]. Activity [PSF05]. Adams [DE06]. Adaptive [KS09c]. Additive [DS05]. Adhesion [NT08]. Adiabatic [AMP05, MP01, NN01]. Adjoint [EL02]. Adjoints [FHW02]. Admissibility [Hat00]. Admissible [FN05]. Affine [Joh03]. after [FMP05]. Age [Wal09a]. Age- [Wal09a]. Agents [BFG04]. Aid [PS00]. Algebraic [BW08]. Algebras [FS09]. Algorithm [CN06, Ish05]. Algorithms [Han08]. Allen [Tan07]. Allocation [DZ08]. Alloys [MPP09, Ste05, Vis09]. Almost [DGG00, HS02]. Along [AH07]. Alternate [BGSZ01]. American [BX09, CC07]. Ampère [JW07]. Amplitude [AIM03, HK00b, LLYZ03, MZ05b, Sun01, YZ09]. Anal [HKS07, Pla09b]. Analysis [ABV06, BK05, BX09, BDV03, BV00a, BNG04, BMR09, CCG05, CFTV03, CJ04, CC07, CCCS06, CE07, DGG08, DLY07, FFH02, Fou06, FMS03, GIK05, KO00, KS09c, Mac04, MB08, Pel01, PRS08, San05, Ste05, Wan08, Zui00, BC06a]. Analyticity [HN05]. Angiogenesis [FFH02]. Angle [II01]. Angular [Cla00]. Anharmonic [Joc00]. Anisotropic [AMSW05, BK04, CN06, FMP09, KN00, KS06a, KP07, Kyr04, MP03, HL02b]. Anisotropy [AG01a]. Annealing [CHK08]. Annulus [Lou06]. Any [Lie05].

Application [Bra05, Bri03, BST05, BLM09, BFK03, CCM07, CR06, CF03, Dal07, DLY07, HM09b, KK08b, LW09, ZD00]. Applications [ACFS09, BL03, BL01, BN08, Car03, COV02, FLT09, HH01, IV04, KK08a, MVNO08, NW03, NS04, SS09b, Tra06]. Approach [Bos00, BV02, DS06, HM09a, JN04, KY04, Len08, MO09, NS09, Ryb01]. Approximate [FC06, Lay07]. Approximation [AMP05, AFM01, ABDG04, AG01b, BN02, BBP05, BGL07, CGZ09, COV02, DP03, FS01, GP04, HS08, II01, IM04, Kei01, KS03b, MPP09, Rou03, TTX03, Wil09, Wri05, Yoo01]. Approximations [BPR00b, BFP08, BC06b, CP00, CL09, KN09, TT00]. Arbitrary [GW07, Lie05]. Arc [BPR00a, BPR00b]. Area [BS05, Yua06]. Argument [LWS05]. Arise [DZ08]. Arising [BV00a, Bla06, CCG05, CF00, DLS05, El 07, KKSY02, LS00, NK08, Osh04, YYW08, Yoc07]. Ascending [SW08]. Aspects [FG09]. Associated [CR03, CHK08, Ped06]. Assumption [CLR02, Pla09b]. Asymptotic [ABDG04, AK03, BF06, Bou08, BLM09, CGL04, CLGZ06, CEGMM09, CT09, Cui08, Dal09, DN09, DG00, DD02b, DKP07, Dun01, FM01, Joc05, KPS09, KY04, Lóp00, Mar06b, Mos05, Nak03, Nis02, NN01, PLN02, QW00, SS03b, SW08, SZ00, Sug09, Tsa07, Wan08, WC09]. Asymptotically [LRRBS09, Nak01]. Asymptotically-Free [Nak01]. Asymptotics [Bjo08, CFG06, GPV00, JN04, Joc00, Kag01, KS09a, MNR05, MV07, Miz01, PS00, dMKST09, vBPW08, EGW06]. Attainable [AG02]. Attraction [Ria04]. Attractors [HZ03, MZ05a, MPS05, MPS06, Rob00]. Aubry [Gom03]. Autocatalysis [CQ07]. Avascular [CE07]. Average [Sun07]. Averaged [HL06]. Averages [BG08]. Averaging [KY04, Wes02]. Away [LLLY04]. Axially [Kan04, SZ07].

Axisymmetric [LW09, SZ05].

B [AH07, FG09]. Background [KS05a]. Backscatter [HHR09, HHR11]. Balance [FL00, GMS09, MP03]. Balian [BCP06]. Banach [FHW02]. Band [NW04]. Band-Limited [NW04]. Bandlimited [KP05]. Based [DY00, DLY07, FKK09, FMS03, HKP07, HSZ03a, HSZ03b, MCP08]. Bases [AS00, KS05b]. Basic [FG09, Grü03]. Basis [BCR01, LYY07, NW02, NW04, Yoo01]. BC [Bel03]. BC-Method [Bel03]. Be [FG09]. Beams [FMP09]. Bearing [WNW07]. Bearings [CJT09]. Before [BV05]. Behavior [BS00a, BS01, BK00, BDK09, CQW03, CEGMM09, CA00, Dal09, DS06, DG00, DD02b, DS05, Gev08, Han01, Hat03, Joc05, LRR06, LP08, LXY00, MV05, Nar08, NN01, PLN02, SS03b, Tak00, Wan00, WZF08]. Behaviors [Sug09]. below [VZ07]. Bence [Ish05]. Bending [Wan08]. Benjamin [MST01]. Besov [Wu05]. Bessel [Dun01]. Between [FM09, Miz03]. BGK [CS06]. Bi [Li05]. Bi-isotropic [Li05]. Bifurcation [CE07, DGG08, Du00, FL00, FH07, Hur06, KKSW08, KA00, Rob00, SS08]. Bifurcations [BPBW05, BW08, KS02a]. Bilaplacian [DDGM07]. Billiard [Pla09a]. Binary [Vis09]. Bioelectric [PSF05]. Biofluid [CCS07]. Biorthogonal [RTW01]. Birefringent [LP00]. Birkhoff [BB05]. Birth [FM04]. Bistable [BCC03, DIN04, Tsa07]. Black [Cré03]. Blatz [CP07b]. Bloch [COV02, MH01]. Blood [MOPMW06]. Blow [CCM07, EGW06, FMP05, Kim06, LP08, Mas01, MV05, WY08, YCL01]. Blow-Up [Kim06, LP08, WY08, YCL01, CCM07, EGW06, FMP05, Mas01, MV05]. Blowup [Esp05, HK06, Jen00, KLVZ09, VZ07, YZ05]. Body [AMR02, BFV08, BCV06, CL04, PP04].

Bogdanov [Sti00]. Boltzmann

[Li11, FKW06, Li09, MT07a, Nis02]. Bolza [CP07b]. Bose [AB06]. Bottle [KRS09]. Bound [DW05, KPS09, LM08, LW08]. Boundaries [BMR09, MV07]. Boundary [Abd03, AG01a, AMR02, AD05, AH07, ABDG04, ANR09, AG02, AMRT09, BX09, BIY08, BFV08, BFG04, BV00a, BGM08, BL01, BW04, Bos03, Bos05, BV02, BFK03, CKK06, CK01a, Cas03, CF00, CF03, CC07, CE07, Cui08, Ebm00, EGG03, FH01, FP02, FS05, FH07, GIK05, GIK08, Gev08, Gor02, GJT06, Gui07, HR08, HP09, HMS03, HMS04, IV04, II01, JZ09a, KY03a, Kel06, Kry03, LLLY04, LS00, LP08, LCS03, MS04, MV05, NS07, NK08, PP04, PLN02, PP00b, PRS08, RW08, SR09, SS09b, WC09, YZ09]. Bounded [Bou00, GM01, Kel06, Mik09, ZS02]. Bounds [CWM08, DP05, Lay07, Lie05, MM05, ORS06, YZ09]. Bourgain’s [MR02]. Bourgain’s-Type [MR02]. Bramble [DL04]. Branches [Mar06a]. Breaking [GZ09, KKSW08, Tov08]. Bridge [Din03, Moo02]. Brittle [BD09]. Brownian [BDK09, DMPD09, FM04, Kry03]. Bubble [BV00a]. Bubbles [AHM09]. Buckley [vDPP07]. Buckling [HM09b, Pel01, Vui03]. Buffered [Tsa07]. Bulk [Bab08, CS07, SS03a]. Bulk-Superconductivity [SS03a]. Burgers [El 07, KMX08, KT01]. Burgers-Type [El 07]. Bursts [PS06]. BV [CRTW05].

C [FG09]. C1 [HNZ06]. Caffarelli [DGP04]. Cahn [EGW06, Tan07]. Calculus [DS08, OW05]. Calderon [Bel03]. Calibration [Cré03]. Call [YYW08]. Camassa [Bjo08, dMKST09]. Canard [KS01]. Capillary [CN00, FS03, FS05, GW07, KP07, Sch05, Wah06]. Capillary-Gravity [Wah06]. Carleman [Zha08]. Cascade [Han08]. Case [Bou08, CY04, DFLM09, FH01, Fou06, GG07, Hat00, Ing00, Pla09a, WXM06]. Cases [BD09]. Casimir [Rei01]. Catastrophic [Dum03]. Cauchy [CD01, CG08, CA00, Mik02, MR02, PP03]. Cell [BFG04, MOPMW06]. Center [FHW02, Kri04, KS00]. Certain [BS08, DD02a, KK02, KK05, MV05, Nis02]. Chain [HKP07]. Changing [WXM06]. Channel [LS01]. Channels [Cla00, EL07]. Chapman [LZ03]. Character [LLYZ03]. Characteristic [Tra06]. Characterization [Auc06, Brü01, KLVZ09, LW09]. Characterizations [Kyr04]. Characterizing [Pro00]. Charge [FG09]. Charged [CCD02]. Charges [EL07]. Chemical [FGHS07]. Chemosensitive [HKS05, HKS07]. Chemostat [WNW07]. Chemotactic [Wan00]. Chemotaxis [BDDS06, Dun00, TW09, Wan00, YCL01]. Chemotaxis-haptotaxis [TW09]. Chip [CF01]. Chromatography [ACFS09]. Chronic [MOPMW06]. circles [FKW06]. Circular [AK06, BV00b]. Clamped [CMM06]. Class [AS06, AG02, BP07, CK01a, CP07b, CD01, DIN04, DFM09, DZ08, Gev08, Gui07, HR06, Le05, RW00b, VZ07, vDPP07, Mus05]. Classical [BPR00b, LZ05, Lil07, WW07]. Close [LMR08, SS03a]. Closure [DE00]. Cluster [WW02]. Coagulation [BL07, LLW03, LW05, MP05]. Coagulation-Fragmentation [LW05]. Coal [MB08]. Coarsening [CNO06, DP05, ORS06, Sle08]. Coast [JLP06]. Coefficient [KS02b]. Coefficients [Ban03, BIY08, CK01a, CY04, CR05, GJX08, KMS07, KK07, Kry09, Mal00, WZF08]. Coercive [San05]. Coexistence [Dun00, HS02]. Coinciding [Bia01]. Collisions [CGL08, HT09]. Column [CM98, CM00]. Combined [TW09]. Combustion [Ala06, BW04, Du00, LZ03]. Comment [Pla09b]. Compactly [DGH00, HS08]. Compactness [BCD05, Rei01, SR09]. Comparisons [BPR00b]. Compartmental [MVNO08]. Competing [HLMP03]. Competition [FKK09, GR03, HS02, MT09]. Competition-Diffusion [MT09]. Complete [LRRBS09, Pro00]. Completeness [Pan09]. Complex [BBP05, KPT09, MT03, UW07]. Complex-Valued [MT03]. Compliance [KS02b]. Composite [BV00b]. Composites [BN02, Gra09]. Compositions [CDD03]. Compressibility [Ste04]. Compressible [AS07, CZZ07, FS00, GJX08, HZ03, Hof06, HW09, HMS03, HMS04, JNS06, JZ09a, KK04b, LXY00, MP01, MV08, NN09, NN01, NYZ04, PLN02, PW08, Per06, PRS08, QW09, WX05]. Computation [DKS02, ZD00]. Computer [AFM01]. Concentrated [BBP05]. Concentrating [DW05]. Concentration [DF06, DH08a, DH08b, MM05, Rei01, Tzi06]. Concentration-Compactness [Rei01]. Concentrations [Lip06]. Concept [BCN02]. Condensates [AB06]. Condition [Cas03, FLP05, Hat00, HSZ03a, HSZ03b, HK00b, II01, Kim09, MW09, MV05, Oga03, SZ07, Kim13]. Conditions [ANR09, AMRT09, Bos03, Cha06, Ebm00, EGG03, GIK05, GIK08, GJT06, HR06, HR08, IV04, Kel06, Lew01, Lew05b, Mar08, MS04, NRJ07, O’D08, Pis04, Xu04]. Conducting [Gra09, JNS06, JZ09a]. Conductivities [KY03a]. Conductivity [AG01a, AK03, Bri03, Gra09]. Conductors [Lip01]. Cone [GH09, KK08b]. Configuration [God09]. Configurations [GV00]. Confined [CCD02]. Confinement [AMP05]. Conic [GH09]. Conjecture [AH07]. Conjugate [HM09b]. Connected [EK08]. Conservation [CWY09, CR05, CHS08, CS06, Dal07, Ha01, HY03, HL02a, HL02b, IV04, Jen00, KL02, Laf04, Lew05a, LS03, LT01, Nak03, NT08, Pan09, Roh05, Rou03, SP07, TT00, TTX03, vBPW08]. Conserved [Lau02]. Conserving [CGL08]. Constant [CG00, LM09]. Constrained [Ped00]. Constraint [BPBW05, CS07, LP08, MST07]. Constraints [BFG04, PS09b]. Contact [BJ09, KS02b]. Containing [CO05, Lew05a]. Continuation [CL00, HK00a, HW03, Isa01, Lil07, MH06, NW03]. Continuity [ACFS09, JW07, KK02, KK05]. Continuous [FLT09, Han08, Ing00, JJS00, Pon07]. Continuum [AC04]. Contracting [MPS06]. Contraction [OW05]. Control [LP08, Rou06, VZ09]. Convected [GP05, GP06]. Convection [CY04, ZY00]. Convergence [ABV06, BL06, BN08, BCM00, BO07, BCD05, CN06, CHK08, God09, GR03, GWZ09, GK08, GK00, HSZ03b, Hor08, HW02, Ish05, KMX08, LYY07, MP03, MP05, NN09, NS05, Pot07, TTX03, aLW09, CMS04, Ped04]. Convex [DL04, DGG00, Hwa04, JS02, KS05a, Bab08]. Convexified [IM04]. Convexity [DS08]. Convolution [LT01]. Coordinates [Lee09, Ste04]. Copolymer [RW03, RW05, RW08]. Cord [BFG04]. Cores [PS06]. Corrected [CM00]. Corrections [Wil09, Wri05]. Correctors [DFM09]. Correlations [BBN08]. Corrigendum [Kim13]. Cortex [CMS04]. Couette [KTvW05]. Coulomb [Ren06]. Coupled [BS08, GW06, Kov08, Sch02]. Coupling [BNR05, HR06]. Couplings [LAJ08]. Crack [BFV08]. Cracks [Gia05, Ron05]. Criteria [BK02]. Criterion [CCM07, DdlL00, Hat00, Kim06, LHK07, LTT03, Ren06]. Critical [AH08, CCP07, CNO06, CA00, DD02b, GPV00, HK06, HK00b, KLVZ09, LM08, LW08, LT01, SS03a, VZ07, WW02]. Cross [BV00b, CJ04, GJV03, Le05]. Cross-Diffusion [CJ04, GJV03]. Cross-Section [BV00b]. Crossings [KL08]. Crystal [Pon07]. Crystalline [CN06, GGH05]. Crystals [CGLL02, CP07a, GV00, LP07, Pan06]. Current [BNR05]. Currents [Alm08, JS02]. Curvature [BCQ01, CN06, II01, IM04, Ish05, Nar08]. Curve [Du00]. Curved [JLP06, MV07]. Curves [DKS02]. Cusp [HP09]. Cylinder [FS00, PP00a]. Cylindrical [JZ09a].

D [Kim13, Bos03, CF01, FH01, Kim09, LY05, MPP09, ST01]. Dafermos [LS03]. Damkohler [MDvD06]. Damping [DP09]. Darcy [Váz03]. Data [BIY08, Esp05, GH08, Lor00, NS07, NW02, NW04]. Davey [VA03]. Dead [PS06]. Decay [AGSS00, Bjo08, BY04, Deu09, HZ09, JM08, KK08b, Nak03, PP00b, Sch09, WY08]. Decomposition [JT09]. Deconvolution [DE06, Lay07]. Decrease [SS03a]. Deep [Hur06]. Deep-Water [Hur06]. Definite [BBW09]. Defocusing [KPT09]. Deformable [TY05]. Deformations [LL05]. Degasperis [WY08]. Degeneracies [DD02a]. Degeneracy [Xu03]. Degenerate [BK04, BK00, BFK03, GW06, HLMP03, PV05, Ria04, Sch08, Tak00, WZF08]. Degenerate/Singular [PV05]. Degree [CGLL02]. Delay [BK02, CLGZ06, LRR06, STZ02]. Delayed [HW03, LWS05, OW07, SZ00, WLW09]. Delays [CLGZ06]. Delivery [FP02]. Denoising [All08]. Densely [BN02]. Density [BIY08, DZ03, GP05, GP06, GJX08, PV05]. Density-Dependent [GJX08]. Dependent [BK02, Ble08, BMR09, CR05, Deu09, FMS03, GWZ09, GJX08, Sac04]. Depth [Sun01]. Derivation [CMM06, MCP08, PW08]. Derivative [CKS^{+}01, CKS^{+}02, GH08, HNS08, KK08b]. Derivatives [AGSS00, BG08, PP00b]. Derived [IK01]. Dermal [FP02]. Derrida [JM08]. Describing [BFG04, FM09, Joc02]. Description [AK06, Mei08]. Design [Ped00]. Deslauriers [dVGH03]. Destabilization [DIN04]. Detecting [AMR02]. Determination [AD05, BFV08, CGL08, CY04, KY03a, NS07, Sin06]. Determining [AG01a, BIY08, CK01a, FPR04]. Deterministic [Sch05]. Development [AK00]. Devices [GG07]. Diameter [AK03]. Diblock [RW03, RW05, RW08]. Dielectric [AKS06, SZ05]. Diffeomorphisms [ESG05]. Difference [BW08, Mal00, PR01]. Difference-Algebraic [BW08]. Differentiable [DP03, SZ04]. Differential [BPBW05, BK02, BLM09, CLGZ06, GH03, GG07, GR03, HKP07, KS06a, KKSY02, KK07, KW01, KO00, LWS05, LTT03, LR09, MVNO08, NOV09, Pal09, WCD07, WLW09]. Diffractive [AR03, BL03]. Diffusion [AK00, Abd03, BDR09, BNR05, Ble08, Bos00, BDDS06, CCK07, CJ04, CQ07, CEGMM09, DE09, DFG02, FGHS07, GPV00, GJV03, GIK05, GIK08, GH05, HKM08, HS02, JZ09b, KL02, LRRBS09, Le05, Li08, MT07a, MB08, MT09, NRJ07, Osh04, Roh05, SZ00, Váz03, Wan00, WXM06, WW02, YCL01]. Diffusion-Dispersion [KL02, Roh05]. Diffusions [BX09, FLT09]. Diffusive [FGQ04, KT01, MP01, Sch00]. Diffusivity [BDR09]. Dihedral [EK08]. Dimension [Bab08, BGM08, CD08, HIN04, Mar06a, Mas01, Ryb01, Tzi06]. Dimensional [Alì03, ABV06, BCR01, Bel03, BCV06, BKL04, Cas03, CWY09, CGZ09, CY04, CN00, CS06, DFLM09, DG09, ET09, FLP05, Fil07, HK00a, HT09, HZ03, HP09, HL02a, HL02b, JNS06, KS09a, Kim02, KSS02, LZ05, LP02, Lim02, LP09, MT03, MV08, MH01, Nak03, NS09, NS04, PLN02, Pin00, Pla09a, RW03, Sac04, SZ09, Xu04, Yua06]. Dimensions [FLT09, Kan04, KS01, Lie05, VZ07, Zha03a]. Diode [GSZ04]. Dirac [Bou08]. Directional [Bri03, KS09c, Zui00].

Dirichlet [AMRT09, CY04, EK08, HN05, Ing00, KY03a, KM01, MS04, ZY00]. Dirichlet-to-Neumann [Ing00, KY03a]. Discontinuities [Tra06]. Discontinuity [AFM01, CWY09, Mor03]. Discontinuous [CR05, KS02b]. Discrete [AC04, CT09, Ing00, JJS00, Kat05, KPS09, LAJ08, Mac04, Nis02, Pon07, RL06, VA03, FKW06]. Discrete-Time [LAJ08]. Discretization [Lim02]. Discretizations [GH03]. Disk [KS09a]. Dislocation [BCLM08, El 07]. Dislocations [CL05, FM09, GM05, Pon07]. Disorder [CP07a]. Dispersion [Ban03, KL02, Kur04, MJZ04, Roh05, WZ05]. Dispersion-Managed [Kur04]. Dispersive [CL09, GZ09, KS07, MR02, Pro02, Sze05, Mus05]. Displacement [BF06, DS08]. Dissipation [Lay07]. Dissipative [BNR05, Cha06, HNS08, MR02, SS03b, Tra06, Wu05, WY08]. Distance [CDJ08, DS08, OW05]. Distribution [AH07, Bri03, GW07, The04, Van01, Van85]. Distributional [JJS00]. Divergence [Kry09]. Divergent [LY09]. Domain [ABDG04, AC07, CCK07, Fil07, FLM01, GG00, GP05, GP06, Hwa04, Kel06, Kim02, Mik09]. Domains [Ble08, CLR02, DL04, DG00, Ebm00, EK08, Han01, HR08, HN05, KK04a, KS09b, MV07, NRJ07, Pla09b, Ria04, WCD07]. Dominant [MDvD06]. Doniach [BK05]. Doping [WXM06]. Doppler [Pal09]. Doring [LZ03]. Double [MPS05, MPS06, Sac04]. Double-Well [Sac04]. Doubly [EK08, Vis07]. Down [GK08]. Drift [BNR05, BDK09, WXM06]. Drift-Diffusion [BNR05]. Driven [DDT04, DFG02, GSR05, LR09]. Droplet [Grü03, ORS06]. Drug [FP02]. Dual [CDY04]. Duality [BV02]. Dubuc [dVGH03]. Ducts [Yua06]. Duffin [Nik01]. Dynamic [BJ09, EGG03, KS02b, MN00, MP08, SZ09]. Dynamical [BS08, DS06, IWY08, LAJ08, MZ05a, MP05]. Dynamics [AG01b, BCC03, BV05, BV00a, CK01b, CFG06, DLS05, DKP07, El 07, Fil07, FM09, GW07, GSZ04, HW02, JZ09b, KK08a, KGB06, Lim02, NT08, Tri08, Wan00, WCD07]. Dynamo [GV05].

e-MHD [PW08]. Ecological [GH03]. Ecology [FKK09]. Eddy [DE06]. Edge [CDY04, KS02a]. Effect [AM05, GV05, Gra09, Pan03, WNW07]. Effective [BBP05, Gra09, NRJ07, WCD07]. Effects [DD02a, Gia05, Wan00, Xu03]. Eguchi [HIN04]. Eigenvalue [BCN05, EK08, HKK01, Rob04]. Eigenvalues [EL02, JN04, JLP06, KM01, KS03a, PS08a]. Eikonal [MT03]. Einstein [AB06]. Elastic [AMR02, Are01, BL06, BF06, BFV08, BJ09, CP07b, CCS07, CMM06, DKS02, Gra08, HM09b, LL05, Pon07]. Elasticity [Hor08, KN00, Kim00, MH06, NW03, SS09b, UW07, Wan08]. Elasto [Mie04, NK08]. Elasto-Plasticity [Mie04, NK08]. Elastodynamics [Rie03]. Elastoplasticity [Ste08]. Electric [Alm08, HHR09, HHR11]. Electrical [Brü01, KKSY02]. Electromagnetic [AN00, AKS06, Joc05]. Electronic [FG09]. Electrostatic [GG07, Li09, Li11]. Element [BFP08]. Elementary [HM09a]. Ellipse [BPR00a]. Elliptic [AGGM05, BCV06, Bla06, BV00b, CF03, CE07, DRY07, DG00, DLS05, Du00, Ebm00, GW09, KMS07, KY03b, KK04a, KK07, LYY05, Lip01, Lóp00, MG05, NK08, Osh07, PR01, PS06]. Elliptic-Hyperbolic [CF03]. Elliptical [BPR00b]. Emission [MN00]. Energetic [MT07b]. Energies [AC04, Bab08, DF06, Ort06]. Energy [BL06, BCM00, CGL08, CL05, CS07, CM06, CR03, CO05, GP05, GP06, HLTT08, Joc05, KTvW05, KS09b, Lay07, Li09, Li11, LL00, MS06, MM05, Pon07, Rei01, RL05, ST08, VZ07]. Energy-Casimir [Rei01].

Energy-Conserving [CGL08]. Enhanced [MB08]. Ensemble [DN09]. Entire [MT09, WLW09]. Entries [JN04]. Entropy [BK04, CD01, Hat00, Pan09, vDPP07]. Equation [AMSW05, Abd03, AH08, AP06, ABDG04, AMRT09, BL07, BV08, BNP06, BD06a, BC05, BD02, BFK03, CCG05, CW00, CQW03, CO06, CHK05, CCP07, CDM08, CT09, DW05, DDT04, DG09, Esp05, FH07, GPV00, GG00, GH08, GJT06, GW09, GZ09, HIN04, Han01, HNS08, JW07, Joc05, Joh09, JP00, JM08, JM09, JLM01, Kag01, KK08a, KMX08, KK08b, KKS00, KS06a, KPS09, KT01, KKSY02, KY03b, Kim05, KN09, KPT09, KP08, Kry01, Kry03, Kur04, LLW03, Lau02, LW05, LRR06, Len08, LL06, Li09, Li11, LP09, LW09, MT03, Mal00, MS04, Mik09, MJZ04, MR02, Moo02, Nak01, Nis02, NS04, OT07, PP03, Piv00, Sac04, SS09a, Sze05, Tak00, Tzi06, VA03, VZ07, Wu05, WY08, Yin04, dMKST09, vDPP07, EGW06, FKW06]. Equations [AK00, ACFS09, AS07, AGSS00, AG01b, BT07, Ban03, BDR09, BS00a, BS01, BCLM08, BPBW05, BW08, BK04, BV05, BS00b, BK00, Bjo08, Bos00, BC06b, BG08, BKL04, BLM09, BSB03, CCM07, CK01a, Car03, Cha06, CM01, CD01, CL03, CMZ08, CCS07, CT08, CP00, CKS^{+}01, CKS^{+}02, Coo00, CMJ09, CA00, CF06b, DGP04, DS06, DLS05, DGG08, DS05, DFM09, DMPD09, EGG03, ET09, Fou06, FS00, FC06, GZ00, GH03, GG07, Gom03, GP04, GR03, GWZ09, GK05, HKP07, HK06, HZ03, Hof06, HL06, HZ08, HKS05, HKS07, Ift02, IL05, JJS00, JNS06, JZ09a, KK02, KK05, Kel06, Kim02, KK04a, Kim06, KK07, Kim09, Kim13, KW01, KKSW08, KA00, KS07, KPT09, KO00, LWS05, Li05, LY05, Lil07, LP02, LW08, Lip01, LTT03, Loe06, LS01, LCS03, LR09]. Equations [LXY00, Mar08, Mar06b, MV08, MH01, MP05, MS06, MB08, Mik02, MR04, Miz01, Miz03, MST01, MST07, MT09, MVNO08, NYZ04, Nob09, NOV09, PW08, Per06, Pin00, PRS08, PS06, QW09, SW09, Sch00, Sch08, SZ07, ST01, SZ00, Spi03, Tan07, VZ09, Váz03, WX05, WCD07, WLW09, WZF08, YCL01, YZ05, Zha03b, Zha08, vdBHV01, BC06a, Mus05]. Equidistributed [CF06a]. Equilibria [CJT09, FL00, MR08, Ort06, PS08b, Ria04]. Equilibrium [GV00, GWZ09, HZ09, NS05, Wal09a]. Equivalence [JLM01]. Equivariant [AHM09]. Ergodicity [GK08]. Erratum [HHR11, HKS07, KK05, Li11, Van01]. Error [BN02, BGL07, Laf04, NW02, NW04, TT00, Wil09]. Errors [JWW07]. Estimate [BY04, CCM07, Laf04]. Estimates [BGL07, CQ07, CD08, HH01, JW07, KM01, KK08b, Lor00, NW02, NW04, PV05, SS09b, TT00, Wil09, Zha08]. Estimation [Rou06]. Euclidean [BGL07]. Euler [ABR00, Alì03, BCN02, CL03, CZZ07, CT08, ET09, HL06, Kim02, Lil07, Loe06, NT08, PW08, Yua06]. Euler/Euler [NT08]. Eulerian [DS08, OW05]. European [YYW08]. Evaluation [BCR01, CDD03]. Evans [BGSZ01, KS02a]. Evolution [AMRT09, BFG04, CGLL02, CWY09, DKS02, GL08, GP05, GP06, BC06a]. Evolutionary [MT07b]. Exact [Du00, Gra09]. Exactly [MC01]. Example [BCP06, Dum03]. Exchange [Joc05]. Excited [KW01]. Exclusions [Tor04]. Exercise [BX09, CC07]. Existence [AS06, ABR00, Alì03, BCLM08, BR06, BS00b, Bos05, Cas03, CDJ08, CWM05, CdMGGK09, CT08, CJT09, CEGMM09, CDM08, DZ03, DH08a, DKS02, EM09, ET09, Fig07, FM09, GH05, Gra08, Hat03, HKP07, Joc02, JLP06, JM08, Kat05, Kur04, LZ05, MT03, MV08, MN05, MS06, Mie04, MPP09, MB08, Mus05, Per06, PS09b, Pis04, Rög05, RL05, ST08, Sch09, SZ09, ST01, Sun01, WW07, YCL01]. Expanding [GGH05, MPS05]. Expansion [Bou00, KY04]. Expansions [AL06, AK03, CDD03, Dun01, GK00, Lóp00].

Explicit

[BCN05, Brü01, CWM08, DZ08, FG09, ZD00]. Exponent [GPV00, GW09, HK00b]. Exponential [AGGM05, DDGM07, HZ03, KMS07, MV05, NOV09, QW00]. Exponents [LM08]. Extended [HM09b]. Extending [BGL07, KS01]. Exterior [GG00, KN00, MS06]. External [DO09]. Extinction [HS02]. Extra [FM04]. Extracting [CDY04]. Extremal [BD09, EK08].

F [FG09]. Faber [DK08]. Factorization [RTW01]. Families [CL04]. Family [CL00, FPR04]. Far [Sin06]. Fast [BCR01, GPV00, Váz03]. Fast-Diffusion [GPV00]. FDEs [FHW02]. Feedback [HW03, STZ02]. Feline [FLM01]. FeLV [FLM01]. Fermion [NS05]. Ferromagnetism [Joc02]. Fiber [BD09, Gra09, GK08, LP00]. Fiber-Reinforced [Gra09]. Fibers [BV00b, Bri03]. Field [BF06, DP05, DD02b, DO09, GSR05, GM05, GP04, Joc05, KSS02, LP07, Lip06, SS03a, Sin06, Wan08]. Field-Induced [LP07]. Fields [Lee09, LW09, Pan03, Sch02]. Film [BL06, BS00b, CD08]. Films [CM06, DD02b, GW06, Hor08, Pan03]. Filtration [BFK03]. Finance [FLT09]. Finite [BFP08, FLT09, FS01, Grü03, HZ03, KGB06, Mie04, Sun07, YZ05, dVGH03]. Finite-Dimensional [HZ03]. Finite-Time [YZ05]. First [Abd03, CCCS06, EK08]. Fixed [AIM03, Pro01]. Flat [LYY07, Tak00]. Flow [AC08, AHM09, BR03, BNP06, BDV03, CP07a, CN06, CGZ09, CGP05, DMP06, EKK09, FLP05, FS03, FS05, GH09, GGH05, HR06, HZ03, Hof06, HP09, KTvW05, MP01, MDvD06, Nar08, NN01, Oga03, Rög05, Sch08, IM04]. Flows [Ala06, AS07, Amo01, AHT03, BFP08, BMR09, CCG05, CR06, CR03, CZZ07, CF01, Deu09, DE06, EM09, ESG05, FMS03, FS00, FM01, GJX08, HW09, JZ09a, Kan04, KK04b, Lin03, LY09, MR04, MT08, Ort06, Per06, SS03b, Sch09, Sun01]. Fluid [AG01b, Dal09, DLS05, Gra08, HT09, HP09, MP08, MT08, NN09, Per06, Rou07, ST08, Sun01, WC09]. Fluids [BP07, BMR09, CR06, CM01, CL03, FMS03, FGM05, FS00, HZ09, JNS06, ZZ08]. Flux [AC07, Cas03]. Focus [BD06b]. Focusing [Dum03, KLVZ09, VZ07, SZ05]. Fokker [Kag01]. Fold [KS01]. Food [OW07]. Food-limited [OW07]. Föppl [CMM06]. Forced [BB08, DS05, HK00a, LLYZ03]. Forces [CL05, WX05]. Form [Kry09, Pal09, QXM08, Xu04]. Formal [FHW02]. Formalism [Fra07, LL05]. Formation [CdMGGK09, CL03, CWY09, HP09, HW03, Ren00, Vis09]. Forming [DS09]. Forms [BW08]. Formula [BF06, BCQ01, QW00]. Formulation [Dal07, IV04, MT07b]. Formulations [Wan08]. Four [BCR01]. Four-Dimensional [BCR01]. Fourier [FN05, FS09, PN08]. Fourth [AGGM05, GB04, GJT06, GW09, JP00, SW09, vdBHV01]. Fourth-Order [JP00, GJT06]. Fractal [KMX08]. Fractional [DMPD09, TTX03]. Fragmentation [LW05]. Frame [Che00]. Frames [Joh03, NPW06]. Framework [BL03]. Free [AFM01, BFG04, BV00a, BW04, BFK03, CKK06, Cas03, CF00, CF03, CE07, Cui08, FH01, FP02, FH07, Gor02, Gui07, HMS03, HMS04, KS03b, LS00, Li09, Li11, MS04, Mor03, Nak01, PP04, PLN02, RW08, WNW07, WC09]. Free-Boundary [Cas03]. Free-Discontinuity [Mor03]. Freezer [KRS09]. Frenkel [QXM08]. Frequency [AN00, AIM03, Fou06, KS05b, Mac04]. Freudenthal [HNZ06]. Friction [FLP05, Kim09, Kim13, KS02b, Ren06]. Frictional [DP09]. Friedrichs [BL01]. Front [Laf04]. Front-Tracking [Laf04].

Fronts

[DIN04, GR00, GR03, Sch05, Tan07, Tsa07]. Fuel [The04]. Full [QW09]. Fully [Gor02, IK01, KA00, Loe06]. Function [BPR00a, CDY04, DGH00, FPR04, KS02a, Kyr04, PS00, YZ09, Yoo01]. Functional [ABV06, BNP06, BDK09, BSB03, CRTW05, CO05, CMS04, DGG08, DO09, GR03, KTvW05, Lew05a, LTT03, MVNO08, NOV09]. Functionals [BN08, CH06, CF06a, MM05, Mor03, Rei01, Tzo06, Zag05]. Functions [Auc09, BCR01, BGSZ01, Can01, CGL04, CDD03, CLR02, CDY04, Dun01, EL02, Han08, KO00, MV05, NW02, NW04, Pla09b, Pro00, ZS02]. Fundamental [HKK01].

Gamma [BL06]. Gamma-Convergence [BL06]. Gap [CGL08, MM05]. Gas [BV05, HW02, HMS03, HMS04, PLN02]. Gas-Solid [HMS03, HMS04]. Gases [Bou00, GSR05]. Gauss [IM04]. Gaussian [DLY07, RW00b]. Gaussian-Based [DLY07]. General [AC04, BL03, Bri03, FHW02, Gra09, KS05b, MZ05b, WXM06]. Generalizations [BD02]. Generalized [BPR00a, CHK05, CCP07, Cré03, DS06, Dun01, Joh09, KTvW05, KO00, LL06, Mar06b, Miz01, Miz03]. Generating [Mor03]. Generic [FL00, Fra07, KL08]. Gennes [Pan03, Pan06]. Genuinely [Gev08]. Genus [Vui03]. Geometric [All08, AR03, BK02, Bos00, KS01, Len08, MH01]. Geometrical [CS07]. Geometry [BN02, BBP05, BD06b, TY05]. Geophysical [AG01b]. Geostrophic [AH08, SS03b, Wu05, Cha06]. Gibbs [CHK08]. Gierer [CK01b, DKP07]. Ginzburg [Spi03, BCM00, DF06, DO09, GWZ09, LL00, Pan06, SS03a]. Glass [Vis09]. Global [AH08, ABR00, Alì03, AS07, BCLM08, BNP06, BC05, CdMGGK09, CO05, CHK05, CKS^{+}01, CKS^{+}02, CA00, DP09, ET09, GH09, GH05, HK00a, HK00b, HL06, Hur06, HKS05, HKS07, JP00, LZ05, LY09, LTT03, MZ05a, MH06, Mar06a, MV08, MR02, NS05, Per06, ST08, ST01, SZ00, WW07, WZF08, Wu05, YCL01, EGW06]. Gluing [MNR05]. Goes [IL05]. Gordon [OT07, Kat05, OT07]. Gradient [BNP06, CF06a, GL08, HK00a, MH06, Ort06, Zha03a]. Gradients [Lip01]. Granular [GJV03, Gor02]. Graph [Piv00]. Gravitational [LMR08]. Gravity [CN00, GW07, Wah06]. Gravity-Capillary [GW07]. Green [PS00, YZ09]. Greene [DdlL00]. Gross [KKSW08, Mar06a]. Ground [FG09, God09, LMR08, Rou06]. Group [KO00, ZD00]. Grow [KS09a]. Grow-Up [KS09a]. Growth [AC04, CF03, CE07, Cui08, DF06, FH07, Gia05, MW09, WC09].

Half [ANR09, BBW09, Kan04, Nak03, NN09, Nis02, WX05]. Half-Line [BBW09]. Half-Plane [WX05]. Half-Space [ANR09, Kan04]. Hall [Gra09]. Hamilton [BS00a, DS06, DS05, EGG03, Gom03, Pin00]. Hamiltonian [BPBW05, BB05, DL00, KW01, LL05]. Hamiltonians [Pis04]. Hand [BG08]. Haptotaxis [WW07, TW09]. Hardening [Ste08]. Hardy [VZ09, ZS02]. Harmonic [AMSW05, AHM09, Auc09, BFP08, Car03, CWM08, HLTT08, HSZ03a, HSZ03b, KKS00, MV05, Oga03]. Harmonics [ZD00]. Hartree [CO06]. Hartree-Type [CO06]. Hats [Tak00]. Hausdorff [CD08]. Having [CF06a, CLR02, Pla09b]. Healing [CF00]. Heart [PSF05]. Heat [BFP08, CK01a, ESG05, JNS06, JZ09a, KKS00, KN09, Kry01, Kry03, Lip01, Oga03]. Heat-Conducting [JNS06, JZ09a]. Hele [EM09, CR06, GP05, GP06]. Helical [ET09]. Helicity [Lay07]. Helium [God09]. Helmholtz [Fou06]. Hemivariational [MO09]. Heterogeneous [FLM01, KP08, Lip06]. Heuristic [DE09].

High [AK03, Bri03, Fou06, Mac04].

High-Conductivity [Bri03].

High-Frequency [Mac04]. High-Order [AK03]. Higher [AGSS00, DFLM09, HK00a, KN09, LY05, MJZ04, Mos05, NS04, VZ07, Wil09]. Higher-Gradient [HK00a]. Highly [ABDG04, FLM01, WX05]. Hilbert [Auc09, DL04, DMPD09]. Hill [GZ00]. Hilliard [EGW06]. Hinged [PS09a]. Hölder [Han08, HN05, KK02, KK05]. Holes [Fil07]. Holm [Bjo08, dMKST09]. Homoclinic [MPS05, MPS06, Rob00, SW09]. Homogeneous [Bra05, Mal00]. Homogenization [BD09, BGM08, BF01, Bri03, CDD06, CDG08, COV02, Dal07, DFM09, FS09, GP04, JT09, KMS07, Lip06, Lou06, MCP08, Mei08, MT07b, Nes07, Ped06, PR01, Sch08, Vis07]. Homogenized [BR03, BDK09]. Hopf [FL00, KA00, SS08]. Hopping [KL08]. Hulst [BL07, LLW03]. Hunter [BC05, Len08, Yin04]. Hybrid [BNR05]. Hydrodynamic [Xu09]. Hydrodynamics [CFdlL04]. Hydrostatic [AG01b, BKL04]. Hyperbolic [BR03, Bia01, CD01, CF03, CWY09, Cui08, DFM09, FL00, FC06, GR00, Gev08, GWZ09, HY03, HL00, HK00b, HL02a, HL02b, Joc02, Kim00, KA00, LRR06, Li08, MS04, Ren00, Tra06, vBPW08]. Hyperbolic-Parabolic [CD01, vBPW08]. Hyperelastic [CHK05]. Hyperelastic-rod [CHK05]. Hypergeometric [BPR00a]. Hypoplastic [Gor02]. Hypothesis [JWW07]. Hysteresis [AM05, EKK09, GJS09, KSS02]. Hysteresis-Type [GJS09].

Icosahedral [ZD00]. Ideal [CCM07, Lin03]. Identification [KY03b, KP05, LS00]. Identity [ZD00]. II [DH08b, FL00, HSZ03b, HL02b, MT03, VA03]. Ill [MST01]. Ill-Posedness [MST01]. Image [All08, CRTW05, DFLM09, GB04, SZ04]. Images [GM01]. Immediate [FMP05]. Impact [AS06, CLR02, Pla09b, QT05]. Impedance [Brü01, HHR09, HHR11, KKSY02, Sin06]. Imperfect [DFM09]. Implicit [Li09, Li11]. Inclusion [AMR02, AD05]. Inclusions [Brü01, UW07]. Incompressible [FLP05, FM01, GV05, Kim06, LZ05, Loe06, PW08, Rou07, ST08]. Increasing [CR06]. Increasingly [LYY07]. Incremental [Mie04, NK08]. Individual [FKK09]. Induced [Ble08, LP07, MPP09, Pon07]. Inequalities [Ban03, BDK09, DGP04, Lip01, MT07b, Nik01, VZ09]. Inequality [BPR00a, DK08, MO09, Oga03, YYW08]. Infinite [FLT09, JN04, MH01, ST08, Sun01, Van01, Van85]. Infinitely [Kur04]. Infinitesimal [NK08]. Infinity [IL05, Mar08]. Inflow [QW09]. Influence [BFG04]. Inhibitor [WNW07]. Inhomogeneities [AK03, AIM03, BF06]. Inhomogeneous [CEGMM09, FMP09, GSR05, HH01, PRS08, dPKW07]. Initial [Bos05, FS05, GH08, LP02, NP00, PP00b]. Initial-Boundary [Bos05, PP00b]. Initial-Value [NP00]. Initiation [FFH02]. Innovation [Sun07]. Instabilities [LP07]. Instability [BD02, CCP07, DFG02, LP00, Lin03, OT07]. Installment [YYW08]. Instantaneous [HP09, STZ02]. Integrable [GMS09, KS02a]. Integral [AC04, Bab08, MM05]. Integrals [CDD06, DGG00, Lóp00, RW00b]. Integrodifference [HZ08]. Intensity [CDY04]. Interacting [CKK06, CCS07, FGQ04, Rou06]. Interaction [CS07, CO05, God09, Gra08, HT09, MC01, Miz03, SZ09]. Interactions [WZ05]. Interface [AK00, BL06, DO09, LXY00, NRJ07, O’D08]. Interfaces [DFM09, Osh04]. Interfacial [Sle08]. Interior [CK01b, dPKW07].

Intermediate [CGL05]. Interpolants [LYY07]. Interpolation [DY00, NW02, NW04, Yoo01, dVGH03]. Interpolatory [DLY07]. Interval [Van01, dVGH03, Van85]. Invariance [Kri04]. Invariant [BCV06, BSB03, CL00, Kag01, Kim05, BW08]. Inverse [BT07, BIY08, BBW09, CCCS06, HH01, HHR09, HHR11, IWY08, KK08a, KKS00, Li05, O’D08, Piv00, Ryb01]. Inversion [CN01]. Inviscid [FLP05, Kim09, Kim13]. Involving [BPR00a, BPR00b]. Ion [EL07]. Ionized [Rou07]. Irregular [BN02]. Isentropic [BV05, CL03, CGZ09, GJX08]. Isolas [BKL^{+}09]. Isolated [KGB06]. Isopolar [Zwo01]. Isosceles [CL04]. Isospectral [Amo01]. Isothermal [Hat00, HW02, Xu09]. isotropic [Li05, HL02a]. Issues [MST01]. Iterated [Abd03].

J [HKS07, Pla09b]. Jacobi [BS00a, DS06, DN09, DS05, EGG03, Gom03, JN04, Pin00]. Josephson [CFTV03]. Jouguet [LZ03]. Journal [CJT09]. Jump [BX09, FLT09]. Jump-Diffusions [FLT09]. Junction [CHS08, CG08]. Junctions [CFTV03, FM09, MP08]. Justification [AG01b, DdlL00].

Kac [CGL08]. KAM [DdlL00]. Kantorovich [AHT03]. Karman [Kim05, BJ09]. Kawahara [BD02]. KDV [Tov08, KPT09, Mar06b, Miz03, Wri05]. Keller [KS09a, BDDS06, Sug09]. Kernels [AL06, Auc09]. Khokhlov [KP08]. Killing [BFG04]. Kinetic [BV05, BG08, CS06, Dal07, HL00, IV04, JT09, NS05]. Kinetics [FGHS07]. Kirchhoff [BGM08]. Klein [OT07]. Knot [KS03b]. Ko [CP07b]. Kohn [DGP04]. Koiter [BL01]. Kolmogorov [PP03]. Kontorova [QXM08]. Korteweg [CCP07, DDT04, Joh09, MC01, Miz01]. KP [MST07]. KP-Type [MST07]. Krahn [DK08]. Kuznetsov [KP08, LP09]. Kuznetsov-Type [KP08].

l [BCN02]. Lack [BCD05]. Ladders [BKL^{+}09]. Lagerstrom [HM09a]. Lagrange [Lil07]. Lagrangian [CF06b, HL06]. Lakes [DE09]. Lamé [BIY08, IWY08]. Lamellar [RW03, RW05]. Landau [GWZ09, BCM00, DF06, DO09, LL00, Pan06, SS03a, Spi03]. Laplace [ABDG04, IL05]. Laplacian [AMRT09, CQW03]. Large [BS00a, BL03, BK00, BDK09, DS06, DLS05, DS05, DE06, FG09, Hat03, Kim09, KS09b, Lew01, MZ05b, Miz01, Moo02, Nar08, Sun01, WW02, YZ09, Kim13]. Large-Amplitude [MZ05b, YZ09]. Large-Eddy [DE06]. Large-Time [DS05]. Larmor [FS01]. Lasers [BL03]. Lateral [Kry03]. Lattice [BCC03, FKW06, KGB06, SZ09, WLW09, Zho01]. Lattices [AB06, CFG06, HLTT08, KS05b]. Law [BN08, CR06, Dal07, HL02a, HL02b, MP03, Pan09, Spi03, Váz03]. Lawrence [BK05]. Laws [CWY09, CR05, CHS08, CS06, FL00, FLT09, GMS09, Ha01, HY03, IV04, Jen00, KL02, Laf04, Lew05a, LS03, LT01, Nak03, NT08, Roh05, Rou03, Sch05, SP07, TT00, TTX03, vBPW08]. Lay [GK08]. Lay-Down [GK08]. Layer [CCK07, Lim02, LCS03]. Layered [BK05, Ing00]. Layers [JZ09a, MB08, SR09, YZ09, dPKW07]. Leads [CMM06]. Least [LLYZ03]. Lebowitz [JM08]. Left [BBW09]. Left-Definite [BBW09]. Lemma [BL01, DL04]. Length [BPR00a, BPR00b]. Leukemia [FLM01, MOPMW06]. Level [KL08]. Leverett [vDPP07]. Lewis [BB05]. Lewis-Type [BB05]. Li [FG09]. Lieberman [MW09]. Lifespan [CM01]. Lifshitz [Lau02, NP00]. Lifting [Kei01]. Light [DGG08, KK08b]. Like [FGM05, Tor04]. Limit [ABR00, BW04, CL03, DO09, DE09, FGHS07, FLP05, Fil07, FG09, HW09, HLMP03, JNS06, Joc02, Kim09, Kim13, LZ05, LS01, Lou06, MP03, MT07a, MT08, Osh07, WX05, WXM06, Xu09, Zha03b, EGW06, GM05]. Limited [NW04, OW07]. Limiting [DH08b, Oga03]. Limits [AC04, FGQ04, HK00a, IL05, KL02]. Line [BBW09]. Linear [BFV08, BL01, BC06b, BD02, BST05, BFK03, BDDS06, DPR09, Don00, FC06, JLP06, JLM01, KMS07, Kim00, Li08, LP02, Mal00, Mar06b, MW09, MC01, PP00b, QT05, Ren00, San05, STZ02, Yoc07, vBPW08, BS01, BK04, BBF06, DH08a, DH08b, FS03, FS05, Ria04, Rob04]. Linearized [CGNT08, GIK05, Hor08, LMR08, SS09b, WX05]. Lines [FL00]. Liouville [BBW09, Esp05, Piv00]. Lipschitz [BFV08, CZZ07, FMS03, HN05, Ron05]. Liquid [CGLL02, CP07a, LP07, Pan06, Tri08]. Liquid-Vapor [Tri08]. Littlewood [ZS02]. Local [AK00, CM02, CN01, CDJ08, Cré03, GH08, GK00, Han01, KY03a, KS07, LL05, Roh05, TY05, Tzo06]. Localization [AP06, GG00]. Localized [BKL^{+}09, DPR09, JLP06, NPW06]. Locally [AP06, NW02]. Logarithm [Abd03]. Logarithmic [Oga03]. Logistic [TW09]. Long [BS01, BD06a, CQW03, CO05, CT08, Joc00, Kag01, MS06, YZ09, dMKST09, vBPW08]. Long-Range [CO05]. Long-Time [BS01, CT08, Kag01, MS06, YZ09, vBPW08, dMKST09]. Look [MR08]. Loops [MPS05, MPS06]. Lorenz [MPS05, MPS06, Rob00]. Losing [CCM07]. Loss [LWS05]. Lotka [LRRBS09, MT09]. Love [BGM08]. Low [Ala06, AN00, GH08, HW09, Mus05, BCP06]. Low-Frequency [AN00]. Lower [Bab08, Xu04]. Lubricated [CJT09]. Lubrication [Wil09]. Lyapunov [CH06, Lew05a].

Mach [Ala06, CKK06, HW09]. Macro [HZ09]. Macroelements [LS02]. Macroscopic [MCP08, WCD07]. Magma [SW08]. Magnetic [DD02b, GSR05, KKSY02, LP07]. Magnetics [KK08a]. Magnetohydrodynamic [EKK09, HW09]. Magnetohydrodynamics [GV05]. Magnetostrictive [RL05]. Managed [Kur04]. Management [CCCS06, MJZ04]. Manakov [DE00]. Manifolds [BW08, Bel03, CG00, FHW02, Kag01, Kri04, KS00]. Many [Kur04]. Map [AHM09, BFP08, CY04, QT05]. Mapping [SZ04]. Mappings [CG00]. Maps [CDJ08, DdlL00, FHW02, Fig07, Ing00, Kri04]. Martensitic [CM06, Hor08]. Masks [Han08]. Mass [KLVZ09, MST07, Tzi06]. Mass-Critical [KLVZ09]. Material [BMR09, Gor02, KS06a]. Materials [BD09, CP07b, GJV03, RL05, Yoc07]. Math [HKS07, Pla09b]. Mathematical [AG01b, BMR09, CCG05, CC07, FLM01, FFH02, FLT09, IM01, PS09b]. Mather [Gom03]. Matrices [JN04, LK01]. Matsumura [HIN04]. Maxwell [DW05, GWZ09, KK02, KK05, Li05, MH01, PW08]. Mean [BCQ01, CN06, DP05, FPR04, GP04, II01, Ish05, Lee09, Nar08]. Mean-Field [DP05]. Measurable [KK07, Lor00]. Measure [LHK07, Rie03, Rig06, Zha03b]. Measurements [AG01a, AMR02, AD05, BFV08, CK01a, Sin06]. Measures [BF01, CP07a, CHK08, Kim05, Mac04, Ped04, Ped06, Zha03a]. Mechanical [BSB03]. Mechanism [LP00]. Media [AS07, BDV03, CL09, FS03, FS05, GG00, HL02a, HL02b, KP08, Li05, Lip06, MP01, Mei08, NN01, Sch05, Tor04]. Median [DY00]. Median-Interpolation [DY00]. Mediated [HKM08]. Medium [AP06, BV00b, HH01, KS05a]. Meinhardt [CK01b, DKP07]. Meissner [BCM00]. Membrane [ABV06]. Membranes [CMM06]. Memory [DFG02, MPP09, Ste05, Yoc07]. MEMS [GG07]. Merriman [Ish05]. Metaphor [DE09]. Metastability [KT01]. Method [Bel03, BF01, CDD06, CDG08, CDY04, FMS03, GK05, HSZ03b, Pro02]. Methods [BCR01, BST05, CWM05, CP00, GW07, MS06, SR09]. Metric [Kov08]. Metrics [KY03a]. MHD [CCM07, PW08]. Micro [HZ09]. Micro-Macro [HZ09]. Microlocal [Sze05]. Micromagnetism [Joc05]. Mild [Bos05]. Minimal [BD06a, CLR02, KLVZ09, Pla09b, Tor04]. Minimal-Mass [KLVZ09]. Minimax [GH03]. Minimization [BST05, CR03, Li09, Li11]. Minimizers [BCM00, CO05, CF06a, DGG00, Mie04, RL05, Tzo06]. Minimizing [AHT03]. Minimum [CQ07, Zag05]. Miscible [AS07]. Mixed [Ebm00, HR08, Roh05]. Mixtures [CNO06, FGM05, Rou07]. Mode [LP00]. Model [AM05, ABR00, Alì03, AC08, BR03, BL06, BR06, BL01, Bou00, BFK03, BDDS06, BS05, CGL08, CdMGGK09, CF00, CJ04, Cré03, CS06, DFLM09, DD02a, DKP07, DE06, FKK09, FLM01, FFH02, FM09, GM05, GL08, God09, Gor02, GR03, HM09a, HZ09, IM01, JZ09b, Joc00, KS09a, KGB06, LZ05, LZ03, Lie05, LT01, MCP08, MP08, MPP09, MB08, MC01, NS05, Nis02, ORS06, OW07, Pan06, PV05, QXM08, SS03a, Sch05, SZ09, Ste05, Tov08, WW07, WNW07, Xu09, ZZ08, aLW09, dPKW07]. Modeling [BBF06, CFTV03, Cui08, FH07, PSF05, WC09]. Modelling [CCG05, CQ07, CE07, DH08a, DH08b, YCL01]. Models [BNR05, CGL05, DP05, DZ03, FM00, GH05, HR06, HKP07, KSS02, LRRBS09, Lay07, Wal09a, WXM06, Wan08, FKW06]. Modes [DPR09, HLTT08, Kur04, SZ05]. Modified [KPT09, LP09]. Modulational [Pro02]. Moduli [BMR09]. Molecular [Li09, Li11]. Moment [KN09]. Moments [KKS00]. Momentum [CGL08]. Monge [AHT03, JW07]. Monostable [CFG06, CDM08, WLW09]. Monotone [KS03b]. Monotonicity [BPR00b]. Morphogen [aLW09]. Morphological [NS04]. Morphology [RW08]. Moser [CN01]. Motility [Wan00]. Motion [BCQ01, DMPD09, FM04, II01, Ish05, PP04, Spi03]. Motions [QT05]. Movement [HKS05, HKS07]. Moving [Ha01, HY03]. Mullins [Rög05]. Multiclass [BR03]. Multidimensional [BV05, CJ04, Cla00, Cou02, GH09, GL07, Hof06, IV04, NN09]. Multifractal [Fra07]. Multilevel [Kyr04]. Multiphase [AC08]. Multiple [DRY07, HKM08, Moo02, ST01]. Multiplicity [Du00, RW00a]. Multiresolution [KS09c]. Multiscale [ABV06, BNG04, CDD03, PSF05, ZZ08]. Multivariate [BG04]. Multiwavelet [Kei01]. Mumford [CMS04]. Mutilated [Can01]. Mutually [FGQ04]. Myelogenous [MOPMW06].

n [BCN02, FG09]. Nash [CN01]. Natural [GM01, MW09]. Navier [Kim13, AGSS00, ANR09, Bra05, BKL04, CCS07, DG09, FN05, FLP05, FM01, HZ03, Hof06, Ift02, JNS06, JZ09a, Kel06, Kim06, Kim09, KK04b, LW09, LXY00, MV08, MR04, Mik09, NYZ04, Per06, PRS08, QW09, SZ07, WX05, WZF08]. Near [HZ09, KS02a, LM09, Tak00, Tzo06, HLMP03, SS09b]. Near-Periodic [Tzo06]. Negative [GW09, STZ02]. Neighborhoods [Zha03a]. Nematic [CGLL02]. Nernst [EL07]. Network [BN02, BBP05, CGP05]. Networks [DMP06, HW03, LAJ08]. Neumann [LZ03, BBF06, BV02, CY04, HN05, Ing00, KY03a, KN00, MV05]. Neuronal [CMS04]. Neurons [HW03]. Neutral [CLGZ06, MVNO08, NOV09, WXM06]. Newtonian [BP07, CCG05, MT08, Rou07]. Nirenberg [DGP04]. Nishiura [FH01].

NLS [CGNT08, KLVZ09]. No [Pot07]. Noise [DDT04, DS05, JWW07, LR09]. Non [BP07, CCG05, Rou07]. Non-Newtonian [BP07, CCG05, Rou07]. Nonautonomous [LRRBS09, LWS05, NOV09, HS02]. Noncharacteristic [YZ09]. Nonclassical [HL00]. Nonconservative [El 07]. Nonconvex [BL06, BNP06, CM02, CP07b, IM01, LZ03, Ort06, Pis04, Rie03, TTX03, Zag05]. Nondecaying [Coz09]. Nonelliptic [Lil07]. Nonexistence [AS00, JS02, Mar08]. Nonhomogeneous [AMRT09, BB08, GJT06, JJS00, Kim06, Rob04]. Nonhyperbolic [KS01]. Noninteracting [Lew01]. Noninvariance [Kri04]. Nonlinear [AMSW05, AR03, BDR09, BW08, BGHR03, BD06a, Bla06, Bou08, BY04, BDDS06, CCD02, CCG05, CKK06, Car03, Cas03, CW00, CD01, CCS07, CGL05, CCP07, CA00, DRY07, DP03, DG00, DKP07, DY00, DE09, Dum03, DZ08, Ebm00, ESG05, FS09, GIK08, Gev08, GH08, GJT06, GW09, HK06, HK00b, HW03, HKS05, HKS07, IK01, Joc00, Joh09, JP00, JM09, KK08a, KPS09, KKSY02, KKSW08, KA00, KP08, Kur04, LW08, LT01, Loe06, MH06, Mar08, MJZ04, MG05, Moo02, Nak01, NK08, NYZ04, OT07, PP00a, PP03, Sac04, SS09a, Sch00, ST01, Sze05, Tzi06, VZ07, Vis07, ZY00, Mus05]. Nonlinearities [HNS08]. Nonlinearity [AGGM05, CDM08, DDGM07, LY05, LW08, WLW09]. Nonlocal [AMRT09, ABV06, Bos00, CEGMM09, CDM08, DPR09, DH08a, DH08b, GR03, JZ09b, Kur04, RW00a, RW08, Roh05, Sle08, Tzo06, BC06a]. Nonmonotone [HZ08]. Nonnegative [GW06, JP00]. Nonresonance [Xu04]. Nonresonant [Ha01]. Nonsectorial [GK08]. Nonsmooth [BLM09, Ebm00, HR08]. Nonstationary [DLY07, FM01]. Nonsymmetric [LK01, Rob00, San05]. Nontwist [DdlL00].

Nonuniform [Sun07]. Nonuniformly [Cou02]. Nonuniqueness [JM08]. Nonviscous [DS05]. Nonzero [Mar08]. Normal [Alm08, BW08, KS02b, Xu04]. Normality [CGL04]. Norms [IK01]. Note [Ren00]. Nozzle [CGZ09]. Nozzles [LY09]. Nuclear [FG09, The04]. Null [HK00b]. Number [Ala06, CWM08, HW09, Lie05]. Numbers [MDvD06].

o [BCN02, FG09]. Observability [VZ09, Zha08]. Obstacle [Fil07, HKK01, NS07, Sin06]. Obstacles [MS06]. Occurring [Lip01]. ODEs [Ria04]. Off [CNO06]. Off-Critical [CNO06]. Ohnishi [FH01]. Oil [BDV03, Sch08]. Oki [HIN04]. Oldroyd [LZ05, MT08]. One [ABV06, Bri03, CGZ09, CS06, DFLM09, DG09, FM04, HIN04, HK00a, JNS06, KK07, KSS02, LP02, Lim02, Mar06a, MV08, Nak03, NS09, PLN02, Ryb01, Sac04, SZ09, Tzi06]. One-Dimensional [ABV06, CGZ09, CS06, DFLM09, DG09, HK00a, JNS06, KSS02, Lim02, MV08, Nak03, NS09, PLN02, Sac04, SZ09]. One-Directional [Bri03]. Ono [MST01]. Onset [Gui07]. Oort [BL07, LLW03]. Operator [DPR09, EL02, JN04, JLP06, KY03a, KM01, MW09, Pro01]. Operators [Amo01, BO07, CGNT08, EL02, HN05, Kov08, KP05, Mac04, O’D08, PR01, Rob04, San05, Ste04]. Optics [AR03, BL03, CGL05, Dum03, Joc00, LP00]. Optimal [BX09, BCP06, BD06b, BS05, CDJ08, CC07, CP00, Fig07, Ish05, Lip01, Lor00, MP03, Ped00, Ron05, SP07]. Optimally [GL07]. Optimization [The04]. Optimize [HKK01]. Options [BX09, CC07, YYW08]. Orbital [LW08]. Orbits [CL04]. Order [AFM01, AK03, Amo01, AGGM05, CP07a, DFLM09, FM00, GB04, GP06, GW09, HS08, HR06, JP00, JM09, KMS07, Kei01, KN09, KS07, Li06, LY05, MJZ04, Mos05, SW09, Wil09, vdBHV01, GJT06]. Order-Disorder [CP07a]. Ordering [NOV09]. Orders [Yoo01]. Orientation [CGLL02]. Orthogonal [QW00, Van01, Van85]. Orthonormal [Don00, PS00, ZD00]. Oscillating [ABDG04, GV05]. Oscillations [CL00, CMS04, MOPMW06, Moo02]. Oscillator [BLM09, Joc00]. Oscillators [QT05]. Oscillatory [WX05]. Oseen [Deu09, LS01]. Osher [Ish05]. Ostrovsky [CMJ09, LL06]. Ostwald [NP00]. Overcrowding [BDDS06]. Oversampling [Joh03].

Packed [BN02]. Packet [DMP06]. Packets [Che00]. Packings [BLRW06]. Pair [GH05]. Pairs [LYY05]. Pan [AH07]. Parabolic [BS01, BK04, BBF06, BK00, BFK03, CCG05, CD01, CJ04, CA00, Dal07, DGP04, FS03, FS05, GJV03, GWZ09, GJT06, GK05, HR08, JP00, JM09, KK08a, KK04a, KA00, LRR06, Lor00, MG05, PP00b, Tak00, vBPW08, vdBHV01]. Parabolic-Elliptic [MG05]. Parabolic-Hyperbolic [GWZ09, KA00]. Parallel [Zha03a]. Parameter [Pan03, Pan06]. Parameters [BK02, FL00, Li06]. Parametric [Coo00]. Parametrically [KW01]. Parametrization [RTW01]. Part [MB08]. Partial [AMP05, CRTW05, DdlL00, DGG00, GG07, HKP07, KKSY02, KW01, LR09, WCD07]. Particle [CP00, FGQ04, NS09, Rou06]. Particles [CCD02]. Partitions [HNZ06, Rig06]. Passage [CCCS06]. Past [GH09]. Path [Red02]. Path-Tracking [Red02]. Patlak [KS09a]. Pattern [CWY09, DS09, HW03]. Patterns [BKL^{+}09, Lew01, Osh04, Pro02, QW09]. PDE [CFTV03, DE09]. PDEs [BB05, Bla06, DZ08, FS09, GB04, IK01, KMS07]. Peclet [MDvD06]. Penalization [SR09]. Pencil [JLP06]. Pendulum [LLYZ03]. Perfect [ST08]. Perforated [WCD07]. Period [BD06a]. Periodic [AP06, BS01, BPBW05, BD06a, Bri03, BST05, BLM09, CDD06, CDG08, Coo00, CL04, DDT04, DL00, DPR09, EM09, FG09, GH08, JT09, JZ09b, Joh09, KMS07, KM01, LLYZ03, LY05, Lou06, MOPMW06, Moo02, PN08, Pin00, QT05, Sun01, Tor04, Tov08, Tzo06, Wah06, Wal09b, Yin04, HS02]. Periodic/Nonautonomous [HS02]. Periodicity [CO05, GJS09]. Permanence [GH03, LRRBS09, Sch02]. Permanent [Bos03, EL07, JS02]. Perpendicular [Pan03]. Persistence [GZ09, HS02, Nob09, Xu04]. Persistent [MZ05a]. Perturbation [DPR09, FM00, Gom03, KS01, LHK07, MW09, MH01]. Perturbations [Sch00]. Perturbed [Bos00, DRY07, GH09, Mor03, Sti00, Tov08, Tzo06]. Phase [DP05, FM00, FS03, FS05, GM05, Gui07, HIN04, Hat00, Hat03, HK00a, KSS02, LS00, Lip01, MPP09, Mos05, PV05, Sch08, SZ09, Tri08, Vis09, Wan08, dPKW07]. Phase-Field [GM05, KSS02]. Phase-Transition [PV05]. Phenomena [DH08a, DH08b]. Phenomenon [BLRW06, DS09, Tzi06]. Physical [CGL08, CF06b]. Physics [BK00, Bla06, LS00]. Phytoplankton [DH08a, DH08b]. Piecewise [DP03, SZ09]. Piercing [PP00a]. Piezoelectric [MN05]. Pitaevskii [Mar06a, KKSW08]. Placement [HKK01]. Planar [AK06, AKS06, BCV06, HM09b, JS02, LYY05, MV05, NN09]. Planck [EL07, Kag01]. Plane [Bjo08, Coz09, EK08, Gor02, GSZ04, KTvW05, Kel06, Lin03, MNR05, MR08, Ron05, Tor04, VA03, WX05]. Plane-Like [Tor04]. Planning [BS05]. Plasma [BK00, LS00]. Plasmid [WNW07]. Plasmid-Bearing [WNW07]. Plasmid-Free [WNW07]. Plasticity [GL08, Mie04, NK08]. Plate [BGM08, Gra08, Kim05]. Plates [BJ09, PS09a]. PML [HSZ03b]. Pockets [GZ00]. Poincaré [BSB03]. Point [FM04, Fou06, HM09b, LM09, Pro01]. Points [KS01, Sti00, FKW06]. Pointwise [AGSS00, TT00, YZ09]. Poisson [Kag01, Li11, AMP05, AL06, ABR00, Alì03, BNR05, Bos03, Bos05, CH06, EL07, Hwa04, LMR08, Li09, MV07, MT07a, NT08, Zha03b]. Pol [BLM09]. Polarizational [LP00]. Pole [HSZ03a, HSZ03b]. Poles [KN00]. Polycrystals [BBN08]. Polydisperse [BBF06]. Polygon [KK04b]. Polyharmonic [BCR01]. Polyhedral [Han01]. Polymer [NS04]. Polymeric [HZ09, Sch09, ZZ08]. Polynomial [BX05, LYY07]. Polynomials [AL06, BG04, DP03, Dun01, PS00, QW00, Van01, Van85]. Polytropic [LMR08]. Population [CJ04, KK08a, KGB06, OW07, Wal09a]. Populations [MOPMW06]. Pore [MDvD06]. Porous [AS07, BDV03, FS03, FS05, GG00, MP01, Mei08, NN01, Sch05]. Posed [MS04]. Posedness [AH08, Amb03, ACFS09, CMZ08, CKS^{+}01, CKS^{+}02, El 07, GH08, HL06, LP09, LCS03, MST01, PRS08, aLW09]. Positive [BY04, LM08, Wal09a]. Positivity [PS09a, PN08]. Posteriori [Laf04]. Potato [CF01]. Potential [AMSW05, Car03, Coo00, HM09b, LY09, Sac04, SZ09, YZ05]. Potentials [AK03, GZ00, VZ09, Zwo01]. Power [BN08, CR06]. Power-Law [BN08, CR06]. pp [HKS07, Pla09b]. Practical [Wil09]. Predator [AM05, DD02a]. Predator-Prey [DD02a]. Prescale [DGH00]. Prescribed [Rig06]. Presence [Alm08, AK03, BF06]. Preserving [PS09a]. Pressure [BMR09, BFK03, CL03, Sch05]. Pressureless [Bou00, GSR05, NT08]. Prevention [BDDS06]. Prey [AM05, DD02a]. Prey-Predator [AM05]. Price [CdMGGK09]. Pricing [YYW08]. Primitive [AG01b]. Principal [CG00]. Principle [BCP06, CN01, Rei01, Ste08]. Principles [BN08, EL02, KY04, KS06b].

Probing [UW07]. Problem [Abd03, AFM01, AHT03, AGGM05, Bel03, BV08, BIY08, BCN05, BBW09, BFG04, BCV06, BV00a, BW04, Bos05, BFK03, CR06, Cas03, CF00, CD01, CF03, CCCS06, CG08, CL04, CA00, CE07, Cui08, DG09, DH08a, DH08b, FH01, FP02, FH07, Gor02, GP06, HH01, HHR09, HHR11, HM09a, Hat00, Hat03, HMS03, HMS04, HL02a, HL02b, IWY08, KK08a, KKS00, KN00, KN09, LP08, LYY05, Li05, LP02, MV07, MW09, MS04, Mik02, MM05, MR02, Mos05, NP00, Osh07, PP00a, PP04, PP03, PP00b, Piv00, PS08b, QW09, RW03, RW05, RW08, Ren06, Ryb01, Vis07, WC09, Xu03, Zag05, ZY00]. Problems [AS06, BT07, BBF06, BJ09, Bri03, BV02, BST05, CKK06, CM02, CP07b, CR03, CEGMM09, DK08, Du00, DZ08, Ebm00, FS05, GG00, Gui07, GJS09, HT09, HSZ03a, HSZ03b, IK01, IL05, JT09, LK01, Lor00, Mar06b, Mor03, MR08, NS04, PS09b, Pin00, PRS08, RW00a, SR09, Tra06]. Procedure [Ort06]. Procesi [WY08]. Process [DFG02, GK08, IM01]. Processes [NRJ07]. Processing [GB04]. Profile [DH08b, WXM06]. Profiles [FM01, MP01, MZ05b]. Projections [BX05]. Propagation [Grü03, KL08, LLLY04, Mei08]. Properties [AG02, DN09, DGH00, GZ09, Li06, PN08, Var08]. Property [BPR00b, PS09a]. Pseudodifferential [Sch00]. Pseudoparticle [CP00]. Pulse [CMJ09, DKP07, MJZ04, RL06]. Pulses [AR03, BL03, CL09]. Purely [LR09]. Put [CC07]. Pyramid [DY00]. Pyramidal [Tan07].

q [BCN02]. Quadratic [SZ09]. Quadrilateral [LS02]. Qualitative [Wan00]. Quantization [JWW07]. Quantum [JP00, JM09]. Quasi [AH08, Bab08, BS01, BW08, BK04, BBF06, BST05, BFK03, Cha06, CGZ09, CDY04, DH08a, DH08b, DGG00, FS03, FS05, Gia05, GL08, Joc02, JLM01, Li08, LY05, LP02, MW09, MO09, PP00b, Pro02, Ren00, Ria04, Rob04, SS03b, WXM06, Wu05, Yoc07, vBPW08]. Quasi-Convex [DGG00, Bab08]. Quasi-Dual [CDY04]. Quasi-Geostrophic [AH08, SS03b, Wu05, Cha06]. Quasi-invariant [BW08]. Quasi-Linear [BST05, BFK03, JLM01, Li08, LP02, MW09, PP00b, Ren00, Yoc07, vBPW08, BS01, BK04, BBF06, DH08a, DH08b, FS03, FS05, Ria04, Rob04]. Quasi-neutral [WXM06]. Quasi-Periodic [LY05]. Quasi-Static [Gia05, MO09, GL08]. Quasi-Stationary [Joc02]. Quasi-Steady [Pro02]. Quasiconvex [CDD06]. Quasiconvexity [Ped00]. Quasilinear [HR08, MS06, PS06]. Quasiminimal [Rig06]. Quintic [Tzi06].

Racah [AL06]. Radial [BCR01, BGL07, DRY07, GG00, LYY07, NW04, Yoo01]. Radially [Nar08]. Radiation [KS00]. Radiosity [Han01]. Radius [AHM09, FS01]. Radon [AK06, BX05]. Raising [Kei01]. Raman [WZ05]. Random [BG04, BBN08, BLRW06, Dal09, DO09]. Range [AK06, BGL07, CO05, Nak01]. Rapid [CL00]. Rapidly [CT08, Dal09]. Rarefaction [JNS06, KMX08, Lew05a, Lew05b, Nak03, NYZ04]. Rarefactions [Bia01]. Ratchets [BDK09]. Rate [Bou00, BMR09, God09, GK08, Hat00, Ish05, KS09a, MP03, Mas01, NN09, Sun07, TTX03]. Rates [CNO06, DP05, JM08, Lay07, ORS06]. Reachable [BV08]. Reaction [AK00, Bos00, CCK07, CQ07, FGHS07, FC06, JZ09b, KS00, NRJ07, Osh04, SZ00, WW02, WZ05, YCL01]. Reaction-Diffusion [AK00, Bos00, CCK07, FGHS07, JZ09b, NRJ07, Osh04, SZ00, WW02, YCL01]. Reaction-Hyperbolic [FC06]. Reactive [MDvD06]. Reactor [The04]. Real [Auc09].

Realization [Bra05]. Receptor [MCP08]. Receptor-Based [MCP08]. Reconstruction [BX05, Mac04, Pal09, Ron05, Sun07]. Recovery [AIM03]. Rectangular [FMP09]. Reduction [Bab08, BGM08, Rei01]. Refinable [Han08, Pro00]. Refined [CKS^{+}02, KS09a]. Refinement [JJS00]. Regimes [Bos03]. Regular [CM01, Kov08, Kry03]. Regularity [Abd03, BV00b, CM02, Cha06, CRTW05, DGG00, Fig07, GH08, Hwa04, Kan04, KK04b, Le05, LW09, MN05, NK08, Oga03, SZ07, SS09b, Sug09, Mus05]. Regularization [All08, Cré03, FMP05, HK00a, LS03]. Regularizations [RW00b]. Regularized [CMJ09, DKP07]. Reinforced [BD09, Gra09]. Related [ABV06, CP07b, CM01, CR03, DGP04, Dun01, Hat03, IL05, KK02, KK05, Kur04, LP08, LYY05, Mar06b, MST01, Mos05]. Relation [LLW03]. Relations [Gra09, HL00]. Relaxation [ABR00, CMM06, Li08, MZ05b, MM05, TT00, Xu09]. Relaxation-Time [Xu09]. Remark [Zwo01]. Remarks [HK06, MST07]. Renormalization [BC06b, Pro02]. Renormalized [BK04, CL05, MG05]. Representation [AC04, Bab08, BCQ01, Can01, GL07, ZD00]. Reproducing [Auc09]. Repulsive [Car03, HM09b]. Residual [IWY08, NW03]. Resistance [CLR02, PP00a, Pla09b]. Resolution [MN00]. Resolvent [KN00]. Resonance [Coo00, GZ00, KKSY02, dPKW07]. Resonant [Bou08, HY03, MPS05, MPS06]. Respect [BF01, KK07, Kri04, Lor00, Sch00]. Response [Pot07]. Restoration [CRTW05, DFLM09]. Result [AC04, BV00b, CKS^{+}02, GH05, Grü03, Hor08, Ift02, Mas01]. Results [ACFS09, BCLM08, DGP04, EM09, MN05, PS09b, Wes02]. Retarded [LWS05]. Reverse [Dun01]. Reversible [BPBW05, Xu04]. Reynolds [Wil09]. Ridgelets [Can01, Don00]. Riemann [HL02a, HL02b, Pin00]. Riemannian [CMS04, KY03a]. Right [BG08, II01]. Right-Angle [II01]. Right-Hand [BG08]. Rigorous [CMM06, MDvD06, ORS06, PW08]. Rings [Kat05]. Ripening [NP00]. Rise [BBN08]. Risk [CCCS06]. Rivers [DE09]. Road [CGP05]. Robin [DK08]. Robust [GH03]. Rod [GZ09, HM09b, LL05, Vui03, CHK05]. Rodlike [ZZ08]. Role [TW09]. Roll [Nob09]. Rotating [AB06, CT08, Dal09, Lou06]. Rough [Are01, CWM05, MV07, Pla09a]. Rupture [CD08].

s [BCN02, BCN02, Du00]. S-Shaped [Du00]. Safronov [BL07, LLW03]. Saint [Nob09]. Samples [JS02]. Sampling [BG04, Mac04, Sun07]. Satisfying [BPBW05, MW09, MV05]. Saturated [Gor02]. Saxton [BC05, Len08, Yin04]. SBV [Bab08]. Scalar [CM02, IV04, KL02, LZ03, LTT03, MP03, NT08, Pan09, Roh05, Zag05]. Scale [BF01, BCD05, MT07b, Zui00]. Scaling [CM06, CGL04, FGQ04, GR00, MP05]. Scattered [NW02, NW04]. Scattered-Data [NW04]. Scattering [AN00, AIM03, AKS06, Are01, BV08, BBW09, CW00, CWM05, CWM08, HSZ03a, HSZ03b, KS05a, LW05, NS07, O’D08, Pla09a, Ryb01]. Schaeffer [Nik01]. Schauder [Lor00]. Scheme [II01]. Schemes [DLY07, KS09c, Zho01]. Scholes [Cré03]. Schrödinger [AMP05, AMSW05, AP06, Ban03, BNR05, BO07, Car03, CW00, CKS^{+}01, CKS^{+}02, CT09, DW05, DPR09, GH08, HNS08, HK06, KM01, KPS09, KKSW08, Kov08, Kur04, LM08, LY05, LP02, LW08, Mar08, Mar06a, MJZ04, Nak01, O’D08, Rou06, Sac04, SS09a, Spi03, Sze05, Tzi06, VZ09, VZ07, Zha03b]. Schrödinger-Type [O’D08]. Schrödinger/Gross [KKSW08]. Screw [Pon07]. Second [AFM01, FM00, HR06, KMS07, MH06, Rob04, SS03a]. Second-Gradient [MH06]. Second-Order [HR06]. Section [BV00b]. Sector [GGH05]. Seeding [BR06]. Segel [BDDS06, KS09a, Sug09]. Segregation [Ble08]. Seismic [Mei08]. Sekerka [Rög05]. Selection [LHK07, Ort06]. Selective [CRTW05]. Self [BL07, CW00, Dum03, EL02, GGH05, HM09b, LS03, SZ05, Zho01]. Self-Adjoint [EL02]. Self-Focusing [Dum03, SZ05]. Self-Potential [HM09b]. Self-Similar [BL07, CW00, LS03, Zho01, GGH05]. Selfdual [MNR05]. Sel’kov [Lie05]. Semiclassical [LW08, Zha03b]. Semiconductor [MT07a]. Semiconductors [ABR00, WXM06, Xu09]. Semicontinuity [Bab08]. Semidiscrete [BGHR03]. Semiflow [Kri04]. Semigeostrophic [CF06b, Loe06]. Semigroup [KL08]. Semilinear [AGGM05, CFTV03, Du00, DMPD09, FHW02, MR02]. Semirelativistic [CO06]. Semistrong [DKP07]. Semisubmerged [PP04]. Sensitivity [PRS08]. Separate [Zha03a]. Separated [NRJ07]. Separation [HIN04]. Sequences [Mor03, dVGH03]. Sequential [FN05, Pel01]. Set [CD08, FN05, Pan09, Vui03]. Sets [AG02, BV08, CR06, Gom03, IM04, Pla09a]. Setting [Mac04]. Shabat [KS03a]. Shah [CMS04]. Shallow [CMZ08, CT08]. Shallow-Water [CT08]. Shape [CM98, CM00, MPP09, Ste05, Yoc07]. Shape-Memory [MPP09]. Shaped [Du00, MS06]. Shapes [Tan07]. Sharp [BL06, BY04, CQ07, DO09, MM05, Oga03]. Sharp-Interface [DO09]. Shaw [CR06, EM09, GP05, GP06]. Shaw-Type [GP05]. Shear [BMR09, FMS03, FS00, Gor02]. Shear-Dependent [FMS03]. Shear-Rate [BMR09]. Shearlet [KS09c]. Shearlets [GL07]. Sheets [Amb03, CZZ07]. Shelf [JLP06]. Shell [ABV06, BL01, CCS07]. Shell-Membrane [ABV06]. Shock [BGSZ01, BGHR03, CWY09, GH09, HMS04, LLLY04, Lew01, LK01, MZ05b, SS08, Tra06]. Shocks [BV05, Bia01, CKK06, CL03, Cou02, CS06, HL00, LY09, Rou03, SP07, Yua06]. Short [AR03, BL03, CL09, CMJ09, HS06, Nak01]. Short-Range [Nak01]. Shrinking [Váz03]. SIAM [HKS07, Pla09b]. Side [BG08]. Sign [WXM06]. Sign-Changing [WXM06]. Signal [BNG04]. Signals [Sun07]. Signorini [Ren06]. Similar [BL07, CW00, LS03, Zho01, GGH05]. Simple [Piv00]. Simulated [CHK08]. Simulation [DE06]. Sine [Kat05]. Sine-Gordon [Kat05]. Single [CLR02, Pla09b, QXM08]. Single-Impact [CLR02, Pla09b]. Single-Wave-Form [QXM08]. Singular [Ban03, BDR09, DE09, Esp05, FGHS07, FM00, HK00a, KS01, LHK07, MW09, MH01, Osh07, PV05, PS06, Ria04, SR09, VZ09, ZY00]. Singularities [CF01, CFdlL04, Don00, Red02, Ren00]. Singularly [Bos00, DRY07, Mor03, Sti00, Tov08, Tzo06]. Sixth [JM09]. Sixth-Order [JM09]. Size [Gia05]. SLEP [NS04]. Slip [BMR09]. Slippage [Grü03]. Slyozov [Lau02, NP00]. Small [AK03, AIM03, Bou08, Fil07, HK00b, KPS09, KS09b, Pan06]. Smoluchowski [LLW03, MP05]. Smooth [ABR00, Alì03, BC06b, CT08, O’D08, Pro00]. Smoothing [CRTW05, KS07, Sze05]. Smoothness [FHW02]. Snakes [BKL^{+}09]. Sobolev [Bra05, Can01, KK04a, Oga03, Yoo01]. Solenoidal [LW09]. Solid [Bla06, BMR09, GV00, HMS03, HMS04].

Solidification [BLRW06]. Solitary [BD02, CGNT08, CMJ09, GW07, LL06, Miz01, Miz03, SW08]. Solitons [Kur04, Mar06b, MC01, SS09a]. Solution [ABDG04, Bos05, Bou00, Cas03, Cui08, Han01, LLYZ03, Mal00, MB08, MT09, Pro00, WY08]. Solutions [AK00, AS06, ABR00, Alì03, AS07, AGSS00, ANR09, AG02, BL07, BS00a, BS01, BK05, BPBW05, BCN05, BNP06, BK04, BR06, Bia01, BD06a, BW04, BC05, BD02, BLM09, CFTV03, CR06, CW00, CDJ08, CdMGGK09, CM01, CD01, CL03, CQW03, CWY09, CT08, CJT09, CHK05, CEGMM09, CDM08, CF06b, DP09, DRY07, DDGM07, DS06, DDT04, DLS05, DL00, DZ03, DZ08, Esp05, ET09, FN05, FM09, GIK05, GW06, GIK08, GV05, Gev08, GGH05, GG00, Gom03, Gra08, GB04, GJS09, Hat03, HZ09, HKP07, Hof06, HK00b, HW02, HKS05, HKS07, IL05, JJS00, Joc00, Joc02, Joh09, JP00, JM08, JLM01, KK02, KK05, KMX08, KLVZ09, KK04b, Le05, LZ05, LP08, Lew05a, LY05, Lil07, LHK07, LS03, Lip01, MT03, MV08, Mik09, Miz01, MJZ04, MG05]. Solutions [Moo02, Nak03, Nak01, NN01, Pan09, PN08, PP00b, Per06, Pis04, QXM08, RW00a, RW03, RW05, RW08, Rie03, Rög05, SW09, ST08, SZ07, SP07, SS09b, Sug09, Tak00, Tov08, Váz03, WW07, Wal09a, Wan00, WLW09, WW02, Wu05, WC09, YCL01, Yin04, vDPP07, Mus05]. Solvability [Abd03, Bla06, BJ09, KS07, PP00a, PP04]. Solvable [MC01]. Solvation [Li09, Li11]. Solvent [Li09, Li11]. Solving [HSZ03a, HSZ03b]. Some [ACFS09, BB05, CEGMM09, CA00, DGP04, DN09, DZ03, Lin03, Osh07, ST08, Sch00, Var08, Wes02, Du00]. Source [BT07, Ha01, HY03, KY03b, TW09]. Sources [Fou06, GMS09]. Space [ANR09, BS01, BGL07, CRTW05, CF06b, DMPD09, FM01, HIN04, Kan04, KK04a, Nak03, NN09, Nis02, RTW01, Yoo01]. Space-Time [BS01]. Spaces [Auc06, Auc09, Bra05, FHW02, Han08, HNZ06, Kry01, Kyr04, MR02, Wu05]. Sparse [CDD03, GL07]. Spatial [CL04, Deu09, FKK09, FLM01, GW07, JZ09b, LR09, PP00b]. Spatially [Wal09a]. Spatially-Structured [Wal09a]. SPDEs [Kry09]. Special [AG01a, BD06b]. Species [HLMP03]. Spectra [CGNT08, RW03]. Spectral [Auc06, BC06a, BBW09, CGL08, HS08, LL05, Yoo01]. Spectrum [DN09, VA03]. Spectrums [BL03]. Speed [CQ07, Grü03]. Speeds [HZ08, ST01]. Speer [JM08]. Spheres [BG08, BGL07, DW05, FPR04, NW02, NPW06]. Spherical [RW08, UW07]. Spherically [GJX08, WZF08]. Spike [CK01b]. Spline [AS00, GK00, HNZ06]. Splines [BCR01, KS03b]. Splitting [Che00, GK05]. Splitting-up [GK05]. Spohn [JM08]. Spread [FLM01]. Spreading [GW06, Grü03, HZ08]. Square [BD06b]. Squirt [CFdlL04]. Stability [Alm08, AG02, BS08, BGHR03, BK02, BFV08, Bia01, Bou08, BLM09, CCD02, CD01, CLGZ06, CZZ07, Cou02, CT09, Cui08, DG09, DGG08, DGH00, DKP07, EM09, FN05, GR00, GIK05, GIK08, Gom03, Ha01, HY03, HH01, HM09b, Joh09, KPS09, KA00, LL05, LWS05, LL06, Lew01, Lew05b, Li08, LK01, LS03, LW08, LTT03, Mar06b, MZ05b, Nis02, NYZ04, NS04, PN08, Pro02, Pro01, PS08b, QXM08, QW09, RW05, Ron05, Sch00, SW08, SZ00, STZ02, Tsa07, WW02, WC09, YZ09]. Stable [AD05, Cou02, DDGM07, KP07, LRRBS09, MJZ04, Osh04, Sin06]. Stage [JZ09b, KGB06]. Stage-Structured [KGB06]. Standard [Lóp00]. Standing [Bou08, CT09, OT07]. Star [AS00, MS06]. Star-Shaped [MS06]. Star-Supported [AS00]. Stars [DGG08]. State [Alm08, AK03, Bla06, God09, HKM08, LMR08, LP08, Lie05, LL00, aLW09]. States [CL03, DW05, FG09, KPS09, LW05, LM08, LW08, MZ05a, Rou06]. Static [CFTV03, Gia05, MO09, GL08]. Stationary [CP07a, Cui08, Dal09, GIK05, GIK08, GG07, Joc02, NN09, Osh04, Tak00, WC09]. Steady [AK03, CZZ07, FMS03, LW05, Lie05, LY09, MZ05a, Pro02, Wah06, Wal09b, Yua06, aLW09]. Steady-State [AK03, Lie05]. Stefan [MR08, PS08b, Vis07]. Stefan-Type [MR08, Vis07]. Stems [CKK06]. Step [GZ00]. Stewartson [VA03]. Sticky [NS09]. Stieltjes [FS09]. Stiff [KS06a]. Stochastic [BDR09, BDK09, CCK07, DMPD09, IM04, Kim00, Kim02, Kim05, LP08, LR09, Mik02, MR04, Mik09, SS09a, WCD07, Zha08]. Stokes [FN05, Kim13, AGSS00, ANR09, BDK09, Bra05, BKL04, CCS07, FGM05, FH07, FM01, HZ03, Hof06, Ift02, JNS06, JZ09a, Kel06, Kim06, Kim09, KK04b, LW09, LXY00, MV08, Mik02, MR04, Mik09, NYZ04, Per06, PRS08, QW09, SZ07, WX05, WZF08]. Stokes-Like [FGM05]. Stokesian [EM09]. Stolz [DE06]. Stone [IM01]. Stopping [GK00]. Stored [Pon07]. Strain [CG00, GL08]. Strains [BBN08, Mie04]. Stratified [Wal09b]. Stress [CM02, CDY04, Dal09, IWY08, MB08, NW03]. Stress-Enhanced [MB08]. Stretching [ET09]. Strichartz [Ban03]. Strictly [HL00]. String [BB08]. Strong [ANR09, CJ04, GSR05, Lew05a, Lew05b, MV08, Mik09, NYZ04, OT07, PS00, Rou03]. Strongly [AMSW05, BFK03]. Structure [HNZ06, HT09, JZ09b, LMR08, Pin00, Yin04]. Structured [KGB06, Wal09a]. Structures [AKS06, BF01]. Study [RW00b, Vui03]. Sturm [BBW09, Piv00]. Subcritical [PP04]. Subdivision [DLY07, KS09c, Pro01, Zho01, dVGH03]. Subject [HM09b]. Successor [QT05]. Sufficient [Pis04, SZ07]. Suns [DGG08]. Super [FM04]. Super-Brownian [FM04]. Superconducting [AC07, DD02b, Pan03]. Superconductivity [AH07, BCM00, LL03, SS03a]. Superconductors [Alm08, BK05, KS09b, LL00]. Supercritical [DF06]. Superlinear [AC04]. Supersonic [CZZ07, GH09, Mar08]. Supply [HKP07]. Support [HS06, KY03b, KS05a]. Supported [AS00, DGH00, HS08, NW02]. Supremal [BNG04]. Surface [AH07, Amb03, CS07, Dal09, GIK05, GIK08, GP05, GP06, KL08, PP00a, PS08b, Sin06]. Surface-Piercing [PP00a]. Surfaces [Are01, CWM05, FGHS07, KP07, Nar08, O’D08, Tor04]. Surfactant [GW06]. Suspension [Din03, Moo02]. Suspensions [BBP05, BBF06]. Switch [BK02]. Switches [GJS09]. Symbols [KP05]. Symmetric [CF06a, GJX08, HS08, Kan04, Lóp00, Nar08, PP00a, SZ07, WZF08]. Symmetrical [Tov08]. Symmetrized [ZD00]. Symmetrizers [Tra06]. Symmetry [DG09, EK08, ET09, FS00, GG00, JZ09a, KKSW08]. Symmetry-Breaking [KKSW08]. Synchronism [BS08]. Synchronization [CCK07, LAJ08]. System [AMP05, ACFS09, ANR09, BK05, BK00, Bla06, Bos03, CCD02, CCK07, CH06, CK01b, CF03, CLGZ06, CWY09, CG08, DP09, DRY07, DLS05, Din03, DE00, El 07, FN05, FGHS07, FGM05, FS03, FS05, GJV03, GW06, Hat03, Hwa04, Isa01, IWY08, Joc02, KK02, KK05, Kim00, KK04b, LM09, LMR08, LS03, Loe06, MP01, Mar06a, MT07a, Mas01, MG05, NW03, NS09, NN01, OT07, STZ02, TW09, Tsa07, UW07, WZ05, YCL01, Yoc07, Yua06]. Systems [AG02, BNR05, BS08, BK02, BBF06, Bia01, BSB03, CKK06, CQ07, DG00, DL00, DIN04, DS09, EL07, FGQ04, FL00, Gev08, Ha01, HY03, HKM08, HL00, HS02, HR08, HK00b, Jen00, JP00, JM09, KS02a, KS03a, Laf04, Le05, Lew05a, Li08, LM08, Lim02, MZ05a, MZ05b, MN05, MVNO08, NK08, NT08, Osh04, Osh07, Pro02, Ren00, Rou03, ST01, Sle08, Sug09, VA03, Wan00, WW02, Xu04, vBPW08].

t [BCN02]. Table [FG09]. Tails [vBPW08]. Tallest [CM98, CM00]. Tangency [HP09]. Tangential [KK08b]. Tapered [Vui03]. TE [SZ05]. TE-Modes [SZ05]. Technique [MNR05]. Techniques [KS02a, MCP08]. Telecommunication [DMP06]. Temperature [BMR09]. Templates [TY05]. Temple [AG02]. Tension [Amb03, PS08b]. Term [KY03b, Tzo06, ZY00]. Terms [AK03, GR03, Roh05]. Test [Abd03, Pot07]. Their [BFP08, KK08b, Lou06, PP00b, RW05, Tra06]. Theorem [BB05, Sug09, ZS02]. Theorems [KP07, MT03]. Theories [LZ03]. Theory [All08, BBW09, BP07, CW00, CMM06, DdlL00, Du00, El 07, Gom03, Gra09, HM09b, HSZ03a, Hur06, KK04a, KS01, MH01, NP00, Váz03]. Thermal [Xu03]. Thermally [MPP09]. Thermistor [Xu03]. Thermocontrol [GJS09]. Thermoelasticity [Isa01, Yoc07]. Thermomechanical [Ste05]. Thermoviscoplasticity [KSS02]. Thin [AKS06, BL06, BF06, BS00b, BF01, CCK07, CM06, CD08, DD02b, FMP09, GW06, Hor08]. Thin-Walled [FMP09]. Third [Amo01, GP06, KS07]. Third-Order [GP06]. Three [CN00, ET09, FS03, FS05, HT09, Kan04, RW03]. Three-Dimensional [ET09, HT09, RW03]. Three-Phase [FS03, FS05]. Threshold [WW02]. Thresholds [LT01]. Tight [NPW06]. Tikhonov [Cré03]. Tilings [Zho01]. Time [BS00a, BS01, BK00, BD06a, BDK09, Ble08, CGLL02, CWM08, CQW03, CT08, CR05, Coo00, DS06, Deu09, DS05, DS09, GWZ09, Hat03, HSZ03a, HSZ03b, Joc00, Kag01, Kri04, KS05b, Lim02, Lor00, LAJ08, MS06, Miz01, Nar08, NK08, Sac04, Xu09, YZ09, YZ05, Zui00, vBPW08, dMKST09]. Time- [Kri04]. Time-Dependent [Ble08, Deu09, GWZ09, Sac04]. Time-Frequency [KS05b]. Time-Harmonic [CWM08, HSZ03a, HSZ03b]. Time-Incremental [NK08]. Time-Periodic [Coo00]. Time-Scale [Zui00]. Time-Varying [LAJ08]. Times [GK00]. Tissues [WC09]. Tomography [Brü01, HHR09, HHR11, KKSY02, MN00]. Tori [BCV06, Xu04]. Torsional [Moo02]. Torus [CL00, MNR05]. Total [All08, CR03, Lau02]. Touching [BV00b]. Trace [Auc06, Ryb01]. Tracing [SP07]. Tracking [Laf04, Red02]. Traffic [BR03, CGP05, HR06]. Trajectories [LRRBS09]. Trajectory [Kry03]. Transform [AK06, Cla00, PN08, RL06, Pal09]. Transformation [BBN08, JN04]. Transformations [MPP09]. Transforms [DY00]. Transition [ABV06, CP07a, Hat00, Hat03, LZ03, Lim02, PV05, Tri08, dPKW07]. Transitions [DP05, FM00, HK00a, Mos05, SZ09, Vis09]. Transmission [NRJ07, PS08a]. Transonic [GH09, LY09, Yua06]. Transport [ACFS09, BT07, BC06b, BG08, CDJ08, CP00, Fig07, GP04, HLTT08, HKM08, HKS05, HKS07, aLW09]. Trapping [BDV03, Sch08]. Traveling [BCC03, BS00b, CFG06, CCP07, CN00, Din03, FC06, GB04, HZ08, Joh09, Li08, Mar08, Nis02, OW07, SW09, SZ00, Tan07, Tsa07, Var08, BC06a]. Travelling [CQ07, GR03, Kat05, Mar06a, PN08, vdBHV01]. Travelling-Wave [PN08]. Travelling-Waves [Mar06a]. Trees [Kov08]. Trick [Che00]. Trigonometric [BG04]. Triple [CWY09, EL02]. Triple-Shock [CWY09]. Truncated [KN09]. Truncation [FMS03]. Tumor [BFG04, CF03, Cui08, FH07]. Tumors [CE07, WC09]. Turbulence [Lay07]. Turbulent [DE06, MR04]. Twist [LLYZ03]. Two [Bel03, BIY08, BF01, BKL04, BCD05, CCK07, Cas03, CWY09, CY04, CN00, FLP05, Fil07, Fou06, GZ00, HP09, KS09a, Kim02, KS01, LS00, LZ05, LP09, Lip01, MT03, MT07b, Pin00, Pla09a, RW00a, Sch08, Sun01, Yua06, Zha03a]. Two-Dimensional [Bel03, BKL04, Cas03, CWY09, CY04, FLP05, Fil07, HP09, KS09a, Kim02, LZ05, LP09, MT03, Pin00, Pla09a, Yua06]. Two-Fluid [Sun01]. Two-Layer [CCK07]. Two-Phase [LS00, Lip01, Sch08]. Two-Scale [BF01, BCD05, MT07b]. Two-Step [GZ00]. Type [CN01, CO06, El 07, GP06, GJS09, HNS08, KP08, MR02, MST07, MT08, MR08, Nik01, O’D08, Oga03, Vis07, BB05, GP05].

u [BCN02]. Unbounded [CWM05, GMS09, JN04, Li06]. Uncertainty [BCP06]. Undercompressive [BS00b]. Underdamped [QXM08]. Undulating [Lou06]. Unfolding [CDD06, CDG08, Sti00]. Uniform [Dun01, HS02, JWW07, MP05, QW00, San05]. Uniformly [MZ05a]. Unilateral [BFG04, PS09b]. Unique [Bla06, NW03]. Uniqueness [AG02, Are01, BCLM08, BR06, BW04, Bos05, BC06b, CWM05, CdMGGK09, CO05, CFG06, CDM08, DK08, ET09, Fig07, GV00, Hof06, Ift02, Isa01, KY03b, KP07, LM09, LM08, LY09, MV08, MN05, Ren06, Mus05]. Univariate [DLY07]. Universal [DP05]. Unreachable [BV08]. unstable [EGW06]. Unsteady [BMR09, CGZ09, Gra08]. Unstirred [WNW07]. Up-to-the [BL01]. Upper [ORS06]. Upscaling [MDvD06]. Urban [BS05]. Using [CDD03, Cré03, GL07, MCP08, RW00b].

Vacuum [CL03, LXY00]. Validity [Fra07]. Value [Abd03, Bos05, Ebm00, FS05, Lee09, LP02, MS04, NP00, PP00b, PRS08]. Valued [MT03]. Values [FPR04]. Vanishing [AHM09, CL03, Coz09, FS00, Ift02, JNS06, MP03, MO09]. Vapor [Tri08]. Variable [CGLL02, CLGZ06, KK07, Mal00, Yua06]. Variable-Area [Yua06]. Variables [GR00]. Variation [All08, BV02, CR03, GM01, ZS02]. Variational [AFM01, BN08, CM02, CWM05, EL02, FN05, IK01, IL05, KS06b, Li06, MT07b, Pel01, RW00a, Ste08, YYW08]. Varying [DG00, LAJ08]. Vector [Lee09, LW09, Sch02]. Velocity [BG08, Coz09, Nis02, Wes02, FKW06]. Venant [Nob09]. Version [BL01, Loe06]. Vertical [Ift02]. Vertices [Han01]. Very [ANR09, BCN02]. via [CK01a, CDD06, Gra09, HK00a, KY03a, KS02a, LZ05, Lil07, Mei08, MS06, MT07b, MO09, NT08, QT05]. Vibrations [BB08]. Virus [FLM01]. Viscoelastic [CM01, Lim02, MT08]. Viscoelastodynamic [PS09b]. Viscoplasticity [Nes07]. Viscosity [AS06, BBP05, CR06, Coz09, FMS03, FS00, Gom03, GJX08, HW02, Ift02, JNS06, JLM01, LK01, LS01, MP03, Pis04, WX05, WZF08]. Viscous [BGSZ01, Bjo08, Bou00, CMZ08, DS05, FL00, GW06, Gra08, HW09, HMS03, HMS04, LLLY04, LK01, Nak03, NN09, PLN02, Rou03, SS08, SP07, TTX03]. Vision [AFM01]. Visual [CMS04]. Vlasov [Bos03, Bos05, CH06, Hwa04, Kag01, LMR08]. VMO [KK07, Kry09]. Volatility [Cré03]. Voltage [AK03]. Volterra [DGG08, LRRBS09, MT09]. Volume [Lau02, LW05]. Vortex [AB06, Amb03, CZZ07, DZ03, Fil07, LM09, LYY05, LL00, Spi03]. Vortex-Wave [LM09]. Vortices [AC07, LL03, MNR05]. Vorticity [BR06, Coz09, ET09, GW07, Hur06, LM09, Var08, Wah06]. Vries [CCP07, DDT04, Joh09, MC01, Miz01]. vs [BDDS06].

Wagner [Lau02, NP00]. Waiting [DS09]. Walled [FMP09]. Walls [CZZ07].

Wasserstein [CDJ08, DS08, NS09, OW05]. Water [BST05, CMZ08, CT08, CN00, GW07, Hur06, KS06b, Var08, Wal09b, Wri05]. Wave [Are01, BD06a, BD02, CKK06, CWM08, CQ07, CHK05, CCP07, Coo00, GH09, GB04, GZ09, HMS04, Joh09, KK08b, LM09, LL00, LL03, Mei08, MS06, Moo02, PP00a, PN08, QXM08, QW09, SW09, ST01, Tsa07, VZ09, YZ05, Zha08]. Wave-Number-Explicit [CWM08]. Wavefronts [OW07]. Waveguide [JLP06]. Wavelet [Che00, Cla00, GK00, Joh03, RTW01, Ste04, Zui00]. Wavelets [HS06, HS08, dVGH03]. Waves [AKS06, BCC03, BGSZ01, BGHR03, BS00b, Bou08, BY04, BST05, CGNT08, CFG06, CMJ09, CN00, CT09, Din03, FC06, Gor02, GW07, HZ08, HW03, Hur06, JNS06, JLP06, KMX08, Kat05, KT01, KS06b, LLLY04, LL06, Lew01, Lew05b, Li08, LK01, Lou06, Mar08, Mar06a, MC01, Miz01, Miz03, Nak03, NN09, Nis02, NYZ04, Nob09, OT07, SS08, SW08, SZ00, Sun01, Tra06, UW07, Var08, Wah06, Wal09b, Wri05, vdBHV01, BC06a]. Weak [AS07, ANR09, BCN02, BR06, Cas03, CHK05, Cou02, CS06, DZ03, ET09, FN05, GP04, Gra08, Grü03, Hof06, JLM01, MB08, Miz03, Pan09, Per06, Rög05, SR09]. Weaker [Xu04]. Weakly [Kov08, MT08, Sch02, SZ04, WY08]. Wear [KS02b]. Wearing [IM01]. Weight [QW00]. Weighted [CR03, Han08, Van85, Van01]. Weights [Kry01]. Well [AH08, Amb03, ACFS09, CMZ08, CKS^{+}01, CKS^{+}02, El 07, GH08, HKK01, HL06, LP09, LCS03, MS04, PRS08, Sac04, aLW09]. Well-Posed [MS04]. Well-Posedness [AH08, Amb03, ACFS09, CMZ08, CKS^{+}01, CKS^{+}02, El 07, GH08, HL06, LP09, LCS03, PRS08, aLW09]. Wells [Zha03a]. White [DDT04, JWW07]. Whitham [KPT09]. Whole [FM01]. Wigner [Mac04, Zha03b]. Willmore [DG09]. Wilson [KS05b].

Windows [BD06b]. Without

[AS06, FL00, Joc05, STZ02]. Wound

[CF00]. Wriggled [RW05].

X. [AH07].

Young [LHK07, CP07a, Ped04, Ped06, Rie03, Zha03a]. Young-Measure [LHK07].

Zabolotskaya [KP08]. Zakharov [KS03a, LP09, Mas01, OT07]. Zeldovich [LZ03]. Zero [BPBW05, KL02, LS01, Van01, WX05, Van85]. Zero-Hamiltonian [BPBW05]. Zero-Viscosity [WX05].

References

Aftalion:2006:VLR

[AB06] Amandine Aftalion and Xavier Blanc. Vortex lattices in rotating Bose–Einstein condensates. SIAM Journal on Mathematical Analysis, 38(3):874– 893, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abdulla:2003:FBV

[Abd03] Ugur G. Abdulla. First boundary value problem for the diffusion equation I. Iterated logarithm test for the boundary regularity and solvability. SIAM Journal on Mathematical Analysis, 34(6):1422–1434, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                                     //epubs.siam.org/sam-bin/
                                                     dbq/article/41504.

Amirat:2004:AAS

[ABDG04] Y. Amirat, O. Bodart, U. De Maio, and A. Gaudiello. Asymptotic approximation of the solution of the Laplace equation in a domain with highly oscillating boundary. SIAM Journal on Mathematical Analysis, 35(6):1598–1616, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41487.

Ali:2000:GER

[ABR00] G. Alì, D. Bini, and S. Rionero. Global existence and relaxation limit for smooth solutions to the Euler–Poisson model for semiconductors. SIAM Journal on Mathematical Analysis, 32(3):572–587, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35517.

Ansini:2006:MAC

[ABV06] Nadia Ansini, Andrea Braides, and Vanda Valente. Multiscale analysis by Γ-convergence of a one-dimensional nonlocal functional related to a shell-membrane transition. SIAM Journal on Mathematical Analysis, 38(3):944–976, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alicandro:2004:GIR

[AC04] Roberto Alicandro and Marco Cicalese. A general integral representation result for continuum limits of discrete energies with superlinear growth. SIAM Journal on Mathematical Analysis, 36(1):1–37, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42647.

Antontsev:2007:FSV

[AC07] S. N. Antontsev and N. V. Chemetov. Flux of superconducting vortices through a domain. SIAM Journal on Mathematical Analysis, 39(1):263– 280, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amadori:2008:MMF

[AC08] Debora Amadori and Andrea Corli. On a model of multiphase flow. SIAM Journal on Mathematical Analysis, 40 (1):134–166, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ambrosio:2009:SNW

[ACFS09] Luigi Ambrosio, Gianluca Crippa, Alessio Figalli, and Laura V. Spinolo. Some new well-posedness results for continuity and transport equations, and applications to the chromatography system.

SIAM Journal on Mathematical Analysis, 41(5):1890–1920, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alessandrini:2005:SDI

[AD05] G. Alessandrini and M. Di Cristo. Stable determination of an inclusion by boundary measurements. SIAM Journal on Mathematical Analysis, 37 (1):200–217, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44191.

Ambrosio:2001:VAS

[AFM01] Luigi Ambrosio, Loris Faina, and Riccardo March. Variational approximation of a second order free discontinuity problem in computer vision. SIAM Journal on Mathematical Analysis, 32(6):1171–1197, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36832.

Alessandrini:2001:DCS

[AG01a] Giovanni Alessandrini and Romina Gaburro. Determining conductivity with special anisotropy by boundary measurements. SIAM Journal on Mathematical Analysis, 33 (1):153–171, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36956.

Azerad:2001:MJH

[AG01b] Pascal Azérad and Francisco Guillén. Mathematical justification of the hydrostatic approximation in the primitive equations of geophysical fluid dynamics. SIAM Journal on Mathematical Analysis, 33(4):847–859, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37596.

Ancona:2002:USS

[AG02] Fabio Ancona and Paola Goatin. Uniqueness and stability of L^{∞} solutions for Temple class systems with boundary and properties of the attainable sets. SIAM Journal on Mathematical Analysis, 34(1):28–63, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38342.

Arioli:2005:SFO

[AGGM05] Gianni Arioli, Filippo Gazzola, Hans-Christoph Grunau, and Enzo Mitidieri. A semilinear fourth order elliptic problem with exponential nonlinearity. SIAM Journal on Mathematical Analysis, 36(4):1226–1258, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41853.

Amrouche:2000:PDS

[AGSS00] Cherif Amrouche, Vivette Girault, Maria Elena Schonbek, and Tomas P. Schonbek. Pointwise decay of solutions and of higher derivatives to Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 31(4):740– 753, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34617.

Almog:2007:DSS

[AH07] Yaniv Almog and Bernard Helffer. The distribution of surface superconductivity along the boundary: On a conjecture of X. B. Pan. SIAM Journal on Mathematical Analysis, 38(6):1715–1732, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abidi:2008:GWP

[AH08] Hammadi Abidi and Taoufik Hmidi. On the global well-posedness of the critical quasi-geostrophic equation. SIAM Journal on Mathematical Analysis, 40(1):167–185, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Angenent:2009:RVB

[AHM09] S. B. Angenent, J. Hulshof, and H. Matano. The radius of vanishing bubbles in equivariant harmonic map flow from D^{2} to S^{2}. SIAM Journal on Mathematical Analysis, 41(3): 1121–1137, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Angenent:2003:MFM

[AHT03] Sigurd Angenent, Steven Haker, and Allen Tannenbaum. Minimizing flows for the Monge–Kantorovich problem. SIAM Journal on Mathematical Analysis, 35(1):61– 97, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41092.

Ammari:2003:RSI

[AIM03] Habib Ammari, Ekaterina Iakovleva, and Shari Moskow. Recovery of small inhomogeneities from the scattering amplitude at a fixed frequency. SIAM Journal on Mathematical Analysis, 34(4):882–900, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39278.

Abdulla:2000:IDL

[AK00] Ugur G. Abdulla and John R. King. Interface development and local solutions to reaction-diffusion equations. SIAM Journal on Mathematical Analysis, 32(2):235–260, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32986.

Ammari:2003:HOT

[AK03] Habib Ammari and Hyeonbae Kang. High-order terms in the asymptotic expansions of the steady-state voltage potentials in the presence of conductivity inhomogeneities of small diameter. SIAM Journal on Mathematical Analysis, 34(5):1152–1166, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39923.

Ambartsoumian:2006:RDP

[AK06] Gaik Ambartsoumian and Peter Kuchment. A range description for the planar circular Radon transform. SIAM Journal on Mathematical Analysis, 38(2):681–692, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:2006:SEW

[AKS06] Habib Ammari, Hyeonbae Kang, and Fadil Santosa. Scattering of electromagnetic waves by thin dielectric planar structures. SIAM Journal on Mathematical Analysis, 38(4):1329– 1342, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Albanese:2006:PKE

[AL06] Claudio Albanese and Stephan Lawi. Poisson kernels as expansions in q-Racah polynomials. SIAM Journal on Mathematical Analysis, 38(3):977– 984, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alazard:2006:LMN

[Ala06] Thomas Alazard. Low Mach number flows and combustion. SIAM Journal on Mathematical Analysis, 38(4):1186– 1213, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ali:2003:GES

[Alì03]  G. Alì.  Global existence

of smooth solutions of the N-dimensional Euler–Poisson model. SIAM Journal on Mathematical Analysis, 35(2): 389–422, 2003. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        39322.

Allard:2008:TVR

[All08] William K. Allard. Total variation regularization for image denoising, I. geometric theory. SIAM Journal on Mathematical Analysis, 39(4):1150–1190, ???? 2008. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Almog:2008:SNS

[Alm08] Y. Almog. The stability of the normal state of superconductors in the presence of electric currents. SIAM Journal on Mathematical Analysis, 40 (2):824–850, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

anowski:2009:WPC

[aLW09] Piotr Krzy anowski, Philippe Laurençot, and Dariusz Wrzosek. Well-posedness and convergence to the steady state for a model of morphogen transport. SIAM Journal on Mathematical Analysis, 40(5):1725– 1749, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aiki:2005:PPM

[AM05] Toyohiko Aiki and Emil Minchev. A prey-predator model with hysteresis effect. SIAM Journal on Mathematical Analysis, 36(6):2020–2032, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44018.

Ambrose:2003:WPV

[Amb03] David M. Ambrose. Wellposedness of vortex sheets with surface tension. SIAM Journal on Mathematical Analysis, 35(1):211–244, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40386.

Amour:2001:IFT

[Amo01] L. Amour. Isospectral flows of third order operators. SIAM Journal on Mathematical Analysis, 32(6):1375–1389, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34724.

Abdallah:2005:AAS

[AMP05] Naoufel Ben Abdallah, Florian Méhats, and Olivier Pinaud. Adiabatic approximation of the Schrödinger– Poisson system with a partial confinement. SIAM Journal on Mathematical Analysis, 36(3):986–1013, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43791.

Alessandrini:2002:DIE

[AMR02] Giovanni Alessandrini, Antonino Morassi, and Edi Rosset. Detecting an inclusion in an elastic body by boundary measurements. SIAM Journal on Mathematical Analysis, 33(6):1247–1268, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38894.

Andreu:2009:NLE

[AMRT09] F. Andreu, J. M. Mazón, J. D. Rossi, and J. Toledo. A nonlocal p-Laplacian evolution equation with nonhomogeneous Dirichlet boundary conditions. SIAM Journal on Mathematical Analysis, 40(5): 1815–1851, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abdallah:2005:NSE

[AMSW05] Naoufel Ben Abdallah, Florian Méhats, Christian Schmeiser, and Rada M. Weishäupl. The nonlinear Schrödinger equation with a strongly anisotropic harmonic potential. SIAM Journal on Mathematical Analysis, 37(1):189– 199, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61455.

Ammari:2000:LFE

[AN00] H. Ammari and J.-C. Nédélec.

Low-frequency electromagnetic scattering. SIAM Journal on Mathematical Analysis, 31(4):836–861, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34360.

Amrouche:2009:SVW

[ANR09] Chérif Amrouche, Šárka Nečasová, and Yves Raudin. From strong to very weak solutions to the Stokes system with Navier boundary conditions in the half-space. SIAM Journal on Mathematical Analysis, 41(5): 1792–1815, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Allaire:2006:LSE

[AP06] Grégoire Allaire and Mariapia Palombaro. Localization for the Schrödinger equation in a locally periodic medium. SIAM Journal on Mathematical Analysis, 38(1):127–142, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alterman:2003:DNG

[AR03] Deborah Alterman and Jeffrey Rauch. Diffractive nonlinear geometric optics for short pulses. SIAM Journal on Mathematical Analysis, 34(6):1477–1502, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40358.

Arens:2001:UEW

[Are01] T. Arens. Uniqueness for elastic wave scattering by rough surfaces. SIAM Journal on Mathematical Analysis, 33(2):

461–476, 2001. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35947.

Alfeld:2000:NSS

[AS00] Peter Alfeld and Larry L.

Schumaker. Nonexistence of star-supported spline bases. SIAM Journal on Mathematical Analysis, 31(2):455–465, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34234.

Ahn:2006:ESC

[AS06] Jeongho Ahn and David E.

Stewart. Existence of solutions for a class of impact problems without viscosity. SIAM Journal on Mathematical Analysis, 38(1):37–63, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amirat:2007:GWS

[AS07] Y. Amirat and V. Shelukhin.

Global weak solutions to equations of compressible miscible flows in porous media. SIAM Journal on Mathematical Analysis, 38(6):1825–1846, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Auchmuty:2006:SCT

[Auc06] Giles Auchmuty. Spectral characterization of the trace spaces H^{s}(∂Ω). SIAM Journal on Mathematical Analysis, 38(3): 894–905, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Auchmuty:2009:RKH

[Auc09] Giles Auchmuty. Reproducing kernels for Hilbert spaces of real harmonic functions. SIAM Journal on Mathematical Analysis, 41(5):1994–2009, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Babadjian:2008:LSQ

[Bab08] Jean-François Babadjian. Lower semicontinuity of quasi-convex bulk energies in SBV and integral representation in dimension reduction. SIAM Journal on Mathematical Analysis, 39 (6):1921–1950, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Banica:2003:DSI

[Ban03]  Valeria Banica.  Disper-

sion and Strichartz inequalities for Schrödinger equations with singular coefficients. SIAM Journal on Mathematical Analysis, 35(4):868–883, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41502.

Bambusi:2005:BLT

[BB05] Dario Bambusi and Massimiliano Berti. A Birkhoff–Lewis-Type theorem for some Hamiltonian PDEs. SIAM Journal on Mathematical Analysis, 37(1):83–102, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43610.

Baldi:2008:FVN

[BB08] Pietro Baldi and Massimiliano Berti. Forced vibrations of a nonhomogeneous string. SIAM Journal on Mathematical Analysis, 40(1):382–412, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berres:2006:NPQ

[BBF06] Stefan Berres, Raimund Bürger, and Hermano Frid. Neumann problems for quasilinear parabolic systems modeling polydisperse suspensions. SIAM Journal on Mathematical Analysis, 38(2):557–573, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berlyand:2008:RCT

[BBN08] Leonid Berlyand, Oscar Bruno, and Alexei Novikov. Rise of correlations of transformation strains in random polycrystals. SIAM Journal on Mathematical Analysis, 40(4):1550–1584, ???? 2008. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Berlyand:2005:NAE

[BBP05] Leonid Berlyand, Liliana Borcea, and Alexander Panchenko. Network approximation for effective viscosity of concentrated suspensions with complex geometry. SIAM Journal on Mathematical Analysis, 36(5):1580–1628, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42470.

Bennewitz:2009:ISS

[BBW09] C. Bennewitz, B. M. Brown, and R. Weikard. Inverse spectral and scattering theory for the half-line leftdefinite Sturm–Liouville problem. SIAM Journal on Mathematical Analysis, 40(5):2105– 2131, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressan:2005:GSH

[BC05] Alberto Bressan and Adrian Constantin. Global solutions of the Hunter–Saxton equation. SIAM Journal on Mathematical Analysis, 37(3):996–1026, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/62303.

Bates:2006:SAT

[BC06a] Peter W. Bates and Fengxin Chen. Spectral analysis of traveling waves for nonlocal evolution equations. SIAM Journal on Mathematical Analysis, 38(1):116–126, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bouchut:2006:URS

[BC06b] François Bouchut and Gianluca Crippa. Uniqueness, renormalization, and smooth approximations for linear transport equations. SIAM Journal on Mathematical Analysis, 38(4):1316–1328, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bates:2003:TWB

[BCC03] Peter W. Bates, Xinfu Chen, and Adam J. J. Chmaj. Traveling waves of bistable dynamics on a lattice. SIAM Journal on Mathematical Analysis, 35(2):520–546, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37400.

Briane:2005:LCT

[BCD05] Marc Briane and Juan Casado-Díaz. Lack of compactness in two-scale convergence. SIAM Journal on Mathematical Analysis, 37(2):343–346, 2005. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/62135.

Barles:2008:GER

[BCLM08] Guy Barles, Pierre Cardaliaguet, Olivier Ley, and Régis Monneau. Global existence results and uniqueness for dislocation equations. SIAM Journal on Mathematical Analysis, 40(1): 44–69, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonnet:2000:CMM

[BCM00] A. Bonnet, S. J. Chapman, and R. Monneau. Convergence of Meissner minimizers of the Ginzburg–Landau energy of superconductivity as κ → +∞. SIAM Journal on Mathematical Analysis, 31(6):1374–1395, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34616.

Bellout:2002:CVW

[BCN02] Hamid Bellout, Emil Cornea, and Jindrich Necas. On the concept of very weak L^{2} S o l u t i o n s t o Euler’s E q u a t i o n s. SIAM Journal on Mathematical Analysis, 33(5):995–1006, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38429.

Bellettini:2005:ESE

[BCN05] Giovanni Bellettini, Vicent Caselles, and Matteo Novaga. Explicit solutions of the eigenvalue problem div (\frac{Du|}{D}u|) =

u in R^{2}. SIAM Journal on Mathematical Analysis, 36(4): 1095–1129, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43000.

Benedetto:2006:OEB

[BCP06] John J. Benedetto, Wojciech Czaja, and Alexander M. Powell. An optimal example for the Balian–Low uncertainty principle. SIAM Journal on Mathematical Analysis, 38(1):333– 345, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buckdahn:2001:RFM

[BCQ01] R. Buckdahn, P. Cardaliaguet, and M. Quincampoix. A representation formula for the mean curvature motion. SIAM Journal on Mathematical Analysis, 33(4):827–846, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38033.

Beatson:2001:FER

[BCR01] R. K. Beatson, J. B. Cherrie, and D. L. Ragozin. Fast evaluation of radial basis functions: Methods for four-dimensional polyharmonic splines. SIAM Journal on Mathematical Analysis, 32(6):1272–1310, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36176.

Biasco:2006:DEI

[BCV06] Luca Biasco, Luigi Chierchia, and Enrico Valdinoci. Ndimensional elliptic invariant tori for the planar (N +1)-body problem. SIAM Journal on Mathematical Analysis, 37(5): 1560–1588, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bridges:2002:LIS

[BD02] Thomas J. Bridges and Gianne Derks. Linear instability of solitary wave solutions of the Kawahara equation and its generalizations. SIAM Journal on Mathematical Analysis, 33(6):1356–1378, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36149.

Biasco:2006:TPS

[BD06a] Luca Biasco and Laura Di Gregorio. Time periodic solutions for the nonlinear wave equation with long minimal period. SIAM Journal on Mathematical Analysis, 38(4):1090– 1125, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burchard:2006:GOW

[BD06b] Almut Burchard and Jochen Denzler. On the geometry of optimal windows, with special focus on the square. SIAM Journal on Mathematical Analysis, 37(6):1800–1827, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barchiesi:2009:HFR

[BD09] Marco Barchiesi and Gianni Dal Maso. Homogenization of fiber reinforced brittle materials: The extremal cases. SIAM Journal on Mathematical Analysis, 41(5):1874–1889, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burger:2006:KSM

[BDDS06] Martin Burger, Marco Di Francesco, and Yasmin Dolak-Struss. The Keller–Segel model for chemotaxis with prevention of overcrowding: Linear vs. nonlinear diffusion. SIAM Journal on Mathematical Analysis, 38(4):1288–1315, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blanchet:2009:SSD

[BDK09] Adrien Blanchet, Jean Dolbeault, and Micha Kowalczyk. Stochastic Stokes’ drift, homogenized functional inequalities, and large time behavior of Brownian ratchets. SIAM Journal on Mathematical Analysis, 41(1):46–76, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barbu:2009:SND

[BDR09] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Stochastic nonlinear diffusion equations with singular diffusivity. SIAM Journal on Mathematical Analysis, 41(3):1106– 1120, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertsch:2003:AOT

[BDV03] M. Bertsch, R. Dal Passo, and C. J. Van Duijn. Analysis of oil trapping in porous media flow. SIAM Journal on Mathematical Analysis, 35(1):245– 267, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40737.

Belishev:2003:CPT

[Bel03] M. I. Belishev. The Calderon problem for two-dimensional manifolds by the BC-method. SIAM Journal on Mathematical Analysis, 35(1):172–182, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41391.

Bouchitte:2001:HTS

[BF01] Guy Bouchitté and Ilaria Fragalà. Homogenization of thin structures by two-scale method with respect to measures. SIAM Journal on Mathematical Analysis, 32(6):1198–1226, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37026.

Beretta:2006:AFD

[BF06] Elena Beretta and Elisa Francini. An asymptotic formula for the displacement field in the presence of thin elastic inhomogeneities. SIAM Journal on Mathematical Analysis, 38(4):1249–1261, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertuzzi:2004:FBP

[BFG04] Alessandro Bertuzzi, Antonio Fasano, and Alberto Gandolfi. A free boundary problem with unilateral constraints describing the evolution of a tumor cord under the influence of cell killing agents. SIAM Journal on Mathematical Analysis, 36(3):882–915, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40606.

Burger:2003:FBP

[BFK03] Raimund Bürger, Hermano Frid, and Kenneth H. Karlsen. On a free boundary problem for a strongly degenerate quasi-linear parabolic equation with an application to a model of pressure filtration. SIAM Journal on Mathematical Analysis, 34(3):611– 635, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40100.

Barrett:2008:HMH

[BFP08] John W. Barrett, Xiaobing Feng, and Andreas Prohl. On p-harmonic map heat flows for 1 ≤ p < ∞ and their finite element approximations. SIAM Journal on Mathematical Analysis, 40(4):1471–1498, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beretta:2008:DLC

[BFV08] Elena Beretta, Elisa Francini, and Sergio Vessella. Determination of a linear crack in an elastic body from boundary measurements — Lipschitz stability. SIAM Journal on Mathematical Analysis, 40(3): 984–1002, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bass:2004:RSM

[BG04] Richard F. Bass and Karlheinz Gröchenig. Random sampling of multivariate trigonometric polynomials. SIAM Journal on Mathematical Analysis, 36 (3):773–795, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43231.

Bournaveas:2008:ASK

[BG08] Nikolaos Bournaveas and Susana Gutiérrez. Averages over spheres for kinetic transport equations with velocity derivatives in the right-hand side. SIAM Journal on Mathematical Analysis, 40(2):653–674, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benzoni-Gavage:2003:NSS

[BGHR03] S. Benzoni-Gavage, P. Huot, and F. Rousset. Nonlinear stability of semidiscrete shock waves. SIAM Journal on Mathematical Analysis, 35(3): 639–707, January 2003. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brownlee:2007:ERE

[BGL07] R. A. Brownlee, E. H. Georgoulis, and J. Levesley. Extending the range of error estimates for radial approximation in Euclidean space and on spheres. SIAM Journal on Mathematical Analysis, 39(2): 554–564, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blanchard:2008:BHR

[BGM08] Dominique Blanchard, Antonio Gaudiello, and Taras A. Mel’nyk. Boundary homogenization and reduction of dimension in a Kirchhoff–Love plate. SIAM Journal on Mathematical Analysis, 39(6):1764– 1787, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benzoni-Gavage:2001:AEF

[BGSZ01] Sylvie Benzoni-Gavage, Denis Serre, and Kevin Zumbrun. Alternate Evans functions and viscous shock waves. SIAM Journal on Mathematical Analysis, 32(5):929–962, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36183.

Bianchini:2001:SSH

[Bia01]  Stefano Bianchini.  Stabil-

ity of L^{∞} solutions for hyperbolic systems with coinciding shocks and rarefactions. SIAM Journal on Mathematical Analysis, 33(4):959–981, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37790.

Bellassoued:2008:IPD

[BIY08] M. Bellassoued, O. Imanuvilov, and M. Yamamoto. Inverse problem of determining the density and two Lamé coefficients by boundary data. SIAM Journal on Mathematical Analysis, 40(1):238–265, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bock:2009:SDC

[BJ09] Igor Bock and Jiří Jarušek.

Solvability of dynamic contact problems for elastic von Kármán plates. SIAM Journal on Mathematical Analysis, 41(1):37–45, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bjorland:2008:DAV

[Bjo08]  Clayton Bjorland.  Decay

asymptotics of the viscous Camassa–Holm equations in the plane. SIAM Journal on Mathematical Analysis, 40(2): 516–539, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertsch:2000:SDP

[BK00] M. Bertsch and S. Kamin. A system of degenerate parabolic equations from plasma physics: The large time behavior. SIAM Journal on Mathematical Analysis, 31(4):776–790, 2000. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33661.

Beretta:2002:GSS

[BK02] Edoardo Beretta and Yang Kuang. Geometric stability switch criteria in delay differential systems with delay dependent parameters. SIAM Journal on Mathematical Analysis, 33(5):1144–1165, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37608.

Bendahmane:2004:RES

[BK04] Mostafa Bendahmane and Kenneth H. Karlsen. Renormalized entropy solutions for quasi-linear anisotropic degenerate parabolic equations. SIAM Journal on Mathematical Analysis, 36(2):405–422, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42893.

Bauman:2005:ASL

[BK05] Patricia Bauman and Yangsuk Ko. Analysis of solutions to the Lawrence–Doniach system for layered superconductors. SIAM Journal on Mathematical Analysis, 37(3):914– 940, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44459.

Bresch:2004:TDH

[BKL04] Didier Bresch, Alexandre Kazhikhov, and Jérôme Lemoine. On the two-dimensional hydrostatic Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 36(3):796– 814, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42224.

Beck:2009:SLI

[BKL^{+}09] Margaret Beck, Jürgen Knobloch, David J. B. Lloyd, Björn Sandstede, and Thomas Wagenknecht. Snakes, ladders, and isolas of localized patterns. SIAM Journal on Mathematical Analysis, 41(3):936–972, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blouza:2001:BVF

[BL01] Adel Blouza and Hervé Le Dret. An up-to-the boundary version of Friedrichs’s lemma and applications to the linear Koiter shell model. SIAM Journal on Mathematical Analysis, 33(4):877–895, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38001.

Barrailh:2003:GFD

[BL03] Karen Barrailh and David Lannes. A general framework for diffractive optics and its applications to lasers with large spectrums and short pulses. SIAM Journal on Mathematical Analysis, 34(3):636–674, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39897.

Belik:2006:GCS

[BL06] Pavel Belík and Mitchell Luskin. The gamma-convergence of a sharp interface thin film model with nonconvex elastic energy. SIAM Journal on Mathematical Analysis, 38(2): 414–433, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bagland:2007:SSS

[BL07] Véronique Bagland and Philippe Laurencot. Self-similar solutions to the Oort–Hulst– Safronov coagulation equation. SIAM Journal on Mathematical Analysis, 39(2):345–378, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blanc:2006:USS

[Bla06] X. Blanc. Unique solvability of a system of nonlinear elliptic PDEs arising in solid state physics. SIAM Journal on Mathematical Analysis, 38(4): 1235–1248, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blesgen:2008:DIS

[Ble08] Thomas Blesgen. On diffusion induced segregation in time-dependent domains. SIAM Journal on Mathematical Analysis, 39(4):1260–1294, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buica:2009:ASP

[BLM09] Adriana Buică, Jaume Llibre, and Oleg Makarenkov. Asymptotic stability of periodic solutions for nonsmooth differential equations with application to the nonsmooth van der Pol oscillator. SIAM Journal on Mathematical Analysis, 40(6): 2478–2495, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bowen:2006:SPR

[BLRW06] L. Bowen, R. Lyons, C. Radin, and P. Winkler. A solidification phenomenon in random packings. SIAM Journal on Mathematical Analysis, 38(4): 1075–1089, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bulicek:2009:MAU

[BMR09] M. Bulíček, J. Málek, and K. R. Rajagopal. Mathematical analysis of unsteady flows of fluids with pressure, shear-rate, and temperature dependent material moduli that slip at solid boundaries. SIAM Journal on Mathematical Analysis, 41(2):665–707, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berlyand:2002:ENA

[BN02] Leonid Berlyand and Alexei Novikov. Error of the network approximation for densely packed composites with irregular geometry. SIAM Journal on Mathematical Analysis, 34 (2):385–408, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        39714.

Bocea:2008:CPL

[BN08] Marian Bocea and Vincenzo Nesi. Γ-convergence of powerlaw functionals, variational principles in L^{∞}, and applications. SIAM Journal on Mathematical Analysis, 39(5):1550– 1576, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braga-Neto:2004:SMS

[BNG04] Ulisses Braga-Neto and John Goutsias. Supremal multiscale signal analysis. SIAM Journal on Mathematical Analysis, 36(1):94–120, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40994.

Bellettini:2006:GSG

[BNP06] G. Bellettini, M. Novaga, and E. Paolini. Global solutions to the gradient flow equation of a nonconvex functional. SIAM Journal on Mathematical Analysis, 37(5):1657– 1687, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baro:2005:CCD

[BNR05] M. Baro, H. Neidhardt, and J. Rehberg. Current coupling of drift-diffusion models and Schrödinger–Poisson systems: Dissipative hybrid models. SIAM Journal on Mathematical Analysis, 37(3):941– 981, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61169.

Brasche:2007:CSO

[BO07] Johannes F. Brasche and Kateřina Ožanová. Convergence of Schrödinger operators. SIAM Journal on Mathematical Analysis, 39(1):281–297, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bose:2000:GAS

[Bos00]  Amitabha Bose.  A ge-

ometric approach to singularly perturbed nonlocal reaction-diffusion equations. SIAM Journal on Mathematical Analysis, 31(2):431–454, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34255.

Bostan:2003:PRV

[Bos03] M. Bostan. Permanent regimes for the 1D Vlasov–Poisson system with boundary conditions. SIAM Journal on Mathematical Analysis, 35(4):922–948, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41642.

Bostan:2005:EUM

[Bos05] M. Bostan. Existence and uniqueness of the mild solution for the 1D Vlasov–Poisson initial-boundary value problem. SIAM Journal on Mathematical Analysis, 37(1):156– 188, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43464.

Boudin:2000:SBE

[Bou00] Laurent Boudin. A solution with bounded expansion rate to the model of viscous pressureless gases. SIAM Journal on Mathematical Analysis, 32 (1):172–193, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34684.

Boussaid:2008:ASS

[Bou08] Nabile Boussaid. On the asymptotic stability of small nonlinear Dirac standing waves in a resonant case. SIAM Journal on Mathematical Analysis, 40(4):1621–1670, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bothe:2007:TCN

[BP07] Dieter Bothe and Jan Prüss.

L_{p}-theory for a class of non-Newtonian fluids. SIAM Journal on Mathematical Analysis, 39(2):379–421, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beardmore:2005:BPS

[BPBW05] R. E. Beardmore, M. A.

Peletier, C. J. Budd, and M. Ahmer Wadee. Bifurcations of periodic solutions satisfying the zero-hamiltonian constraint in reversible differential equations. SIAM Journal on Mathematical Analysis, 36(5):1461–1488, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41863.

Barnard:2000:IIG

[BPR00a] Roger W. Barnard, Kent Pearce, and Kendall C. Richards. An inequality involving the generalized hypergeometric function and the arc length of an ellipse. SIAM Journal on Mathematical Analysis, 31(3):693–699, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34157.

Barnard:2000:MPI

[BPR00b] Roger W. Barnard, Kent Pearce, and Kendall C. Richards. A monotonicity property involving _{3}F_{2} and comparisons of the classical approximations of elliptical arc length. SIAM Journal on Mathematical Analysis, 32 (2):403–419, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        35050.

Bagnerini:2003:MHH

[BR03] P. Bagnerini and M. Rascle. A multiclass homogenized hyperbolic model of traffic flow. SIAM Journal on Mathematical Analysis, 35(4):949–973, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41149.

Berselli:2006:EUW

[BR06] Luigi C. Berselli and Marco Romito. On the existence and uniqueness of weak solutions for a vorticity seeding model. SIAM Journal on Mathematical Analysis, 37(6): 1780–1799, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brandolese:2005:ARH

[Bra05] Lorenzo Brandolese. Application of the realization of homogeneous Sobolev spaces to Navier–Stokes. SIAM Journal on Mathematical Analysis, 37 (2):673–683, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44440.

Briane:2003:HHC

[Bri03] Marc Briane. Homogenization of high-conductivity periodic problems: Application to a general distribution of one-directional fibers. SIAM Journal on Mathematical Analysis, 35(1):33–60, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39866.

Bruhl:2001:ECI

[Brü01] Martin Brühl. Explicit characterization of inclusions in electrical impedance tomography.

SIAM Journal on Mathematical Analysis, 32(6):1327–1341, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36656.

Barles:2000:LTB

[BS00a] G. Barles and Panagiotis E. Souganidis. On the large time behavior of solutions of Hamilton–Jacobi equations. SIAM Journal on Mathematical Analysis, 31(4):925–939, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35086.

Bertozzi:2000:EUT

[BS00b] A. L. Bertozzi and M. Shearer. Existence of undercompressive traveling waves in thin film equations. SIAM Journal on Mathematical Analysis, 32(1):194–213, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35089.

Barles:2001:STP

[BS01] G. Barles and P. E. Souganidis. Space-time periodic solutions and long-time behavior of solutions to quasilinear parabolic equations. SIAM Journal on Mathematical Analysis, 32(6):1311–1323, 2001. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36934.

Buttazzo:2005:MOP

[BS05] Giuseppe Buttazzo and Filippo Santambrogio. A model for the optimal planning of an urban area. SIAM Journal on Mathematical Analysis, 37(2):514– 530, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43831.

Barrionuevo:2008:SSC

[BS08] José A. Barrionuevo and Jacques A. L. Silva. Stability and synchronism of certain coupled dynamical systems. SIAM Journal on Mathematical Analysis, 40(3):939– 951, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Byatt-Smith:2003:FEP

[BSB03] J. G. B. Byatt-Smith and H. W. Braden. Functional equations and Poincaré invariant mechanical systems. SIAM Journal on Mathematical Analysis, 34(3):736–758, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39374.

Buffoni:2005:MMQ

[BST05] B. Buffoni, É. Séré, and J. F. Toland. Minimization methods for quasi-linear problems with an application to periodic water waves. SIAM Journal on Mathematical Analysis, 36(4):1080–1094, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43276.

Bal:2007:ISP

[BT07] Guillaume Bal and Alexandru Tamasan. Inverse source problems in transport equations. SIAM Journal on Mathematical Analysis, 39(1):57–76, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biro:2000:AFB

[BV00a] Zs. Biro and J. J. L. Velazquez. Analysis of a free boundary problem arising in bubble dynamics. SIAM Journal on Mathematical Analysis, 32(1):142–171, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        35168.

Bonnetier:2000:ERR

[BV00b] Eric Bonnetier and Michael Vogelius. An elliptic regularity result for a composite medium with “Touching” fibers of circular cross-section. SIAM Journal on Mathematical Analysis, 31(3):651–677, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33398.

Bucur:2002:DAB

[BV02] Dorin Bucur and Nicolas Varchon. A duality approach for the boundary variation of Neumann problems. SIAM Journal on Mathematical Analysis, 34(2):460–477, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38957.

Berthelin:2005:KEM

[BV05] F. Berthelin and A. Vasseur.

From kinetic equations to multidimensional isentropic gas dynamics before shocks. SIAM Journal on Mathematical Analysis, 36(6):1807–1835, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43155.

Belishev:2008:RUS

[BV08] Mikhail I. Belishev and Alexsei F. Vakulenko. Reachable and unreachable sets in the scattering problem for the acoustical equation in R^{3}. SIAM Journal on Mathematical Analysis, 39(6):1821–1850, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonder:2004:ULS

[BW04] J. Fernández Bonder and N. Wolanski. Uniqueness of limit solutions to a free boundary problem from combustion. SIAM Journal on Mathematical Analysis, 36(1):172–185, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41293.

Beardmore:2008:NFQ

[BW08] R. Beardmore and K. Webster. Normal forms, quasi-invariant manifolds, and bifurcations of nonlinear difference-algebraic equations. SIAM Journal on Mathematical Analysis, 40(1): 413–441, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bojanov:2005:RPR

[BX05] Borislav Bojanov and Yuan Xu. Reconstruction of a polynomial from its Radon projections. SIAM Journal on Mathematical Analysis, 37(1):238– 250, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/61651.

Bayraktar:2009:AOE

[BX09] Erhan Bayraktar and Hao Xing. Analysis of the optimal exercise boundary of American options for jump diffusions. SIAM Journal on Mathematical Analysis, 41(2):825–860, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressan:2004:SDE

[BY04] Alberto Bressan and Tong Yang. A sharp decay estimate for positive nonlinear waves. SIAM Journal on Mathematical Analysis, 36(2):659–677, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42777.

Crespo:2000:GBC

[CA00] J. A. Aguilar Crespo and I. Peral Alonso. Global behavior of the Cauchy problem for some critical nonlinear parabolic equations. SIAM Journal on Mathematical Analysis, 31(6):1270–1294, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34113.

Candes:2001:RRM

[Can01] Emmanuel J. Candes. Ridgelets and the representation of mutilated Sobolev functions. SIAM Journal on Mathematical Analysis, 33(2):347–368, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36364.

Carles:2003:NSE

[Car03]  Rémi Carles.  Nonlinear

Schrödinger equations with repulsive harmonic potential and applications. SIAM Journal on Mathematical Analysis, 35 (4):823–843, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41693.

Casella:2003:EWS

[Cas03] E. Casella. On the existence of a weak solution to a two-dimensional free-boundary problem with a nonlinear flux condition. SIAM Journal on Mathematical Analysis, 35 (2):376–388, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41041.

Chen:2007:MAO

[CC07] Xinfu Chen and John Chadam. A mathematical analysis of the optimal exercise boundary for American put options. SIAM Journal on Mathematical Analysis, 38(5):1613–1641, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2006:AIF

[CCCS06] Lan Cheng, Xinfu Chen, John Chadam, and David Saunders. Analysis of an inverse first passage problem from risk management. SIAM Journal on Mathematical Analysis, 38(3): 845–873, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caceres:2002:NSC

[CCD02] María J. Cáceres, José A. Carrillo, and Jean Dolbeault. Nonlinear stability in L^{p} for a confined system of charged particles. SIAM Journal on Mathematical Analysis, 34(2):478– 494, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39843.

Cances:2005:MAN

[CCG05] Eric Cancès, Isabelle Catto, and Yousra Gati. Mathematical analysis of a nonlinear parabolic equation arising in the modelling of non-Newtonian flows. SIAM Journal on Mathematical Analysis, 37(1):60–82, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43004.

Caraballo:2007:SSR

[CCK07] Tomás Caraballo, Igor D.

Chueshov, and Peter E. Kloeden. Synchronization of a stochastic reaction-diffusion system on a thin two-layer domain. SIAM Journal on Mathematical Analysis, 38(5):1489– 1507, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cannone:2007:LEI

[CCM07] Marco Cannone, Qionglei Chen, and Changxing Miao. A losing estimate for the ideal MHD equations with application to blow-up criterion. SIAM Journal on Mathematical Analysis, 38(6):1847–1859, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Comech:2007:NIC

[CCP07] Andrew Comech, Scipio Cuccagna, and Dmitry E. Pelinovsky. Nonlinear instability of a critical traveling wave in the generalized Korteweg–de Vries equation. SIAM Journal on Mathematical Analysis, 39(1):1–33, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2007:NSE

[CCS07] C. H. Arthur Cheng, Daniel Coutand, and Steve Shkoller. Navier–Stokes equations interacting with a nonlinear elastic biofluid shell. SIAM Journal on Mathematical Analysis, 39 (3):742–800, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2001:SES

[CD01] Gui-Qiang Chen and Emmanuele DiBenedetto. Stability of entropy solutions to the Cauchy problem for a class of nonlinear hyperbolic-parabolic equations. SIAM Journal on Mathematical Analysis, 33 (4):751–762, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36359.

Chou:2008:EHD

[CD08] Kai-Seng Chou and Shi-Zhong Du. Estimates on the Hausdorff dimension of the rupture set of a thin film. SIAM Journal on Mathematical Analysis, 40(2):790–823, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:2003:SEC

[CDD03] Albert Cohen, Wolfgang Dahmen, and Ronald Devore. Sparse evaluation of compositions of functions using multiscale expansions. SIAM Journal on Mathematical Analysis, 35(2):279–303, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41207.

Cioranescu:2006:HQI

[CDD06] Doina Cioranescu, Alain Damlamian, and Riccardo De Arcangelis. Homogenization of quasiconvex integrals via the periodic unfolding method. SIAM Journal on Mathematical Analysis, 37(5):1435–1453, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cioranescu:2008:PUM

[CDG08] D. Cioranescu, A. Damlamian, and G. Griso. The periodic unfolding method in homogenization. SIAM Journal on Mathematical Analysis, 40(4): 1585–1620, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Champion:2008:WDL

[CDJ08] Thierry Champion, Luigi De Pascale, and Petri Juutinen. The ∞-Wasserstein distance: Local solutions and existence of optimal transport maps. SIAM Journal on Mathematical Analysis, 40(1):1–20, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coville:2008:EUS

[CDM08] Jérôme Coville, Juan Dávila, and Salomé Martínez. Existence and uniqueness of solutions to a nonlocal equation with monostable nonlinearity. SIAM Journal on Mathematical Analysis, 39(5):1693–1709, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chayes:2009:GEU

[CdMGGK09] Lincoln Chayes, María del Mar González, Maria Pia Gualdani, and Inwon Kim. Global existence and uniqueness of solutions to a model of price formation. SIAM Journal on Mathematical Analysis, 41 (5):2107–2135, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Costabel:2004:QDF

[CDY04] Martin Costabel, Monique Dauge, and Zohar Yosibash. A quasi-dual function method for extracting edge stress intensity functions. SIAM Journal on Mathematical Analysis, 35(5):1177–1202, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40486.

Cui:2007:BAE

[CE07] Shangbin Cui and Joachim Escher. Bifurcation analysis of an elliptic free boundary problem modelling the growth of avascular tumors. SIAM Journal on Mathematical Analysis, 39 (1):210–235, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cortazar:2009:EAB

[CEGMM09] Carmen Cortázar, Manuel Elgueta, Jorge García-Melián, and Salomé Martínez. Existence and asymptotic behavior of solutions to some inhomogeneous nonlocal diffusion problems. SIAM Journal on Mathematical Analysis, 41(5): 2136–2164, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2000:FBP

[CF00] Xinfu Chen and Avner Friedman. A free boundary problem arising in a model of wound healing. SIAM Journal on Mathematical Analysis, 32(4):778–800, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        35169.

Cordoba:2001:PCS

[CF01] Diego Cordoba and Charles Fefferman. Potato chip singularities of 3D flows. SIAM Journal on Mathematical Analysis, 33(4):786–789, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38499.

Chen:2003:FBP

[CF03] Xinfu Chen and Avner Friedman. A free boundary problem for an elliptic-hyperbolic system: An application to tumor growth. SIAM Journal on Mathematical Analysis, 35 (4):974–986, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41838.

Cianchi:2006:SFG

[CF06a] Andrea Cianchi and Adele Ferone. On symmetric functionals of the gradient having symmetric equidistributed minimizers. SIAM Journal on Mathematical Analysis, 38(1): 279–308, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cullen:2006:LSS

[CF06b] Michael Cullen and Mikhail Feldman. Lagrangian solutions of semigeostrophic equations in physical space. SIAM Journal on Mathematical Analysis, 37(5):1371–1395, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cordoba:2004:SSH

[CFdlL04] Diego Córdoba, Charles Fefferman, and Rafael de la Llave. On squirt singularities in hydrodynamics. SIAM Journal on Mathematical Analysis, 36 (1):204–213, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.

                                        org/sam-bin/dbq/article/
                                        42409.

Chen:2006:UAT

[CFG06] Xinfu Chen, Sheng-Chen Fu, and Jong-Shenq Guo. Uniqueness and asymptotics of traveling waves of monostable dynamics on lattices. SIAM Journal on Mathematical Analysis, 38(1):233–258, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caputo:2003:ASP

[CFTV03] J.-G. Caputo, N. Flytzanis, A. Tersenov, and E. Vavalis. Analysis of a semilinear PDE for modeling static solutions of Josephson junctions. SIAM Journal on Mathematical Analysis, 34(6):1356–1379, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30367.

Chuaqui:2000:CPS

[CG00] Martin Chuaqui and Julian Gevirtz. Constant principal strain mappings on 2manifolds. SIAM Journal on Mathematical Analysis, 32 (4):734–759, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        35253.

Colombo:2008:CPS

[CG08] Rinaldo M. Colombo and Mauro Garavello. On the Cauchy problem for the psystem at a junction. SIAM Journal on Mathematical Analysis, 39(5):1456–1471, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2004:ANS

[CGL04] Louis H. Y. Chen, Tim N. T.

Goodman, and S. L. Lee. Asymptotic normality of scaling functions. SIAM Journal on Mathematical Analysis, 36 (1):323–346, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40622.

Colin:2005:IMN

[CGL05] Thierry Colin, Gérard Gallice, and Karen Laurioux. Intermediate models in nonlinear optics. SIAM Journal on Mathematical Analysis, 36(5):1664–1688, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42306.

Carlen:2008:DSG

[CGL08] Eric A. Carlen, Jeffrey S.

Geronimo, and Michael Loss. Determination of the spectral gap in the Kac model for physical momentum and energy-conserving collisions. SIAM Journal on Mathematical Analysis, 40(1):327–364, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Calderer:2002:TEN

[CGLL02] M. C. Calderer, D. Golovaty, F.-H. Lin, and C. Liu. Time evolution of nematic liquid crystals with variable degree of orientation. SIAM Journal on Mathematical Analysis, 33(5):1033–1047, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36208.

Chang:2008:SLO

[CGNT08] Shu-Ming Chang, Stephen Gustafson, Kenji Nakanishi, and Tai-Peng Tsai. Spectra of linearized operators for NLS solitary waves. SIAM Journal on Mathematical Analysis, 39 (4):1070–1111, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coclite:2005:TFR

[CGP05] G. M. Coclite, M. Garavello, and B. Piccoli. Traffic flow on a road network. SIAM Journal on Mathematical Analysis, 36(6):1862–1886, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40268.

Chen:2009:IAQ

[CGZ09] Shuxing Chen, Jinbo Geng, and Yongqian Zhang. Isentropic approximation of quasi– one-dimensional unsteady nozzle flow. SIAM Journal on Mathematical Analysis, 41(4): 1693–1712, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chae:2006:NLF

[CH06] Myeongju Chae and Seung-Yeal Ha. New Lyapunov functionals of the Vlasov–Poisson system. SIAM Journal on Mathematical Analysis, 37(6): 1709–1731, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chae:2006:RCD

[Cha06] Dongho Chae. On the regularity conditions for the dissipative quasi-geostrophic equations. SIAM Journal on Mathematical Analysis, 37(5):1649– 1656, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2000:STW

[Che00] Di-Rong Chen. On the splitting trick and wavelet frame packets. SIAM Journal on Mathematical Analysis, 31(4): 726–739, 2000. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32333.

Coclite:2005:GWS

[CHK05] G. M. Coclite, H. Holden, and K. H. Karlsen. Global weak solutions to a generalized hyperelastic-rod wave equation. SIAM Journal on Mathematical Analysis, 37(4): 1044–1069, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
61671.

Cox:2008:CGM

[CHK08] Dennis D. Cox, Robert M. Hardt, and Petr Klouček. Convergence of Gibbs measures associated with simulated annealing. SIAM Journal on Mathematical Analysis, 39(5): 1472–1496, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colombo:2008:CLJ

[CHS08] R. M. Colombo, M. Herty, and V. Sachers. On 2 × 2 conservation laws at a junction. SIAM Journal on Mathematical Analysis, 40(2):605–622, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2004:AMP

[CJ04] Li Chen and Ansgar Jüngel. Analysis of a multidimensional parabolic population model with strong cross-diffusion. SIAM Journal on Mathematical Analysis, 36(1):301–322, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42779.

Ciuperca:2009:ESE

[CJT09] I. Ciuperca, M. Jai, and J. I.

Tello. On the existence of solutions of equilibria in lubricated journal bearings. SIAM Journal on Mathematical Analysis, 40(6):2316–2327, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Canuto:2001:DCC

[CK01a] B. Canuto and O. Kavian.

Determining coefficients in a class of heat equations via boundary measurements. SIAM Journal on Mathematical Analysis, 32(5):963–986, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36525.

Chen:2001:DIS

[CK01b] Xinfu Chen and MichaL Kowalczyk. Dynamics of an interior spike in the Gierer– Meinhardt system. SIAM Journal on Mathematical Analysis, 33(1):172–193, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36495.

Canic:2006:FBP

[CKK06] Suncica Canic, Barbara Lee Keyfitz, and Eun Heui Kim. Free boundary problems for nonlinear wave systems: Mach stems for interacting shocks. SIAM Journal on Mathematical Analysis, 37(6):1947–1977, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colliander:2001:GWP

[CKS^{+}01] J. Colliander, M. Keel, G. Staffilani, H. Takaoka, and T. Tao. Global well-posedness for Schrödinger equations with derivative. SIAM Journal on Mathematical Analysis, 33 (3):649–669, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        38438.

Colliander:2002:RGW

[CKS^{+}02] J. Colliander, M. Keel, G. Staffilani, H. Takaoka, and T. Tao. A refined global well-posedness result for Schrödinger equations with derivative. SIAM Journal on Mathematical Analysis, 34(1):64–86, 2002. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39454.

Chicone:2000:CIT

[CL00] Carmen Chicone and Weishi Liu. On the continuation of an invariant torus in a family with rapid oscillations. SIAM Journal on Mathematical Analysis, 31(2):386–415, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33874.

Chen:2003:FSV

[CL03] Gui-Qiang Chen and Hailiang Liu. Formation of δshocks and vacuum states in the vanishing pressure limit of solutions to the Euler equations for isentropic fluids. SIAM Journal on Mathematical Analysis, 34(4):925– 938, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39935.

Corbera:2004:FPO

[CL04] Montserrat Corbera and Jaume Llibre. Families of periodic orbits for the spatial isosceles 3body problem. SIAM Journal on Mathematical Analysis, 35(5):1311–1346, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40788.

Cermelli:2005:REF

[CL05] Paolo Cermelli and Giovanni Leoni. Renormalized energy and forces on dislocations. SIAM Journal on Mathematical Analysis, 37(4):1131–1160, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/62163.

Colin:2009:SPA

[CL09] Mathieu Colin and David Lannes. Short pulses approximations in dispersive media. SIAM Journal on Mathematical Analysis, 41(2):708–732, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clarkson:2000:ACM

[Cla00]  Eric Clarkson.  Angular

channels in a multidimensional wavelet transform. SIAM Journal on Mathematical Analysis, 32(1):80–102, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30961.

Chen:2006:ASN

[CLGZ06] Wu-Hua Chen, Xiaomei Lu, Zhi-Hong Guan, and Wei Xing Zheng. Asymptotic stability in a neutral delay differential system with variable delays. SIAM Journal on Mathematical Analysis, 37(5):1522– 1534, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Comte:2002:FDH

[CLR02] M. Comte and T. Lachand-Robert. Functions and domains having minimal resistance under a single-impact assumption. SIAM Journal on Mathematical Analysis, 34 (1):101–120, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/

38884. See comment [Pla09b].

Cox:1998:STC

[CM98] Steven J. Cox and C. Maeve McCarthy. The shape of the tallest column. SIAM Journal on Mathematical Analysis, 29(3):547–554, May 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31453. See cor-

rection [CM00].

Cox:2000:STC

[CM00] Steven J. Cox and C. Maeve McCarthy. The shape of the tallest column: Corrected. SIAM Journal on Mathematical Analysis, 31(4):940, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35495.  See

[CM98].

Chemin:2001:ALR

[CM01] Jean-Yves Chemin and Nader Masmoudi. About lifespan of regular solutions of equations related to viscoelastic fluids. SIAM Journal on Mathematical Analysis, 33(1):84– 112, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35931.

Carstensen:2002:LSR

[CM02] Carsten Carstensen and Stefan Müller. Local stress regularity in scalar nonconvex variational problems. SIAM Journal on Mathematical Analysis, 34(2):495–509, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        39643.

Chaudhuri:2006:SET

[CM06] Nirmalendu Chaudhuri and Stefan Müller. Scaling of the energy for thin Martensitic films. SIAM Journal on Mathematical Analysis, 38(2):468– 477, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Costanzino:2009:SWR

[CMJ09] Nicola Costanzino, Vahagn Manukian, and Christopher K. R. T. Jones. Solitary waves of the regularized short pulse and Ostrovsky equations.

SIAM Journal on Mathematical Analysis, 41(5):2088–2106, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Conti:2006:RDF

[CMM06] Sergio Conti, Francesco Maggi, and Stefan Müller. Rigorous derivation of Föppl’s theory for clamped elastic membranes leads to relaxation. SIAM Journal on Mathematical Analysis, 38(2):657–680, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Citti:2004:NOV

[CMS04] Giovanna Citti, Maria Manfredini, and Alessandro Sarti. Neuronal oscillations in the visual cortex: Γ-Convergence to the Riemannian Mumford– Shah functional. SIAM Journal on Mathematical Analysis, 35(6):1394–1419, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39867.

Chen:2008:WPV

[CMZ08] Qionglei Chen, Changxing Miao, and Zhifei Zhang. On the well-posedness for the viscous shallow water equations. SIAM Journal on Mathematical Analysis, 40(2):443–474, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Craig:2000:TTT

[CN00] Walter Craig and David P. Nicholls. Traveling two and three dimensional capillary gravity water waves. SIAM Journal on Mathematical Analysis, 32(2):323–359, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35418.

Castro:2001:LIP

[CN01] Alfonso Castro and J. W. Neuberger. A local inversion principle of the Nash–Moser type. SIAM Journal on Mathematical Analysis, 33(4):989–993, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35769.

Chambolle:2006:CAA

[CN06] Antonin Chambolle and Matteo Novaga. Convergence of an algorithm for the anisotropic and crystalline mean curvature flow. SIAM Journal on Mathematical Analysis, 37(6):1978– 1987, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Conti:2006:CRC

[CNO06] Sergio Conti, Barbara Niethammer, and Felix Otto. Coarsening rates in off-critical mixtures. SIAM Journal on Mathematical Analysis, 37(6):

1732–1741, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2005:PUG

[CO05] Xinfu Chen and Yoshihito Oshita. Periodicity and uniqueness of global minimizers of an energy functional containing a long-range interaction. SIAM Journal on Mathematical Analysis, 37(4):1299–1332, January 2005. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cho:2006:SHT

[CO06] Yonggeun Cho and Tohru Ozawa. On the semirelativistic Hartree-type equation. SIAM Journal on Mathematical Analysis, 38(4):1060– 1074, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cooper:2000:PRW

[Coo00] Jeffery Cooper. Parametric resonance in wave equations with a time-periodic potential. SIAM Journal on Mathematical Analysis, 31(4):821– 835, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34070.

Coulombel:2002:WSN

[Cou02] Jean-François Coulombel. Weak stability of nonuniformly stable multidimensional shocks.

SIAM Journal on Mathematical Analysis, 34(1):142–172, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39280.

Conca:2002:BAH

[COV02] Carlos Conca, Rafael Orive, and Muthusamy Vanninathan. Bloch approximation in homogenization and applications. SIAM Journal on Mathematical Analysis, 33(5):1166–1198, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38220.

Cozzi:2009:VVP

[Coz09] Elaine Cozzi. Vanishing viscosity in the plane for nondecaying velocity and vorticity. SIAM Journal on Mathematical Analysis, 41(2):495– 510, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:2000:OAT

[CP00] Albert Cohen and Benoit Perthame. Optimal approximations of transport equations by particle and pseudoparticle methods. SIAM Journal on Mathematical Analysis, 32 (3):616–636, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.

org/sam-bin/dbq/article/
35035.

Calderer:2007:YMO

[CP07a] Maria-Carme Calderer and Alexander Panchenko. Young measures and order-disorder transition in stationary flow of liquid crystals. SIAM Journal on Mathematical Analysis, 38 (5):1642–1659, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Celada:2007:CNB

[CP07b] P. Celada and S. Perrotta. On a class of nonconvex Bolza problems related to Blatz–Ko elastic materials. SIAM Journal on Mathematical Analysis, 39(3):838–862, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2007:SEM

[CQ07] Xinfu Chen and Yuanwei Qi. Sharp estimates on minimum travelling wave speed of reaction diffusion systems modelling autocatalysis. SIAM Journal on Mathematical Analysis, 39(2):437–448, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2003:LTB

[CQW03] Xinfu Chen, Y. W. Qi, and Mingxin Wang. Long time behavior of solutions to P - Laplacian equation with absorption. SIAM Journal on Mathematical Analysis, 35(1): 123–134, 2003. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40772.

Chen:2003:MPA

[CR03] Yunmei Chen and Murali Rao. Minimization problems and associated flows related to weighted p energy and total variation. SIAM Journal on Mathematical Analysis, 34(5):1084–1104, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40457.

Coclite:2005:CLT

[CR05] Giuseppe Maria Coclite and Nils Henrik Risebro. Conservation laws with time dependent discontinuous coefficients. SIAM Journal on Mathematical Analysis, 36(4):1293–1309, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42000.

Cardaliaguet:2006:VSI

[CR06] Pierre Cardaliaguet and Elisabeth Rouy. Viscosity solutions of increasing flows of sets. application of the Hele– Shaw problem for power-law fluids. SIAM Journal on Mathematical Analysis, 38(1):143– 165, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crepey:2003:CLV

[Cré03] S. Crépey. Calibration of the local volatility in a generalized Black–Scholes model using Tikhonov regularization. SIAM Journal on Mathematical Analysis, 34(5):1183–1206, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40020.

Chen:2005:PRS

[CRTW05] Yunmei Chen, M. Rao, Y. Tonegawa, and T. Wunderli. Partial regularity for a selective smoothing functional for image restoration in BV space. SIAM Journal on Mathematical Analysis, 37(4):1098–1116, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/62388.

Cuesta:2006:WSO

[CS06] C. M. Cuesta and C. Schmeiser.

Weak shocks for a onedimensional BGK kinetic model for conservation laws. SIAM Journal on Mathematical Analysis, 38(2):637–656, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chambolle:2007:IBS

[CS07] Antonin Chambolle and Margherita Solci. Interaction of a bulk and a surface energy with a geometrical constraint. SIAM Journal on Mathematical Analysis, 39 (1):77–102, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2008:LTE

[CT08] Bin Cheng and Eitan Tadmor. Long-time existence of smooth solutions for the rapidly rotating shallow-water and Euler equations. SIAM Journal on Mathematical Analysis, 39(5): 1668–1685, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuccagna:2009:ASS

[CT09] Scipio Cuccagna and Mirko Tarulli. On asymptotic stability of standing waves of discrete Schrödinger equation in Z. SIAM Journal on Mathematical Analysis, 41(3):861– 885, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cui:2008:ASS

[Cui08] Shangbin Cui. Asymptotic stability of the stationary solution for a hyperbolic free boundary problem modeling tumor growth. SIAM Journal on Mathematical Analysis, 40(4): 1692–1724, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cazenave:2000:STS

[CW00] Thierry Cazenave and Fred B.

Weissler. Scattering theory and self-similar solutions for the nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 31(3): 625–650, 2000. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35130.

Chandler-Wilde:2005:EUV

[CWM05] Simon N. Chandler-Wilde and Peter Monk. Existence, uniqueness, and variational methods for scattering by unbounded rough surfaces. SIAM Journal on Mathematical Analysis, 37(2):598– 618, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61552.

Chandler-Wilde:2008:WNE

[CWM08] Simon N. Chandler-Wilde and Peter Monk. Wave-number-explicit bounds in timeharmonic scattering. SIAM Journal on Mathematical Analysis, 39(5):1428–1455, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2009:EDF

[CWY09] Gui-Qiang Chen, Dehua Wang, and Xiaozhou Yang. Evolution of discontinuity and formation of triple-shock pattern in solutions to a two-dimensional hyperbolic system of conservation laws. SIAM Journal on Mathematical Analysis, 41(1):1–25, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2004:DTC

[CY04] Jin Cheng and Masahiro Yamamoto. Determination of two convection coefficients from Dirichlet to Neumann map in the two-dimensional case. SIAM Journal on Mathematical Analysis, 35(6):1371–1393, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42249.

Chen:2007:SCV

[CZZ07] Gui-Qiang Chen, Yongqian Zhang, and Dianwen Zhu. Stability of compressible vortex sheets in steady supersonic Euler flows over Lipschitz walls. SIAM Journal on Mathematical Analysis, 38(5):1660–1693, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dalibard:2007:KFP

[Dal07] Anne-Laure Dalibard. Kinetic formulation for a parabolic conservation law. application to homogenization. SIAM Journal on Mathematical Analysis, 39(3):891–915, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dalibard:2009:ABR

[Dal09] Anne-Laure Dalibard. Asymptotic behavior of a rapidly rotating fluid with random stationary surface stress. SIAM Journal on Mathematical Analysis, 41(2):511–563, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dancer:2002:ECD

[DD02a] E. N. Dancer and Yihong Du.

Effects of certain degeneracies in the predator-prey model. SIAM Journal on Mathematical Analysis, 34(2):292–314, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38759.

Ding:2002:CMF

[DD02b] Shijin Ding and Qiang Du.

Critical magnetic field and asymptotic behavior of superconducting thin films. SIAM Journal on Mathematical Analysis, 34(1):239–256, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37861.

Davila:2007:SSB

[DDGM07] Juan Dávila, Louis Dupaigne, Ignacio Guerra, and Marcelo Montenegro. Stable solutions for the bilaplacian with exponential nonlinearity. SIAM Journal on Mathematical Analysis, 39(2):565–592, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Delshams:2000:KTP

[DdlL00] Amadeu Delshams and Rafael de la Llave. KAM theory and a partial justification of Greene’s criterion for nontwist maps. SIAM Journal on Mathematical Analysis, 31(6):1235–1269, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34908.

DeBouard:2004:PSK

[DDT04] A. De Bouard, A. Debussche, and Y. Tsutsumi. Periodic solutions of the Korteweg–de Vries equation driven by white noise. SIAM Journal on Mathematical Analysis, 36(3):815– 855, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42530.

Docksey:2000:CMS

[DE00] R. G. Docksey and J. N. Elgin. Closure of the Manakov system. SIAM Journal on Mathematical Analysis, 32(1):54– 79, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34367.

Dunca:2006:SAD

[DE06] A. Dunca and Y. Epshteyn.

On the Stolz–Adams deconvolution model for the largeeddy simulation of turbulent flows. SIAM Journal on Mathematical Analysis, 37(6):1890– 1902, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dorfman:2009:LRH

[DE09] Jonathan Dorfman and Lawrence C.

Evans. A “Lakes and rivers” heuristic metaphor for the singular limit of a nonlinear diffusion PDE. SIAM Journal on Mathematical Analysis, 41(4): 1621–1652, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deuring:2009:SDT

[Deu09] Paul Deuring. Spatial decay of time-dependent Oseen flows. SIAM Journal on Mathematical Analysis, 41(3):886–922, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Desenzani:2006:CGL

[DF06] N. Desenzani and I. Fragalà.

Concentration of Ginzburg– Landau energies with supercritical growth. SIAM Journal on Mathematical Analysis, 38(2):385–413, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duffy:2002:MDI

[DFG02] Brian R. Duffy, Pedro Freitas, and Michael Grinfeld. Memory driven instability in a diffusion process. SIAM Journal on Mathematical Analysis, 33(5):1090–1106, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38859.

DalMaso:2009:HOM

[DFLM09] G. Dal Maso, I. Fonseca, G. Leoni, and M. Morini. A higher order model for image restoration: The onedimensional case. SIAM Journal on Mathematical Analysis, 40(6):2351–2391, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Donato:2009:CHC

[DFM09] Patrizia Donato, Luisa Faella, and Sara Monsurrò. Correctors for the homogenization of a class of hyperbolic equations with imperfect interfaces. SIAM Journal on Mathematical Analysis, 40(5):1952–1978, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diaz:2000:ABN

[DG00] Juan Casado Diaz and Adriana Garroni. Asymptotic behavior of nonlinear elliptic systems on varying domains. SIAM Journal on Mathematical Analysis, 31(3):581–624, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32962.

Deckelnick:2009:SSN

[DG09] Klaus Deckelnick and Hans-Christoph Grunau. Stability and symmetry in the Navier problem for the onedimensional Willmore equation. SIAM Journal on Mathematical Analysis, 40(5):2055– 2076, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duzaar:2000:PRA

[DGG00] Frank Duzaar, Andreas Gastel, and Joseph F. Grotowski. Partial regularity for almost minimizers of quasi-convex integrals. SIAM Journal on Mathematical Analysis, 32(3):665– 687, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37453.

Diekmann:2008:SBA

[DGG08] Odo Diekmann, Philipp Getto, and Mats Gyllenberg. Stability and bifurcation analysis of Volterra functional equations in the light of suns and stars. SIAM Journal on Mathematical Analysis, 39(4):1023–1069, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dobric:2000:SPC

[DGH00] V. Dobric, R. F. Gundy, and P. Hitczenko. Stability properties for a compactly supported prescale function. SIAM Journal on Mathematical Analysis, 31(3):574–580, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32746.

DallAglio:2004:RPE

[DGP04] Andrea Dall’Aglio, Daniela Giachetti, and Ireneo Peral. Results on parabolic equations related to some Caffarelli– Kohn–Nirenberg inequalities. SIAM Journal on Mathematical Analysis, 36(3):691–716, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43235.

Du:2008:CPNa

[DH08a] Yihong Du and Sze-Bi Hsu. Concentration phenomena in a nonlocal quasi-linear problem modelling phytoplankton I: Existence. SIAM Journal on Mathematical Analysis, 40(4): 1419–1440, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2008:CPNb

[DH08b] Yihong Du and Sze-Bi Hsu.

Concentration phenomena in a nonlocal quasi-linear problem modelling phytoplankton II: Limiting profile. SIAM Journal on Mathematical Analysis, 40 (4):1441–1470, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ding:2003:TWS

[Din03]  Zhonghai Ding.  Traveling

waves in a suspension bridge system. SIAM Journal on Mathematical Analysis, 35(1): 160–171, 2003. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41269.

Doelman:2004:DFC

[DIN04] Arjen Doelman, David Iron, and Yasumasa Nishiura. Destabilization of fronts in a class of bistable systems. SIAM Journal on Mathematical Analysis, 35(6):1420–1450, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41924.

Daners:2008:UFK

[DK08] Daniel Daners and James Kennedy. Uniqueness in the Faber–Krahn inequality for Robin problems. SIAM Journal on Mathematical Analysis, 39(4):1191–1207, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Doelman:2007:NAS

[DKP07] Arjen Doelman, Tasso J. Kaper, and Keith Promislow. Nonlinear asymptotic stability of the semistrong pulse dynamics in a regularized Gierer– Meinhardt model. SIAM Journal on Mathematical Analysis, 38(6):1760–1787, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dziuk:2002:EEC

[DKS02] Gerhard Dziuk, Ernst Kuwert, and Reiner Schatzle. Evolution of elastic curves in R^{n}: Existence and computation. SIAM Journal on Mathematical Analysis, 33(5):1228–1245, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38370.

Ding:2000:PSH

[DL00] Yanheng Ding and Cheng Lee. Periodic solutions of Hamiltonian systems. SIAM Journal on Mathematical Analysis, 32 (3):555–571, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.

org/sam-bin/dbq/article/
35817.

Dekel:2004:BHL

[DL04] S. Dekel and D. Leviatan. The Bramble–Hilbert lemma for convex domains. SIAM Journal on Mathematical Analysis, 35(5):1203–1212, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41758.

Diaz:2005:LSS

[DLS05] J. I. Díaz, M. Lazzo, and P. G. Schmidt. Large solutions for a system of elliptic equations arising from fluid dynamics. SIAM Journal on Mathematical Analysis, 37(2):490– 513, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44355.

Dyn:2007:AUN

[DLY07] Nira Dyn, David Levin, and Jungho Yoon. Analysis of univariate nonstationary subdivision schemes with application to Gaussian-based interpolatory schemes. SIAM Journal on Mathematical Analysis, 39(2):470–488, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deMonvel:2009:LTA

[dMKST09] Anne Boutet de Monvel, Aleksey Kostenko, Dmitry Shepelsky, and Gerald Teschl. Long-time asymptotics for the Camassa–Holm equation. SIAM Journal on Mathematical Analysis, 41(4):1559–1588, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dapice:2006:PFT

[DMP06] Ciro D’apice, Rosanna Manzo, and Benedetto Piccoli. Packet flow on telecommunication networks. SIAM Journal on Mathematical Analysis, 38(3): 717–740, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duncan:2009:SSE

[DMPD09] T. E. Duncan, B. Maslowski, and B. Pasik-Duncan. Semilinear stochastic equations in a Hilbert space with a fractional Brownian motion. SIAM Journal on Mathematical Analysis, 40(6):2286–2315, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dette:2009:SAP

[DN09] Holger Dette and Jan Nagel. Some asymptotic properties of the spectrum of the Jacobi ensemble. SIAM Journal on Mathematical Analysis, 41(4): 1491–1507, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dirr:2009:SIL

[DO09] Nicolas Dirr and Enza Orlandi. Sharp-interface limit of a Ginzburg–Landau functional with a random external field. SIAM Journal on Mathematical Analysis, 41(2): 781–824, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Donoho:2000:ORL

[Don00] David L. Donoho. Orthonormal ridgelets and linear singularities. SIAM Journal on Mathematical Analysis, 31(5): 1062–1099, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34440.

Davydov:2003:NAD

[DP03] Oleg Davydov and Pencho Petrushev. Nonlinear approximation from differentiable piecewise polynomials. SIAM Journal on Mathematical Analysis, 35(3):708–758, January 2003. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2005:UBC

[DP05] Shibin Dai and Robert L.

Pego. Universal bounds on coarsening rates for mean-field models of phase transitions. SIAM Journal on Mathematical Analysis, 37(2):347–371, 2005. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/61804.

Dafermos:2009:GSS

[DP09] Constantine M. Dafermos and Ronghua Pan. Global BV solutions for the P - system with frictional damping. SIAM Journal on Mathematical Analysis, 41(3):1190– 1205, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

delPino:2007:RIL

[dPKW07] Manuel del Pino, Micha Kowalczyk, and Juncheng Wei. Resonance and interior layers in an inhomogeneous phase transition model. SIAM Journal on Mathematical Analysis, 38(5):1542–1564, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dohnal:2009:LML

[DPR09] Tomáš Dohnal, Michael Plum, and Wolfgang Reichel. Localized modes of the linear periodic Schrödinger operator with a nonlocal perturbation. SIAM Journal on Mathematical Analysis, 41(5):1967–1993, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dancer:2007:MRS

[DRY07] Edward Norman Dancer, Xiaofeng Ren, and Shusen Yan. On multiple radial solutions of a singularly perturbed nonlinear elliptic system. SIAM Journal on Mathematical Analysis, 38(6):2005– 2041, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dirr:2005:LTB

[DS05] Nicolas Dirr and Panagiotis E.

Souganidis. Large-time behavior for viscous and nonviscous Hamilton–Jacobi equations forced by additive noise. SIAM Journal on Mathematical Analysis, 37(3):777–796, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61189.

Davini:2006:GDA

[DS06] Andrea Davini and Antonio Siconolfi. A generalized dynamical approach to the large time behavior of solutions of Hamilton–Jacobi equations. SIAM Journal on Mathematical Analysis, 38(2):478– 502, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daneri:2008:ECD

[DS08] Sara Daneri and Giuseppe Savaré. Eulerian calculus for the displacement convexity in the Wasserstein distance. SIAM Journal on Mathematical Analysis, 40(3):1104–1122, ???? 2008. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Dull:2009:WTP

[DS09] Wolf-Patrick Düll and Guido Schneider. A waiting time phenomenon in pattern forming systems. SIAM Journal on Mathematical Analysis, 41 (1):415–433, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2000:EMS

[Du00] Yihong Du. Exact multiplicity and S-shaped bifurcation curve for some semilinear elliptic problems from combustion theory. SIAM Journal on Mathematical Analysis, 32(4):707– 733, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34358.

Dumas:2003:ECS

[Dum03] Éric Dumas. An example of catastrophic self-focusing in nonlinear optics? SIAM Journal on Mathematical Analysis, 35(1):268–278, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        41448.

Dung:2000:CC

[Dun00] Le Dung. Coexistence with chemotaxis. SIAM Journal on Mathematical Analysis, 32 (3):504–521, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34677.

Dunster:2001:UAE

[Dun01]  T. M. Dunster.  Uniform

asymptotic expansions for the reverse generalized Bessel polynomials, and related functions. SIAM Journal on Mathematical Analysis, 32(5):987–1013, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35906.

deVilliers:2003:DDS

[dVGH03] J. M. de Villiers, K. M.

Goosen, and B. M. Herbst. Dubuc–Deslauriers subdivision for finite sequences and interpolation wavelets on an interval. SIAM Journal on Mathematical Analysis, 35(2):423– 452, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38683.

DAprile:2005:BSC

[DW05] Teresa D’Aprile and Juncheng Wei. On bound states concentrating on spheres for the Maxwell–Schrödinger equation. SIAM Journal on Mathematical Analysis, 37(1):321– 342, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/44279.

Donoho:2000:NPT

[DY00] David L. Donoho and Thomas P.-Y. Yu. Nonlinear pyramid transforms based on medianinterpolation. SIAM Journal on Mathematical Analysis, 31(5):1030–1061, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33029.

Du:2003:EWS

[DZ03] Qiang Du and Ping Zhang. Existence of weak solutions to some vortex density models. SIAM Journal on Mathematical Analysis, 34(6):1279–1299, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40800.

Dupuis:2008:ESC

[DZ08] Paul Dupuis and Jim X. Zhang. Explicit solutions for a class of nonlinear PDEs that arise in allocation problems. SIAM Journal on Mathematical Analysis, 39(5):1627–1667, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ebmeyer:2000:NEP

[Ebm00] Carsten Ebmeyer. Nonlinear elliptic problems under mixed boundary value conditions in nonsmooth domains. SIAM Journal on Mathematical Analysis, 32(1):103–118, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34986.

Elliott:2003:DBC

[EGG03] C. M. Elliott, Y. Giga, and S. Goto. Dynamic boundary conditions for Hamilton– Jacobi equations. SIAM Journal on Mathematical Analysis, 34(4):861–881, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39957.

Evans:2006:BGA

[EGW06] J. D. Evans, V. A. Galaktionov, and J. F. Williams. Blow-up and global asymptotics of the limit unstable Cahn–Hilliard equation. SIAM Journal on Mathematical Analysis, 38(1):64–102, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ElSoufi:2008:EFD

[EK08] Ahmad El Soufi and Rola Kiwan. Extremal first Dirichlet eigenvalue of doubly connected plane domains and dihedral symmetry. SIAM Journal on Mathematical Analysis, 39(4): 1112–1119, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eleuteri:2009:MFH

[EKK09] Michela Eleuteri, Jana Kopfová, and Pavel Krejčí. Magnetohydrodynamic flow with hysteresis. SIAM Journal on Mathematical Analysis, 41(2): 435–464, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ElHajj:2007:WPT

[El 07]  Ahmad El Hajj.  Well-

posedness theory for a nonconservative Burgers-type system arising in dislocation dynamics. SIAM Journal on Mathematical Analysis, 39(3):965– 986, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eschwe:2002:TVP

[EL02] David Eschwé and Heinz Langer. Triple variational principles for eigenvalues of selfadjoint operators and operator functions. SIAM Journal on Mathematical Analysis, 34 (1):228–238, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        38774.

Eisenberg:2007:PNP

[EL07] Bob Eisenberg and Weishi Liu. Poisson–Nernst–Planck systems for ion channels with permanent charges. SIAM Journal on Mathematical Analysis, 38 (6):1932–1966, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Escher:2009:ESR

[EM09] Joachim Escher and Bogdan-Vasile Matioc. Existence and stability results for periodic Stokesian Hele–Shaw flows. SIAM Journal on Mathematical Analysis, 40(5):1992–2006, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evans:2005:DNH

[ESG05] L. C. Evans, O. Savin, and W. Gangbo. Diffeomorphisms and nonlinear heat flows. SIAM Journal on Mathematical Analysis, 37(3):737–751, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61386.

Esposito:2005:BSL

[Esp05] Pierpaolo Esposito. Blowup solutions for a Liouville equation with singular data. SIAM Journal on Mathematical Analysis, 36(4):1310–1345, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43054.

Ettinger:2009:GEU

[ET09] Boris Ettinger and Edriss S. Titi. Global existence and uniqueness of weak solutions of three-dimensional Euler equations with helical symmetry in the absence of vorticity stretching. SIAM Journal on Mathematical Analysis, 41(1): 269–296, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:2006:ATW

[FC06] Avner Friedman and Gheorghe Craciun. Approximate traveling waves in linear reaction-hyperbolic equations. SIAM Journal on Mathematical Analysis, 38(3):741– 758, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fontelos:2002:MAM

[FFH02] Marco A. Fontelos, Avner Friedman, and Bei Hu. Mathematical analysis of a model for the initiation of angiogenesis. SIAM Journal on Mathematical Analysis, 33(6):1330–1355, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38504.

Friesecke:2009:ELN

[FG09] Gero Friesecke and Benjamin D. Goddard. Explicit large nuclear charge limit of electronic ground states for Li, Be, B, C, N, O, F, Ne and basic aspects of the periodic table. SIAM Journal on Mathematical Analysis, 41(2):631–664, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fibich:2007:CKS

[FGHS07] G. Fibich, I. Gannot, A. Hammer, and S. Schochet. Chemical kinetics on surfaces: A singular limit of a reactiondiffusion system. SIAM Journal on Mathematical Analysis, 38(5):1371–1388, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frehse:2005:SLS

[FGM05] Jens Frehse, Sonja Goj, and Josef Málek. On a Stokes-like system for mixtures of fluids. SIAM Journal on Mathematical Analysis, 36(4):1259–1281, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43342.

Feng:2004:DSL

[FGQ04] Shui Feng, Ilie Grigorescu, and Jeremy Quastel. Diffusive scaling limits of mutually interacting particle systems. SIAM Journal on Mathematical Analysis, 35(6):1512–1533, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40952.

Fife:2001:NOF

[FH01] Paul C. Fife and Danielle Hilhorst. The Nishiura–Ohnishi free boundary problem in the 1D case. SIAM Journal on Mathematical Analysis, 33(3): 589–606, 2001. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        37250.

Friedman:2007:BFB

[FH07] Avner Friedman and Bei Hu. Bifurcation for a free boundary problem modeling tumor growth by Stokes equation. SIAM Journal on Mathematical Analysis, 39(1):174–194, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Faria:2002:SCM

[FHW02] Teresa Faria, Wenzhang Huang, and Jianhong Wu. Smoothness of center manifolds for maps and formal adjoints for semilinear FDEs in general Banach spaces. SIAM Journal on Mathematical Analysis, 34 (1):173–203, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        38497.

Figalli:2007:EUR

[Fig07]  Alessio Figalli.  Existence,

uniqueness, and regularity of optimal transport maps. SIAM Journal on Mathematical Analysis, 39(1):126–137, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Filho:2007:VDT

[Fil07] M. C. Lopes Filho. Vortex dynamics in a twodimensional domain with holes and the small obstacle limit. SIAM Journal on Mathematical Analysis, 39(2):422–436, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Finkelshtein:2009:IBM

[FKK09] Dmitri Finkelshtein, Yuri Kondratiev, and Oleksandr Kutoviy. Individual based model with competition in spatial ecology. SIAM Journal on Mathematical Analysis, 41(1): 297–317, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fainsilber:2006:LPC

[FKW06] Laura Fainsilber, Pär Kurlberg, and Bernt Wennberg. Lattice points on circles and discrete velocity models for the Boltzmann equation. SIAM Journal on Mathematical Analysis, 37(6):1903– 1922, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fiedler:2000:GHB

[FL00] Bernold Fiedler and Stefan Liebscher. Generic Hopf bifurcation from lines of equilibria without parameters: II. systems of viscous hyperbolic balance laws. SIAM Journal on Mathematical Analysis, 31(6):1396–1404, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34172.

Fitzgibbon:2001:MMS

[FLM01] W. E. Fitzgibbon, M. Langlais, and J. J. Morgan. A mathematical model of the spread of feline leukemia virus (FeLV) through a highly heterogeneous spatial domain. SIAM Journal on Mathematical Analysis, 33(3):570–588, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37175.

Filho:2005:ILT

[FLP05] M. C. Lopes Filho, H. J. Nussenzveig Lopes, and G. Planas. On the inviscid limit for two-dimensional incompressible flow with Navier friction condition. SIAM Journal on Mathematical Analysis, 36(4):1130–1141, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43234.

Forster:2009:ACL

[FLT09] Barbara Forster, Eva Lütkebohmert, and Josef Teichmann. Absolutely continuous laws of jumpdiffusions in finite and infinite dimensions with applications to mathematical finance. SIAM Journal on Mathematical Analysis, 40(5):2132–2153, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fonseca:2000:SOS

[FM00] Irene Fonseca and Carlo Mantegazza. Second order singular perturbation models for phase transitions. SIAM Journal on Mathematical Analysis, 31(5):1121–1143, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35683.

Fujigaki:2001:APN

[FM01] Yoshiko Fujigaki and Tetsuro Miyakawa. Asymptotic profiles of nonstationary incompressible Navier– Stokes flows in the whole space. SIAM Journal on Mathematical Analysis, 33(3):523–544, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36707.

Fleischmann:2004:SBM

[FM04] Klaus Fleischmann and Carl Mueller. Super-Brownian motion with extra birth at one point. SIAM Journal on Mathematical Analysis, 36(3):740– 772, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41947.

Forcadel:2009:ESM

[FM09] Nicolas Forcadel and Régis Monneau. Existence of solutions for a model describing the dynamics of junctions between dislocations. SIAM Journal on Mathematical Analysis, 40(6): 2517–2535, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fila:2005:IRA

[FMP05] Marek Fila, Hiroshi Matano, and Peter Polácik. Immediate regularization after blowup. SIAM Journal on Mathematical Analysis, 37(3):752– 776, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/61329.

Freddi:2009:AIR

[FMP09] Lorenzo Freddi, François Murat, and Roberto Paroni. Anisotropic inhomogeneous rectangular thin-walled beams. SIAM Journal on Mathematical Analysis, 40(5):1923–1951, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frehse:2003:ASF

[FMS03] Jens Frehse, Josef Málek, and Mark Steinhauer. On analysis of steady flows of fluids with shear-dependent viscosity based on the Lipschitz truncation method. SIAM Journal on Mathematical Analysis, 34(5):1064–1083, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41098.

Feireisl:2005:WSS

[FN05] Eduard Feireisl and Antonín Novotný. Weak sequential stability of the set of admissible variational solutions to the Navier–Stokes–Fourier system. SIAM Journal on Mathematical Analysis, 37(2):619– 650, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61458.

Fouassier:2006:HFA

[Fou06] Elise Fouassier. High frequency analysis of Helmholtz equations: Case of two point sources. SIAM Journal on Mathematical Analysis, 38(2): 617–636, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Filo:2002:FBP

[FP02] Ján Filo and Volker Pluschke. A free boundary problem in dermal drug delivery. SIAM Journal on Mathematical Analysis, 33(6):1430–1454, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38579.

Finch:2004:DFM

[FPR04] David Finch, Sarah K. Patch, and Rakesh. Determining a function from its mean values over a family of spheres. SIAM Journal on Mathematical Analysis, 35(5):1213–1240, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41781.

Fraysse:2007:GVM

[Fra07] A. Fraysse. Generic validity of the multifractal formalism. SIAM Journal on Mathematical Analysis, 39(2):593–607, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frid:2000:VSV

[FS00] Hermano Frid and Vladimir Shelukhin. Vanishing shear viscosity in the equations of compressible fluids for the flows with the cylinder symmetry. SIAM Journal on Mathematical Analysis, 31(5):1144–1156, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34394.

Frenod:2001:FLR

[FS01] Emmanuel Frénod and Eric Sonnendrücker. The finite Larmor radius approximation. SIAM Journal on Mathematical Analysis, 32(6):1227–1247, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36424.

Frid:2003:QLP

[FS03] Hermano Frid and Vladimir Shelukhin. A quasi-linear parabolic system for threephase capillary flow in porous media. SIAM Journal on Mathematical Analysis, 35(4): 1029–1041, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40216.

Frid:2005:IBV

[FS05] Hermano Frid and Vladimir Shelukhin. Initial boundary value problems for a quasilinear parabolic system in three-phase capillary flow in porous media. SIAM Journal on Mathematical Analysis, 36(5):1407–1425, 2005.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43533.

Frid:2009:HNP

[FS09] Hermano Frid and Jean Silva. Homogenization of nonlinear PDEs in the Fourier–Stieltjes algebras. SIAM Journal on Mathematical Analysis, 41(4): 1589–1620, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greer:2004:TWS

[GB04] J. B. Greer and A. L. Bertozzi. Traveling wave solutions of fourth order PDEs for image processing. SIAM Journal on Mathematical Analysis, 36(1):38–68, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        42737.

Gevirtz:2008:BBS

[Gev08] Julian Gevirtz. Boundary behavior of solutions of a class of genuinely nonlinear hyperbolic systems. SIAM Journal on Mathematical Analysis, 40 (4):1291–1336, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gilding:2000:LSE

[GG00] B. H. Gilding and J. Goncerzewicz. Localization of solutions of exterior domain problems for the porous media equation with radial symmetry. SIAM Journal on Mathematical Analysis, 31(4):862–893, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34450.

Ghoussoub:2007:PDE

[GG07] Nassif Ghoussoub and Yujin Guo. On the partial differential equations of electrostatic MEMS devices: Stationary case. SIAM Journal on Mathematical Analysis, 38(5): 1423–1449, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giga:2005:SSE

[GGH05] Mi-Ho Giga, Yoshikazu Giga, and Hidekata Hontani. Selfsimilar expanding solutions in a sector for a crystalline flow. SIAM Journal on Mathematical Analysis, 37(4):1207– 1226, January 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garay:2003:RPE

[GH03] Barnabas M. Garay and Josef Hofbauer. Robust permanence for ecological differential equations, minimax, and discretizations. SIAM Journal on Mathematical Analysis, 34(5):1007–1039, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39281.

Glitzky:2005:GER

[GH05] A. Glitzky and R. Hünlich. Global existence result for pair diffusion models. SIAM Journal on Mathematical Analysis, 36(4):1200–1225, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41759.

Grunrock:2008:LRL

[GH08] Axel Grünrock and Sebastian Herr. Low regularity local wellposedness of the derivative nonlinear Schrödinger equation with periodic initial data. SIAM Journal on Mathematical Analysis, 39(6):1890–1920, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gang:2009:GMT

[GH09] Xu Gang and Yin Huicheng. Global multidimensional transonic conic shock wave for the perturbed supersonic flow past a cone. SIAM Journal on Mathematical Analysis, 41(1): 178–218, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giacomini:2005:SEQ

[Gia05] Alessandro Giacomini. Size effects on quasi-static growth of cracks. SIAM Journal on Mathematical Analysis, 36(6): 1887–1928, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43936.

Garcke:2005:LSA

[GIK05] Harald Garcke, Kazuo Ito, and Yoshihito Kohsaka. Linearized stability analysis of stationary solutions for surface diffusion with boundary conditions. SIAM Journal on Mathematical Analysis, 36(4):1031–1056, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43793.

Garcke:2008:NSS

[GIK08] Harald Garcke, Kazuo Ito, and Yoshihito Kohsaka. Nonlinear stability of stationary solutions for surface diffusion with boundary conditions. SIAM Journal on Mathematical Analysis, 40(2):491–515, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gurevich:2009:PST

[GJS09] P. Gurevich, W. Jäger, and A. Skubachevskii. On periodicity of solutions for thermocontrol problems with hysteresistype switches. SIAM Journal on Mathematical Analysis, 41 (2):733–752, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gualdani:2006:NFO

[GJT06] Maria Pia Gualdani, Ansgar Jüngel, and Giuseppe Toscani. A nonlinear fourthorder parabolic equation with nonhomogeneous boundary conditions. SIAM Journal on Mathematical Analysis, 37(6): 1761–1779, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galiano:2003:PCD

[GJV03] Gonzalo Galiano, Ansgar Jüngel, and Julián Velasco. A parabolic cross-diffusion system for granular materials. SIAM Journal on Mathematical Analysis, 35(3):561–578, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40938.

Guo:2008:SSI

[GJX08] Zhenhua Guo, Quansen Jiu, and Zhouping Xin. Spherically symmetric isentropic compressible flows with densitydependent viscosity coefficients. SIAM Journal on Mathematical Analysis, 39(5):1402– 1427, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gundy:2000:STL

[GK00] Richard F. Gundy and Kazaros Kazarian. Stopping times and local convergence for spline wavelet expansions. SIAM Journal on Mathematical Analysis, 31(3):561–573, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32739.

Gyongy:2005:ASM

[GK05] István Gyöngy and Nicolai Krylov. An accelerated splitting-up method for parabolic equations. SIAM Journal on Mathematical Analysis, 37(4):1070–1097, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43790.

Grothaus:2008:ERC

[GK08] Martin Grothaus and Axel Klar. Ergodicity and rate of convergence for a nonsectorial fiber lay-down process. SIAM Journal on Mathematical Analysis, 40(3):968–983, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2007:OSM

[GL07] Kanghui Guo and Demetrio Labate. Optimally sparse multidimensional representation using shearlets. SIAM Journal on Mathematical Analysis, 39(1):298–318, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giacomini:2008:QSE

[GL08] Alessandro Giacomini and Luca Lussardi. Quasistatic evolution for a model in strain gradient plasticity. SIAM Journal on Mathematical Analysis, 40(3):1201–1245, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gousseau:2001:NIB

[GM01] Yann Gousseau and Jean-Michel Morel. Are natural images of bounded variation? SIAM Journal on Mathematical Analysis, 33(3):634–648, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37115.

Garroni:2005:LPF

[GM05] A. Garroni and S. Müller. Γ-Limit of a phase-field model of dislocations. SIAM Journal on Mathematical Analysis, 36(6):1943–1964, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43768.

Guerra:2009:BLI

[GMS09] Graziano Guerra, Francesca Marcellini, and Veronika Schleper. Balance laws with integrable unbounded sources. SIAM Journal on Mathematical Analysis, 41(3):1164–1189, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goddard:2009:RCC

[God09] Benjamin D. Goddard. Rate of convergence of the configuration interaction model for the helium ground state. SIAM Journal on Mathematical Analysis, 41(1):77–116, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gomes:2003:PTV

[Gom03] Diogo Aguiar Gomes. Perturbation theory for viscosity solutions of Hamilton– Jacobi equations and stability of Aubry–Mather sets. SIAM Journal on Mathematical Analysis, 35(1):135–147, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40596.

Gordon:2002:FBP

[Gor02] Michael S. Gordon. A free boundary problem for a hypoplastic model of plane shear waves in a fully saturated granular material. SIAM Journal on Mathematical Analysis, 34 (3):527–554, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39012.

Goudon:2004:HTE

[GP04] Thierry Goudon and Frédéric Poupaud. Homogenization of transport equations: Weak mean field approximation. SIAM Journal on Mathematical Analysis, 36(3):856–881, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41503.

Gunther:2005:HST

[GP05] Matthias Günther and Georg Prokert. On a Hele–Shaw-Type domain evolution with convected surface energy density. SIAM Journal on Mathematical Analysis, 37(2):372– 410, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44484.

Gunther:2006:HST

[GP06] Matthias Günther and Georg Prokert. On a Hele–Shaw type domain evolution with convected surface energy density: The third-order problem. SIAM Journal on Mathematical Analysis, 38(4):1154– 1185, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galaktionov:2000:AFD

[GPV00] Victor A. Galaktionov, Lambertus A. Peletier, and Juan L. Vazquez. Asymptotics of the fast-diffusion equation with critical exponent. SIAM Journal on Mathematical Analysis, 31(5):1157–1174, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32845.

Gallay:2000:SVS

[GR00] Thierry Gallay and Geneviève Raugel. Scaling variables and stability of hyperbolic fronts. SIAM Journal on Mathematical Analysis, 32(1):1–29, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35133.

Gourley:2003:CTF

[GR03] Stephen A. Gourley and Shigui Ruan. Convergence and travelling fronts in functional differential equations with nonlocal terms: A competition model. SIAM Journal on Mathematical Analysis, 35(3): 806–822, January 2003. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grandmont:2008:EWS

[Gra08] Céline Grandmont. Existence of weak solutions for the unsteady interaction of a viscous fluid with an elastic plate. SIAM Journal on Mathematical Analysis, 40(2):716–737, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grabovsky:2009:ERE

[Gra09] Yury Grabovsky. Exact relations for effective conductivity of fiber-reinforced conducting composites with the Hall effect via a general theory. SIAM Journal on Mathematical Analysis, 41(3):973–1024, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grun:2003:DSU

[Grü03]  Günther Grün.  Droplet

spreading under weak slippage: A basic result on finite speed of propagation. SIAM Journal on Mathematical Analysis, 34(4):992–1006, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40329.

Gallagher:2005:PGD

[GSR05] Isabelle Gallagher and Laure Saint-Raymond. On pressureless gases driven by a strong inhomogeneous magnetic field. SIAM Journal on Mathematical Analysis, 36(4):1159–1176, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43554.

Guo:2004:DPD

[GSZ04] Yan Guo, Chi-Wang Shu, and Tie Zhou. The dynamics of a plane diode. SIAM Journal on Mathematical Analysis, 35(6):1617–1635, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42113.

Guidotti:2007:CFB

[Gui07] Patrick Guidotti. A class of free boundary problems with onset of a new phase. SIAM Journal on Mathematical Analysis, 38(6):1981–2004, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gangbo:2000:UEC

[GV00] Wilfrid Gangbo and Roberto Van der Putten. Uniqueness of equilibrium configurations in solid crystals. SIAM Journal on Mathematical Analysis, 32(3):465–492, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35668.

Gerard-Varet:2005:OSI

[GV05] David Gérard-Varet. Oscillating solutions of incompressible magnetohydrodynamics and dynamo effect. SIAM Journal on Mathematical Analysis, 37(3):815–840, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44460.

Garcke:2006:SST

[GW06] Harald Garcke and Sandra Wieland. Surfactant spreading on thin viscous films: Nonnegative solutions of a coupled degenerate system. SIAM Journal on Mathematical Analysis, 37(6):2025–2048, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Groves:2007:SDM

[GW07] M. D. Groves and E. Wahlén. Spatial dynamics methods for solitary gravity-capillary water waves with an arbitrary distribution of vorticity. SIAM Journal on Mathematical Analysis, 39(3):932–964, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2009:FON

[GW09] Zongming Guo and Juncheng Wei. On a fourth order nonlinear elliptic equation with negative exponent. SIAM Journal on Mathematical Analysis, 40 (5):2034–2054, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grasselli:2009:CEP

[GWZ09] Maurizio Grasselli, Hao Wu, and Songmu Zheng. Convergence to equilibrium for parabolic-hyperbolic time-dependent Ginzburg–Landau–Maxwell equations. SIAM Journal on Mathematical Analysis, 40(5):2007– 2033, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gan:2000:RPH

[GZ00] Shaobo Gan and Meirong Zhang. Resonance pockets of Hill’s equations with twostep potentials. SIAM Journal on Mathematical Analysis, 32(3):651–664, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35684.

Guo:2009:WBP

[GZ09] Zhengguang Guo and Yong Zhou. Wave breaking and persistence properties for the dispersive rod equation. SIAM Journal on Mathematical Analysis, 40(6):2567–2580, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ha:2001:SSC

[Ha01] Seung-Yeal Ha. L^{1} stability for systems of conservation laws with a nonresonant moving source. SIAM Journal on Mathematical Analysis, 33(2):411–439, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37304.

Hansen:2001:LBS

[Han01] Olaf Hansen. The local behavior of the solution of the radiosity equation at the vertices of polyhedral domains in R^{3}. SIAM Journal on Mathematical Analysis, 33(3):718– 750, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37810.

Han:2008:RFC

[Han08] Bin Han. Refinable functions and cascade algorithms in weighted spaces with Hölder continuous masks. SIAM Journal on Mathematical Analysis, 40(1):70–102, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hattori:2000:ERA

[Hat00] Harumi Hattori. The entropy rate admissibility criterion and the entropy condition for a phase transition problem: The isothermal case. SIAM Journal on Mathematical Analysis, 31(4):791–820, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34122.

Hattori:2003:ELT

[Hat03] Harumi Hattori. The existence and large time behavior of solutions to a system related to a phase transition problem. SIAM Journal on Mathematical Analysis, 34(4):774– 804, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39137.

Hahner:2001:NSE

[HH01] Peter Hähner and Thorsten Hohage. New stability estimates for the inverse acoustic inhomogeneous medium problem and applications. SIAM Journal on Mathematical Analysis, 33(3):670–685, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38356.

Hanke:2009:IBP

[HHR09] Martin Hanke, Nuutti Hyvönen, and Stefanie Reusswig. An inverse backscatter problem for electric impedance tomography. SIAM Journal on Mathematical Analysis, 41(5):1948– 1966, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [HHR11].

Hanke:2011:EIB

[HHR11] Martin Hanke, Nuutti Hyvönen, and Stefanie Reusswig. Erratum: An inverse backscatter problem for electric impedance tomography. SIAM Journal on Mathematical Analysis, 43(3): 1495–1497, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/1/
sjmaah/v43/i3/p1495_s1. See

[HHR09].

Hanada:2004:EOM

[HIN04] Takao Hanada, Naoyuki Ishimura, and Masaaki Nakamura. On the Eguchi–Oki–Matsumura equation for phase separation in one space dimension. SIAM Journal on Mathematical Analysis, 36(2):463–478, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40012.

Healey:2000:GCH

[HK00a] Timothy J. Healey and Hansjörg Kielhöfer. Global continuation via higher-gradient regularization and singular limits in forced one-dimensional phase transitions. SIAM Journal on Mathematical Analysis, 31(6):1307–1331, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34006.

Hoshiga:2000:GSA

[HK00b] Akira Hoshiga and Hideo Kubo. Global small amplitude solutions of nonlinear hyperbolic systems with a critical exponent under the null condition. SIAM Journal on Mathematical Analysis, 31(3):486– 513, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32606.

Hmidi:2006:RBC

[HK06] Taoufik Hmidi and Sahbi Keraani. Remarks on the blowup for the L^{2}-critical nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 38(4):1035–1047, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrell:2001:POW

[HKK01] Evans M. Harrell II, Pawel Kröger, and Kazuhiro Kurata. On the placement of an obstacle or a well so as to optimize the fundamental eigenvalue. SIAM Journal on Mathematical Analysis, 33(1):240– 259, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35757.

Hastings:2008:DMT

[HKM08] Stuart Hastings, David Kinderlehrer, and J. Bryce McLeod. Diffusion mediated transport in multiple state systems. SIAM Journal on Mathematical Analysis, 39(4):1208–1230, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herty:2007:ESS

[HKP07] M. Herty, A. Klar, and B. Piccoli. Existence of solutions for supply chain models based on partial differential equations. SIAM Journal on Mathematical Analysis, 39(1):160–173, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hwang:2005:GSN

[HKS05] Hyung Ju Hwang, Kyungkeun Kang, and Angela Stevens. Global solutions of nonlinear transport equations for chemosensitive movement. SIAM Journal on Mathematical Analysis, 36(4):1177–1199, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43188. See erra-

tum [HKS07].

Hwang:2007:EGS

[HKS07] Hyung Ju Hwang, Kyungkeun Kang, and Angela Stevens. Erratum: “Global Solutions of Nonlinear Transport Equations for Chemosensitive Movement” [SIAM J. Math. Anal. 36 (2005), pp. 1177–1199]. SIAM Journal on Mathematical Analysis, 39(3):1018–1021, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [HKS05].

Hayes:2000:NSK

[HL00] Brian T. Hayes and Philippe G.

LeFloch. Nonclassical shocks and kinetic relations: Strictly hyperbolic systems. SIAM Journal on Mathematical Analysis, 31(5):941–991, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31982.

Hwang:2002:DRPa

[HL02a] Woonjae Hwang and W. Brent Lindquist. The 2-dimensional Riemann problem for a 2 × 2 hyperbolic conservation law I. Isotropic media. SIAM Journal on Mathematical Analysis, 34(2):341–358, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39663.

Hwang:2002:DRPb

[HL02b] Woonjae Hwang and W. Brent Lindquist. The 2-dimensional Riemann problem for a 2 × 2 hyperbolic conservation law II. Anisotropic media. SIAM Journal on Mathematical Analysis, 34(2):359–384, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39664.

Hou:2006:GWP

[HL06] Thomas Y. Hou and Congming Li. On global well-posedness of the Lagrangian averaged Euler equations. SIAM Journal on Mathematical Analysis, 38(3): 782–794, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hutson:2003:CSN

[HLMP03] V. Hutson, Y. Lou, K. Mischaikow, and P. Polácik. Competing species near a degenerate limit. SIAM Journal on Mathematical Analysis, 35 (2):453–491, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        40218.

Harris:2008:ETA

[HLTT08] Lisa Harris, Jani Lukkarinen, Stefan Teufel, and Florian Theil. Energy transport by acoustic modes of harmonic lattices. SIAM Journal on Mathematical Analysis, 40(4): 1392–1418, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hastings:2009:EAM

[HM09a] S. P. Hastings and J. B.

McLeod. An elementary approach to a model problem of Lagerstrom. SIAM Journal on Mathematical Analysis, 40(6): 2421–2436, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoffman:2009:ECP

[HM09b] Kathleen A. Hoffman and Robert S. Manning. An extended conjugate point theory with application to the stability of planar buckling of an elastic rod subject to a repulsive self-potential. SIAM Journal on Mathematical Analysis, 41(2):465–494, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2003:GSF

[HMS03] Feimin Huang, Akitaka Matsumura, and Xiaoding Shi. A gas-solid free boundary problem for a compressible viscous gas. SIAM Journal on Mathematical Analysis, 34(6):1331–1355, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40373.

Huang:2004:VSW

[HMS04] Feimin Huang, Akitaka Matsumura, and Xiaoding Shi.

Viscous shock wave to a gassolid free boundary problem for compressible gas. SIAM Journal on Mathematical Analysis, 36(2):498–522, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40943.

Hu:2005:ADN

[HN05] Bei Hu and David P. Nicholls. Analyticity of Dirichlet–Neumann operators on Hölder and Lipschitz domains. SIAM Journal on Mathematical Analysis, 37(1):302–320, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44481.

Hayashi:2008:SED

[HNS08] Nakao Hayashi, Pavel I. Naumkin, and Hideaki Sunagawa. On the Schrödinger equation with dissipative nonlinearities of derivative type. SIAM Journal on Mathematical Analysis, 40(1):278–291, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hecklin:2006:SCS

[HNZ06] G. Hecklin, G. Nürnberger, and F. Zeilfelder. The structure of C1 spline spaces on Freudenthal partitions. SIAM Journal on Mathematical Analysis, 38(2):347–367, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoff:2006:UWS

[Hof06] David Hoff. Uniqueness of weak solutions of the Navier– Stokes equations of multidimensional, compressible flow. SIAM Journal on Mathematical Analysis, 37(6):1742–1760, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hornung:2008:CRT

[Hor08]  Peter Hornung.  A Γ-

convergence result for thin Martensitic films in linearized elasticity. SIAM Journal on Mathematical Analysis, 40(1): 186–214, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoff:2009:IBT

[HP09] David Hoff and Misha Perepelitsa. Instantaneous boundary tangency and cusp formation in two-dimensional fluid flow. SIAM Journal on Mathematical Analysis, 41(2):753– 780, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herty:2006:CCC

[HR06] M. Herty and M. Rascle. Coupling conditions for a class of second-order models for traffic flow. SIAM Journal on Mathematical Analysis, 38(2):

595–616, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hieber:2008:QPS

[HR08] Matthias Hieber and Joachim Rehberg. Quasilinear parabolic systems with mixed boundary conditions on nonsmooth domains. SIAM Journal on Mathematical Analysis, 40(1): 292–305, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hetzer:2002:UPC

[HS02] Georg Hetzer and Wenxian Shen. Uniform persistence, coexistence, and extinction in almost Periodic/Nonautonomous competition diffusion systems. SIAM Journal on Mathematical Analysis, 34(1):204–227, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39069.

Han:2006:WSS

[HS06] Bin Han and Zuowei Shen. Wavelets with short support. SIAM Journal on Mathematical Analysis, 38(2):530–556, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han:2008:CSS

[HS08] Bin Han and Zuowei Shen.

Compactly supported symmetric C^{∞} wavelets with spectral approximation order. SIAM Journal on Mathematical Analysis, 40(3):905–938, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hohage:2003:STHa

[HSZ03a] Thorsten Hohage, Frank Schmidt, and Lin Zschiedrich. Solving time-harmonic scattering problems based on the pole condition I: Theory. SIAM Journal on Mathematical Analysis, 35(1):183–210, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40647.

Hohage:2003:STHb

[HSZ03b] Thorsten Hohage, Frank Schmidt, and Lin Zschiedrich. Solving time-harmonic scattering problems based on the pole condition II: Convergence of the PML method. SIAM Journal on Mathematical Analysis, 35(3):547–560, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40648.

Hillairet:2009:CTD

[HT09] Matthieu Hillairet and Takéo Takahashi. Collisions in three-dimensional fluid structure interaction problems. SIAM Journal on Mathematical Analysis, 40(6):2451–2477, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hur:2006:GBT

[Hur06] Vera Mikyoung Hur. Global bifurcation theory of deepwater waves with vorticity. SIAM Journal on Mathematical Analysis, 37(5):1482–1521, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2002:CVS

[HW02] Feimin Huang and Zhen Wang. Convergence of viscosity solutions for isothermal gas dynamics. SIAM Journal on Mathematical Analysis, 34(3): 595–610, 2002. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40581.

Huang:2003:NWN

[HW03] Lihong Huang and Jianhong Wu. Nonlinear waves in networks of neurons with delayed feedback: Pattern formation and continuation. SIAM Journal on Mathematical Analysis, 34(4):836–860, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38651.

Hu:2009:LMN

[HW09] Xianpeng Hu and Dehua Wang. Low Mach number limit of viscous compressible magnetohydrodynamic flows. SIAM Journal on Mathematical Analysis, 41(3):1272–1294, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hwang:2004:RVP

[Hwa04] Hyung Ju Hwang. Regularity for the Vlasov–Poisson system in a convex domain. SIAM Journal on Mathematical Analysis, 36(1):121–171, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42227.

Ha:2003:SSH

[HY03] Seung-Yeal Ha and Tong Yang.

L^{1} stability for systems of hyperbolic conservation laws with a resonant moving source. SIAM Journal on Mathematical Analysis, 34(5):1226–1251, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39798.

Hoff:2003:FDA

[HZ03] David Hoff and Mohammed Ziane. Finite-dimensional attractors and exponential attractors for the Navier–Stokes equations of compressible flow.

SIAM Journal on Mathematical Analysis, 34(5):1040–1063, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40088.

Hsu:2008:SST

[HZ08] Sze-Bi Hsu and Xiao-Qiang Zhao. Spreading speeds and traveling waves for nonmonotone integrodifference equations. SIAM Journal on Mathematical Analysis, 40(2):776– 789, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:2009:DSM

[HZ09] Lingbing He and Ping Zhang. L^{2} decay of solutions to a micro-macro model for polymeric fluids near equilibrium. SIAM Journal on Mathematical Analysis, 40(5):1905–1922, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iftimie:2002:URN

[Ift02] Dragos Iftimie. A uniqueness result for the Navier–Stokes equations with vanishing vertical viscosity. SIAM Journal on Mathematical Analysis, 33(6):1483–1493, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38212.

Ishii:2001:ASM

[II01] Hitoshi Ishii and Katsuyuki Ishii. An approximation scheme for motion by mean curvature with right-angle boundary condition. SIAM Journal on Mathematical Analysis, 33(2):369–389, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36810.

Ishibashi:2001:FNP

[IK01] Toshihiro Ishibashi and Shigeaki Koike. On fully nonlinear PDEs derived from variational problems of L^{p} norms. SIAM Journal on Mathematical Analysis, 33(3):545–569, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38000.

Ishii:2005:LSL

[IL05] Hitoshi Ishii and Paola Loreti.

Limits of solutions of p-Laplace equations as p goes to infinity and related variational problems. SIAM Journal on Mathematical Analysis, 37(2):411– 437, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43282.

Ishii:2001:MMW

[IM01] Hitoshi Ishii and Toshio Mikami. A mathematical model of the wearing process of a nonconvex stone. SIAM Journal on Mathematical Analysis, 33(4):860–876, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38495.

Ishii:2004:CGC

[IM04] Hitoshi Ishii and Toshio Mikami. Convexified Gauss curvature flow of sets: A stochastic approximation. SIAM Journal on Mathematical Analysis, 36(2):552–579, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42050.

Ingerman:2000:DCD

[Ing00] David V. Ingerman. Discrete and continuous Dirichlet-to-Neumann maps in the layered case. SIAM Journal on Mathematical Analysis, 31(6): 1214–1234, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32658.

Isakov:2001:UCT

[Isa01] Victor Isakov. On the uniqueness of the continuation for a thermoelasticity system. SIAM Journal on Mathematical Analysis, 33(3):509–522, 2001. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36650.

Ishii:2005:ORC

[Ish05] Katsuyuki Ishii. Optimal rate of convergence of the Bence– Merriman–Osher algorithm for motion by mean curvature. SIAM Journal on Mathematical Analysis, 37(3):841–866, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61862.

Imbert:2004:KFM

[IV04] C. Imbert and J. Vovelle. A kinetic formulation for multidimensional scalar conservation laws with boundary conditions and applications. SIAM Journal on Mathematical Analysis, 36(1):214–232, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42468.

Isakov:2008:IPD

[IWY08] Victor Isakov, Jenn-Nan Wang, and Masahiro Yamamoto. An inverse problem for a dynamical Lamé system with residual stress. SIAM Journal on Mathematical Analysis, 39(4): 1328–1343, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jenssen:2000:BSC

[Jen00] Helge Kristian Jenssen. Blowup for systems of conservation laws. SIAM Journal on Mathematical Analysis, 31(4):894– 908, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35233.

Jia:2000:DSN

[JJS00] Rong-Qing Jia, Qingtang Jiang, and Zuowei Shen. Distributional solutions of nonhomogeneous discrete and continuous refinement equations. SIAM Journal on Mathematical Analysis, 32(2):420–434, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35088.

Juutinen:2001:EVS

[JLM01] Petri Juutinen, Peter Lindqvist, and Juan J. Manfredi. On the equivalence of viscosity solutions and weak solutions for a quasi-linear equation. SIAM Journal on Mathematical Analysis, 33(3):699–717, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37217.

Johnson:2006:EEL

[JLP06] E. R. Johnson, Michael Levitin, and Leonid Parnovski.

Existence of eigenvalues of a linear operator pencil in a curved waveguide — localized shelf waves on a curved coast. SIAM Journal on Mathematical Analysis, 37(5):1465– 1481, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jungel:2008:DLS

[JM08] Ansgar Jüngel and Daniel Matthes. The Derrida– Lebowitz–Speer–Spohn equation: Existence, nonuniqueness, and decay rates of the solutions. SIAM Journal on Mathematical Analysis, 39(6): 1996–2015, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jungel:2009:SON

[JM09] Ansgar Jüngel and Josipa-Pina Milišić. A sixth-order nonlinear parabolic equation for quantum systems. SIAM Journal on Mathematical Analysis, 41(4):1472–1490, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Janas:2004:IJM

[JN04] Jan Janas and Serguei Naboko.

Infinite Jacobi matrices with unbounded entries: Asymptotics of eigenvalues and the transformation operator approach. SIAM Journal on Mathematical Analysis, 36(2): 643–658, 2004. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        40607.

Jiang:2006:VVL

[JNS06] Song Jiang, Guoxi Ni, and Wenjun Sun. Vanishing viscosity limit to rarefaction waves for the Navier–Stokes equations of one-dimensional compressible heat-conducting fluids. SIAM Journal on Mathematical Analysis, 38(2):368– 384, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jochmann:2000:LTA

[Joc00]  Frank Jochmann.  Long

time asymptotics of solutions to the anharmonic oscillator model from nonlinear optics. SIAM Journal on Mathematical Analysis, 32(4):887–915, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36093.

Jochmann:2002:ESQ

[Joc02] Frank Jochmann. Existence of solutions and a quasistationary limit for a hyperbolic system describing ferromagnetism. SIAM Journal on Mathematical Analysis, 34 (2):315–340, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.

org/sam-bin/dbq/article/
39229.

Jochmann:2005:ABE

[Joc05] Frank Jochmann. Asymptotic behavior of the electromagnetic field for a micromagnetism equation without exchange energy. SIAM Journal on Mathematical Analysis, 37(1):276–290, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44332.

Johnson:2003:OAW

[Joh03] Brody Dylan Johnson. On the oversampling of affine wavelet frames. SIAM Journal on Mathematical Analysis, 35(3): 623–638, January 2003. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johnson:2009:NSP

[Joh09] Mathew A. Johnson. Nonlinear stability of periodic traveling wave solutions of the generalized Korteweg–de Vries equation. SIAM Journal on Mathematical Analysis, 41(5):1921– 1947, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jungel:2000:GNS

[JP00] Ansgar Jüngel and René Pinnau. Global nonnegative solutions of a nonlinear fourthorder parabolic equation for quantum systems. SIAM Journal on Mathematical Analysis, 32(4):760–777, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36026.

Jimbo:2002:NPC

[JS02] Shuichi Jimbo and Peter Sternberg. Nonexistence of permanent currents in convex planar samples. SIAM Journal on Mathematical Analysis, 33(6):1379–1392, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38602.

Jabin:2009:KDP

[JT09] Pierre-Emmanuel Jabin and Athanasios E. Tzavaras. Kinetic decomposition for periodic homogenization problems. SIAM Journal on Mathematical Analysis, 41(1):360–390, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jian:2007:CEM

[JW07] Huai-Yu Jian and Xu-Jia Wang. Continuity estimates for the Monge–Ampère equation. SIAM Journal on Mathematical Analysis, 39(2):608– 626, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jimenez:2007:WNH

[JWW07] David Jimenez, Long Wang, and Yang Wang. White noise hypothesis for uniform quantization errors. SIAM Journal on Mathematical Analysis, 38 (6):2042–2056, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2009:BLN

[JZ09a] Song Jiang and Jianwen Zhang. Boundary layers for the Navier–Stokes equations of compressible heat-conducting flows with cylindrical symmetry. SIAM Journal on Mathematical Analysis, 41(1):237– 268, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jin:2009:SDN

[JZ09b] Yu Jin and Xiao-Qiang Zhao.

Spatial dynamics of a nonlocal periodic reaction-diffusion model with stage structure. SIAM Journal on Mathematical Analysis, 40(6):2496–2516, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koch:2000:SHB

[KA00] Herbert Koch and Stuart S.

Antman. Stability and Hopf bifurcation for fully nonlinear parabolic-hyperbolic equations. SIAM Journal on Mathematical Analysis, 32(2):360– 384, 2000. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33793.

Kagei:2001:IML

[Kag01]  Yoshiyuki Kagei.  Invari-

ant manifolds and long-time asymptotics for the Vlasov– Poisson–Fokker–Planck equation. SIAM Journal on Mathematical Analysis, 33(2):489– 507, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37136.

Kang:2004:RAS

[Kan04] Kyungkeun Kang. Regularity of axially symmetric flows in a half-space in three dimensions. SIAM Journal on Mathematical Analysis, 35(6):1636–1643, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41442.

Katriel:2005:ETW

[Kat05] Guy Katriel. Existence of travelling waves in discrete Sine-Gordon rings. SIAM Journal on Mathematical Analysis, 36(5):1434–1443, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44017.

Keinert:2001:RMA

[Kei01] Fritz Keinert. Raising multiwavelet approximation order through lifting. SIAM Journal on Mathematical Analysis, 32(5):1032–1049, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34950.

Kelliher:2006:NSE

[Kel06] James P. Kelliher. Navier– Stokes equations with Navier boundary conditions for a bounded domain in the plane. SIAM Journal on Mathematical Analysis, 38(1):210–232, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kyrychko:2006:DSS

[KGB06] Y. Kyrychko, S. A. Gourley, and M. V. Bartuccelli. Dynamics of a stage-structured population model on an isolated finite lattice. SIAM Journal on Mathematical Analysis, 37(5): 1688–1708, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2000:SHS

[Kim00] Jong Uhn Kim. On a stochastic hyperbolic system in linear elasticity. SIAM Journal on Mathematical Analysis, 32 (2):304–322, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        35037.

Kim:2002:SEE

[Kim02]  Jong Uhn Kim.  On the

stochastic Euler equations in a two-dimensional domain. SIAM Journal on Mathematical Analysis, 33(5):1211–1227, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38394.

Kim:2005:IMS

[Kim05] Jong Uhn Kim. Invariant measures for the stochastic von Karman plate equation. SIAM Journal on Mathematical Analysis, 36(5):1689–1703, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43885.

Kim:2006:BCN

[Kim06] Hyunseok Kim. A blow-up criterion for the nonhomogeneous incompressible Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 37(5):1417–1434, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2009:LFL

[Kim09] Namkwon Kim. Large friction limit and the inviscid limit of 2D Navier–Stokes equations under Navier friction condition. SIAM Journal on Mathematical Analysis, 41(4):1653– 1663, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See corrigendum [Kim13].

Kim:2013:CLF

[Kim13] Namkwon Kim. Corrigendum: Large Friction Limit and the Inviscid Limit of 2D Navier–Stokes Equations under Navier Friction Condition. SIAM Journal on Mathematical Analysis, 45(3):1992–1994, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Kim09].

Kang:2002:HCS

[KK02] Kyungkeun Kang and Seick Kim. On the Hölder continuity of solutions of a certain system related to Maxwell’s equations. SIAM Journal on Mathematical Analysis, 34(1):87– 100, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39334. See erra-

tum [KK05].

Kim:2004:SST

[KK04a] Kyeong-Hun Kim and N. V.

Krylov. On the Sobolev space theory of parabolic and elliptic equations in C^{1} domains. SIAM Journal on Mathematical Analysis, 36(2):618–642, 2004. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42114.

Kweon:2004:RSN

[KK04b] Jae Ryong Kweon and R. Bruce Kellogg. Regularity of solutions to the Navier–Stokes system for compressible flows on a polygon. SIAM Journal on Mathematical Analysis, 35(6):1451–1485, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41806.

Kang:2005:EHC

[KK05] Kyungkeun Kang and Seick Kim. Erratum: On the Hölder continuity of solutions of a certain system related to Maxwell’s equations. SIAM Journal on Mathematical Analysis, 36(5):1704–1705, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/61290.  See

[KK02].

Kim:2007:EDE

[KK07] Doyoon Kim and N. V. Krylov. Elliptic differential equations with coefficients measurable with respect to one variable and VMO with respect to the others. SIAM Journal on Mathematical Analysis, 39(2): 489–506, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaltenbacher:2008:IPN

[KK08a] Barbara Kaltenbacher and Michael V. Klibanov. An inverse problem for a nonlinear parabolic equation with applications in population dynamics and magnetics. SIAM Journal on Mathematical Analysis, 39 (6):1863–1889, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Katayama:2008:DET

[KK08b] Soichiro Katayama and Hideo Kubo. Decay estimates of a tangential derivative to the light cone for the wave equation and their application. SIAM Journal on Mathematical Analysis, 39(6):1851–1862, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kawashita:2000:HMI

[KKS00] Mishio Kawashita, Yaroslav V.

Kurylev, and Hideo Soga. Harmonic moments and an inverse problem for the heat equation. SIAM Journal on Mathematical Analysis, 32(3):522– 537, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35303.

Kirr:2008:SBB

[KKSW08] E. W. Kirr, P. G. Kevrekidis, E. Shlizerman, and M. I. Weinstein. Symmetry-breaking bifurcation in nonlinear Schrödinger/ Gross–Pitaevskii equations. SIAM Journal on Mathematical Analysis, 40(2):566–604, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2002:NPD

[KKSY02] Sungwhan Kim, Ohin Kwon, Jin Keun Seo, and Jeong-Rock Yoon. On a nonlinear partial differential equation arising in magnetic resonance electrical impedance tomography. SIAM Journal on Mathematical Analysis, 34(3):511– 526, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39135.

Kondo:2002:ZDD

[KL02] Cezar I. Kondo and Philippe G. LeFloch. Zero diffusiondispersion limits for scalar conservation laws. SIAM Journal on Mathematical Analysis, 33(6):1320–1329, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37426.

Kammerer:2008:PTG

[KL08] Clotilde Fermanian Kammerer and Caroline Lasser. Propagation through generic level crossings: A surface hopping semigroup. SIAM Journal on Mathematical Analysis, 40(1): 103–133, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Killip:2009:CMM

[KLVZ09] Rowan Killip, Dong Li, Monica Visan, and Xiaoyi Zhang. Characterization of minimalmass blowup solutions to the focusing mass-critical NLS. SIAM Journal on Mathematical Analysis, 41(1):219–236, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kappeler:2001:EPD

[KM01] T. Kappeler and B. Mityagin. Estimates for periodic and Dirichlet eigenvalues of the Schrödinger operator. SIAM Journal on Mathematical Analysis, 33(1):113– 152, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36575.

Kamotski:2007:EHL

[KMS07] Vladimir Kamotski, Karsten Matthies, and Valery P. Smyshlyaev. Exponential homogenization of linear second order elliptic PDEs with periodic coefficients. SIAM Journal on Mathematical Analysis, 38(5):1565–1587, ???? 2007.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Karch:2008:CSF

[KMX08] Grzegorz Karch, Changxing Miao, and Xiaojing Xu. On convergence of solutions of fractal Burgers equation toward rarefaction waves. SIAM Journal on Mathematical Analysis, 39(5):1536–1549, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kawashita:2000:PRE

[KN00] Mishio Kawashita and Gen Nakamura. The poles of the resolvent for the exterior Neumann problem of anisotropic elasticity. SIAM Journal on Mathematical Analysis, 31(4): 701–725, 2000. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        31486.

Kim:2009:HOA

[KN09] Yong Jung Kim and Wei-Ming Ni. Higher order approximations in the heat equation and the truncated moment problem. SIAM Journal on Mathematical Analysis, 40(6):2241– 2261, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kunzinger:2000:GAD

[KO00] Michael Kunzinger and Michael Oberguggenberger. Group analysis of differential equations and generalized functions. SIAM Journal on Mathematical Analysis, 31(6): 1192–1213, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
33450.

Kovarik:2008:WCS

[Kov08]  Hynek Kovařík.  Weakly

coupled Schrödinger operators on regular metric trees. SIAM Journal on Mathematical Analysis, 39(4):1135–1149, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kozek:2005:IOB

[KP05] W. Kozek and G. E. Pfander. Identification of operators with bandlimited symbols. SIAM Journal on Mathematical Analysis, 37(3):867– 888, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43343.

Koiso:2007:UTS

[KP07] Miyuki Koiso and Bennett Palmer. Uniqueness theorems for stable anisotropic capillary surfaces. SIAM Journal on Mathematical Analysis, 39(3): 721–741, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kostin:2008:KZK

[KP08] I. Kostin and G. Panasenko. Khokhlov–Zabolotskaya–Kuznetsovtype equation: Nonlinear acoustics in heterogeneous media. SIAM Journal on Mathematical Analysis, 40(2):699– 715, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kevrekidis:2009:ASS

[KPS09] P. G. Kevrekidis, D. E. Pelinovsky, and A. Stefanov. Asymptotic stability of small bound states in the discrete nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 41(5):2010– 2030, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kodama:2009:WED

[KPT09] Yuji Kodama, V. U. Pierce, and Fei-Ran Tian. On the Whitham equations for the defocusing complex modified KdV equation. SIAM Journal on Mathematical Analysis, 40 (5):1750–1782, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krisztin:2004:INC

[Kri04] Tibor Krisztin. Invariance and noninvariance of center manifolds of time- t maps with respect to the semiflow. SIAM Journal on Mathematical Analysis, 36(3):717–739, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41917.

Krejci:2009:BF

[KRS09] Pavel Krejčí, Elisabetta Rocca, and Jürgen Sprekels. A bottle in a freezer. SIAM Journal on Mathematical Analysis, 41 (5):1851–1873, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krylov:2001:HES

[Kry01] N. V. Krylov. The heat equation in L_{q}((0, T), L_{p})-spaces with weights. SIAM Journal on Mathematical Analysis, 32(5):1117–1141, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37203.

Krylov:2003:BTR

[Kry03] N. V. Krylov. Brownian trajectory is a regular lateral boundary for the heat equation. SIAM Journal on Mathematical Analysis, 34(5):1167–1182, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40298.

Krylov:2009:DFS

[Kry09] N. V. Krylov. On divergence form SPDEs with VMO coefficients. SIAM Journal on Mathematical Analysis, 40(6): 2262–2285, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kunze:2000:RRC

[KS00] Markus Kunze and Herbert Spohn. Radiation reaction and center manifolds. SIAM Journal on Mathematical Analysis, 32(1):30–53, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35157.

Krupa:2001:EGS

[KS01] M. Krupa and P. Szmolyan. Extending geometric singular perturbation theory to nonhyperbolic points — fold and Canard points in two dimensions. SIAM Journal on Mathematical Analysis, 33(2):286– 314, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36091.

Kapitula:2002:EBN

[KS02a] Todd Kapitula and Björn Sandstede. Edge bifurcations for near integrable systems via Evans function techniques. SIAM Journal on Mathematical Analysis, 33(5):1117–1143, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37230.

Kuttler:2002:DCN

[KS02b] Kenneth L. Kuttler and Meir Shillor. Dynamic contact with normal compliance wear and discontinuous friction coefficient. SIAM Journal on Mathematical Analysis, 34(1):1–27, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39118.

Klaus:2003:EZS

[KS03a] M. Klaus and J. K. Shaw. On the eigenvalues of Zakharov– Shabat systems. SIAM Journal on Mathematical Analysis, 34(4):759–773, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40306.

Kopotun:2003:MAF

[KS03b] Kirill Kopotun and Alexei Shadrin. On k-monotone approximation by free knot splines. SIAM Journal on Mathematical Analysis, 34(4): 901–924, 2003. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35851.

Kusiak:2005:CSS

[KS05a] Steven Kusiak and John Sylvester. The convex scattering support in a background medium. SIAM Journal on Mathematical Analysis, 36(4):1142–1158, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43357.

Kutyniok:2005:WBG

[KS05b] Gitta Kutyniok and Thomas Strohmer. Wilson bases for general time-frequency lattices. SIAM Journal on Mathematical Analysis, 37(3):685– 711, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43723.

Kawohl:2006:DEA

[KS06a] B. Kawohl and G. Sweers. On the differential equation u_{xxxx} + u_{yyyy} = f for an anisotropic stiff material. SIAM Journal on Mathematical Analysis, 37(6):1828–1853, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kolev:2006:VPW

[KS06b] B. Kolev and D. H. Sattinger. Variational principles for water waves. SIAM Journal on Mathematical Analysis, 38(3): 906–920, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koch:2007:LSL

[KS07] Herbert Koch and Jean-Claude Saut. Local smoothing and local solvability for third order dispersive equations. SIAM Journal on Mathematical Analysis, 38(5):1528–1541, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kavallaris:2009:GRR

[KS09a] Nikos I. Kavallaris and Philippe Souplet. Grow-up rate and refined asymptotics for a two-dimensional Patlak– Keller–Segel model in a disk. SIAM Journal on Mathematical Analysis, 40(5):1852–1881, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kurzke:2009:ESL

[KS09b] Matthias Kurzke and Daniel Spirn. On the energy of superconductors in large and small domains. SIAM Journal on Mathematical Analysis, 40(5): 2077–2104, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kutyniok:2009:ADS

[KS09c] Gitta Kutyniok and Tomas Sauer. Adaptive directional subdivision schemes and shearlet multiresolution analysis. SIAM Journal on Mathematical Analysis, 41(4):1436–1471, ???? 2009. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Krejci:2002:PFM

[KSS02] Pavel Krejcí, Jürgen Sprekels, and Ulisse Stefanelli. Phasefield models with hysteresis in one-dimensional thermoviscoplasticity. SIAM Journal on Mathematical Analysis, 34 (2):409–434, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38760.

Kim:2001:DWM

[KT01] Yong Jung Kim and Athanasios E. Tzavaras. Diffusive N-waves and metastability in the Burgers equation. SIAM Journal on Mathematical Analysis, 33(3):607– 633, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38051.

Kaiser:2005:GEF

[KTvW05] R. Kaiser, A. Tilgner, and W. von Wahl. A generalized energy functional for plane Couette flow. SIAM Journal on Mathematical Analysis, 37 (2):438–454, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44260.

Kurth:2004:EIM

[Kur04]  Michael Kurth.  On the

existence of infinitely many modes of a nonlocal nonlinear Schrödinger equation related to dispersion-managed solitons. SIAM Journal on Mathematical Analysis, 36(3):967– 985, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43153.

Kirr:2001:PEH

[KW01] E. Kirr and M. I. Weinstein.

Parametrically excited Hamiltonian partial differential equations. SIAM Journal on Mathematical Analysis, 33(1):16– 52, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36345.

Kang:2003:BDC

[KY03a] Hyeonbae Kang and KiHyun Yun. Boundary determination of conductivities and Riemannian metrics via local Dirichlet-to-Neumann operator. SIAM Journal on Mathematical Analysis, 34(3):719– 735, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39504.

Kim:2003:UIS

[KY03b] Sungwhan Kim and Masahiro Yamamoto. Uniqueness in identification of the support of a source term in an elliptic equation. SIAM Journal on Mathematical Analysis, 35 (1):148–159, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41270.

Khasminskii:2004:APA

[KY04] R. Z. Khasminskii and G. Yin. On averaging principles: An asymptotic expansion approach. SIAM Journal on Mathematical Analysis, 35(6): 1534–1560, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40397.

Kyriazis:2004:MCA

[Kyr04]  George Kyriazis.  Mul-

tilevel characterizations of anisotropic function spaces. SIAM Journal on Mathematical Analysis, 36(2):441–462, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42568.

Laforest:2004:PEE

[Laf04] M. Laforest. A posteriori error estimate for front-tracking: Systems of conservation laws. SIAM Journal on Mathematical Analysis, 35(5):1347–1370, 2004. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41687.

Lu:2008:SDT

[LAJ08] Wenlian Lu, Fatihcan M. Atay, and Jürgen Jost. Synchronization of discrete-time dynamical networks with time-varying couplings. SIAM Journal on Mathematical Analysis, 39(4): 1231–1259, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurencot:2002:LSW

[Lau02]  Philippe Laurençot.  The

Lifshitz–Slyozov–Wagner equation with conserved total volume. SIAM Journal on Mathematical Analysis, 34(2):257– 272, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38747.

Layton:2007:BEH

[Lay07] William Layton. Bounds on energy and helicity dissipation rates of approximate deconvolution models of turbulence. SIAM Journal on Mathematical Analysis, 39(3):916–931, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lombardo:2003:WPB

[LCS03] Maria Carmela Lombardo, Marco Cannone, and Marco Sammartino. Well-posedness of the boundary layer equations. SIAM Journal on Mathematical Analysis, 35(4): 987–1004, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41205.

Le:2005:RSC

[Le05] Dung Le. Regularity of solutions to a class of cross diffusion systems. SIAM Journal on Mathematical Analysis, 36(6):1929–1942, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42835.

Lee:2009:VFM

[Lee09]  S. L. Lee.  Vector fields

for mean value coordinates. SIAM Journal on Mathematical Analysis, 40(6):2437–2450, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lenells:2008:HSE

[Len08] Jonatan Lenells. The Hunter– Saxton equation: A geometric approach. SIAM Journal on Mathematical Analysis, 40 (1):266–277, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lewicka:2001:SCP

[Lew01] Marta Lewicka. Stability conditions for patterns of noninteracting large shock waves. SIAM Journal on Mathematical Analysis, 32(5):1094–1116, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36750.

Lewicka:2005:LFS

[Lew05a] Marta Lewicka. Lyapunov functional for solutions of systems of conservation laws containing a strong rarefaction. SIAM Journal on Mathematical Analysis, 36(5):1371–1399, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42950.

Lewicka:2005:SCS

[Lew05b]  Marta Lewicka.  Stability

conditions for strong rarefaction waves. SIAM Journal on Mathematical Analysis, 36(4):1346–1369, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42951.

Lilli:2007:SPS

[LHK07] M. Lilli, T. J. Healey, and H. Kielhöfer. Singular perturbation as a selection criterion for young-measure solutions. SIAM Journal on Mathematical Analysis, 39(1):195– 209, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2005:IPM

[Li05] Shumin Li. An inverse problem for Maxwell’s equations in biisotropic media. SIAM Journal on Mathematical Analysis, 37(4):1027–1043, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/44366.

Li:2006:VPU

[Li06] Bo Li. Variational properties of unbounded order parameters. SIAM Journal on Mathematical Analysis, 38(1):16–36, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2008:STW

[Li08] Tong Li. Stability of traveling waves in quasi-linear hyperbolic systems with relaxation and diffusion. SIAM Journal on Mathematical Analysis, 40 (3):1058–1075, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2009:MEF

[Li09] Bo Li. Minimization of electrostatic free energy and the Poisson–Boltzmann equation for molecular solvation with implicit solvent. SIAM Journal on Mathematical Analysis, 40 (6):2536–2566, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Li09].

Li:2011:EME

[Li11] Bo Li. Erratum: “Minimization of Electrostatic Free Energy and the Poisson– Boltzmann Equation for Molecular Solvation with Implicit Solvent”. SIAM Journal on Mathematical Analysis, 43(6): 2776–2777, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/1/
sjmaah/v43/i6/p2776_s1. See

[Li09].

Lieberman:2005:BSS

[Lie05] Gary M. Lieberman. Bounds for the steady-state Sel’kov model for arbitrary p in any number of dimensions. SIAM Journal on Mathematical Analysis, 36(5):1400–1406, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43651.

Lilli:2007:CSN

[Lil07] Markus Lilli. Classical solutions for nonelliptic Euler– Lagrange equations via continuation. SIAM Journal on Mathematical Analysis, 39(3): 801–818, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lim:2002:TDT

[Lim02] H. Lim. Time discretization of transition layer dynamics in one-dimensional viscoelastic systems. SIAM Journal on Mathematical Analysis, 34 (3):573–594, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        39877.

Lin:2003:ISI

[Lin03] Zhiwu Lin. Instability of some ideal plane flows. SIAM Journal on Mathematical Analysis, 35(2):318–356, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        40626.

Lipton:2001:OIG

[Lip01] Robert Lipton. Optimal inequalities for gradients of solutions of elliptic equations occurring in two-phase heat conductors. SIAM Journal on Mathematical Analysis, 32(5):1081–1093, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36662.

Lipton:2006:HFC

[Lip06]  Robert Lipton.  Homoge-

nization and field concentrations in heterogeneous media. SIAM Journal on Mathematical Analysis, 38(4):1048– 1059, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liefvendahl:2001:SVS

[LK01] M. Liefvendahl and G. Kreiss.

Stability of viscous shock waves for problems with nonsymmetric viscosity matrices. SIAM Journal on Mathematical Analysis, 33(4):913–929, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37494.

Lin:2000:VSW

[LL00] FangHua Lin and Tai-Chia Lin. Vortex state of dwave superconductors in the Ginzburg–Landau energy. SIAM Journal on Mathematical Analysis, 32(3):493–503, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35352.

Lin:2003:VWS

[LL03] FangHua Lin and Tai-Chia Lin. Vortices in p-wave superconductivity. SIAM Journal on Mathematical Analysis, 34(5):1105–1127, 2003.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39582.

Lafortune:2005:SSL

[LL05] S. Lafortune and J. Lega. Spectral stability of local deformations of an elastic rod: Hamiltonian formalism. SIAM Journal on Mathematical Analysis, 36(6):1726–1741, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43935.

Levandosky:2006:SSW

[LL06] Steve Levandosky and Yue Liu. Stability of solitary waves of a generalized Ostrovsky equation. SIAM Journal on Mathematical Analysis, 38(3):985– 1011, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lan:2004:PVS

[LLLY04] Chiu-Ya Lan, Huey-Er Lin, Tai-Ping Liu, and Shih-Hsien Yu. Propagation of viscous shock waves away from the boundary. SIAM Journal on Mathematical Analysis, 36 (2):580–617, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
42815.

Lachowicz:2003:OHS

[LLW03] Miroslaw Lachowicz, Philippe Laurençot, and Dariusz Wrzosek. On the Oort–Hulst– Safronov coagulation equation and its relation to the Smoluchowski equation. SIAM Journal on Mathematical Analysis, 34(6):1399–1421, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41447.

Lei:2003:TCL

[LLYZ03] Jinzhi Lei, Xiong Li, Ping Yan, and Meirong Zhang. Twist character of the least amplitude periodic solution of the forced pendulum. SIAM Journal on Mathematical Analysis, 35(4):844–867, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41037.

Li:2008:UPB

[LM08] Congming Li and Li Ma.

Uniqueness of positive bound states to Schrödinger systems with critical exponents. SIAM Journal on Mathematical Analysis, 40(3):1049–1057, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lacave:2009:UVW

[LM09] Christophe Lacave and Evelyne Miot. Uniqueness for the vortex-wave system when the vorticity is constant near the point vortex. SIAM Journal on Mathematical Analysis, 41(3): 1138–1163, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lemou:2008:SLG

[LMR08] Mohammed Lemou, Florian Méhats, and Pierre Raphaël. Structure of the linearized gravitational Vlasov–Poisson system close to a polytropic ground state. SIAM Journal on Mathematical Analysis, 39 (6):1711–1739, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Loeper:2006:FNV

[Loe06] G. Loeper. A fully nonlinear version of the incompressible Euler equations: The semigeostrophic system. SIAM Journal on Mathematical Analysis, 38(3):795–823, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lopez:2000:AES

[Lóp00]  José L. López.  Asymp-

totic expansions of symmetric standard elliptic integrals. SIAM Journal on Mathematical Analysis, 31(4):754–775, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35117.

Lorenzi:2000:OSE

[Lor00]  Luca Lorenzi.  Optimal

Schauder estimates for parabolic problems with data measurable with respect to time. SIAM Journal on Mathematical Analysis, 32(3):588–615, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34284.

Lou:2006:PRW

[Lou06] Bendong Lou. Periodic rotating waves in an undulating annulus and their homogenization limit. SIAM Journal on Mathematical Analysis, 38(3): 693–716, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2000:MPM

[LP00] Yi A. Li and Keith Promislow. The mechanism of the polarizational mode instability in birefringent fiber optics. SIAM Journal on Mathematical Analysis, 31(6):1351–1373, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34996.

Lim:2002:IVP

[LP02] Wee Keong Lim and Gustavo Ponce. On the initial value problem for the one dimensional quasi-linear Schrödinger equations. SIAM Journal on Mathematical Analysis, 34 (2):435–459, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39952.

Lin:2007:MFI

[LP07] Fanghua Lin and Xing-Bin Pan. Magnetic field-induced instabilities in liquid crystals. SIAM Journal on Mathematical Analysis, 38(5):1588–1612, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leonori:2008:BBB

[LP08] Tommaso Leonori and Alessio Porretta. The boundary behavior of blow-up solutions related to a stochastic control problem with state constraint. SIAM Journal on Mathematical Analysis, 39(4):1295–1327, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Linares:2009:WPT

[LP09] Felipe Linares and Ademir Pastor. Well-posedness for the two-dimensional modified Zakharov–Kuznetsov equation. SIAM Journal on Mathematical Analysis, 41(4):1323–1339, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lototsky:2009:SPD

[LR09] Sergey V. Lototsky and Boris L. Rozovskii. Stochastic partial differential equations driven by purely spatial noise. SIAM Journal on Mathematical Analysis, 41(4):1295–1322, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurent:2006:PBH

[LRR06] Thomas Laurent, Brian Rider, and Michael Reed. Parabolic behavior of a hyperbolic delay equation. SIAM Journal on Mathematical Analysis, 38 (1):1–15, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Langa:2009:PAS

[LRRBS09] José A. Langa, James C.

Robinson, Aníbal Rodríguez-Bernal, and Antonio Suárez. Permanence and asymptotically stable complete trajectories for nonautonomous Lotka– Volterra models with diffusion. SIAM Journal on Mathematical Analysis, 40(6):2179–2216, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lee:2000:ITP

[LS00] June-Yub Lee and Jin Keun Seo. Identification of twophase free boundary arising in plasma physics. SIAM Journal on Mathematical Analysis, 31(6):1295–1306, 2000.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35118.

Lombardo:2001:ZVL

[LS01] Maria Carmela Lombardo and Marco Sammartino. Zero viscosity limit of the Oseen equations in a channel. SIAM Journal on Mathematical Analysis, 33(2):390–410, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37201.

Lai:2002:QM

[LS02] Ming-Jun Lai and Larry L. Schumaker. Quadrilateral macroelements. SIAM Journal on Mathematical Analysis, 33(5):1107–1116, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37713.

Lin:2003:SSS

[LS03] Xiao-Biao Lin and Stephen Schecter. Stability of selfsimilar solutions of the Dafermos regularization of a system of conservation laws. SIAM Journal on Mathematical Analysis, 35(4):884–921, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40502.

Liu:2001:CTC

[LT01] Hailiang Liu and Eitan Tadmor. Critical thresholds in a convolution model for nonlinear conservation laws. SIAM Journal on Mathematical Analysis, 33(4):930–945, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38690.

Liz:2003:GSC

[LTT03] Eduardo Liz, Victor Tkachenko, and Sergei Trofimchuk. A global stability criterion for scalar functional differential equations. SIAM Journal on Mathematical Analysis, 35(3): 596–622, January 2003. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurencot:2005:SSC

[LW05] Philippe Laurençot and Christoph Walker. Steady states for a coagulation-fragmentation equation with volume scattering. SIAM Journal on Mathematical Analysis, 37(2):531– 548, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44411.

Lin:2008:OSB

[LW08] Tai-Chia Lin and Juncheng Wei. Orbital stability of bound states of semiclassical nonlinear Schrödinger equations with critical nonlinearity. SIAM Journal on Mathematical Analysis, 40(1):365– 381, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2009:CRA

[LW09] Jian-Guo Liu and Wei-Cheng Wang. Characterization and regularity for axisymmetric solenoidal vector fields with application to Navier–Stokes equation. SIAM Journal on Mathematical Analysis, 41(5): 1825–1850, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lani-Wayda:2005:DLS

[LWS05] Bernhard Lani-Wayda and Klaus R. Schneider. Delayed loss of stability in nonautonomous differential equations with retarded argument. SIAM Journal on Mathematical Analysis, 36(5):1522–1539, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40953.

Luo:2000:IBC

[LXY00] Tao Luo, Zhouping Xin, and Tong Yang. Interface behavior of compressible Navier– Stokes equations with vacuum. SIAM Journal on Mathematical Analysis, 31(6):1175–1191, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33104.

Liang:2005:QPS

[LY05] Zhenguo Liang and Jiangong You. Quasi-periodic solutions for 1D Schrödinger equations with higher order nonlinearity. SIAM Journal on Mathematical Analysis, 36(6):1965–1990, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43501.

Liu:2009:GUT

[LY09] Li Liu and Hairong Yuan.

Global uniqueness of transonic shocks in divergent nozzles for steady potential flows. SIAM Journal on Mathematical Analysis, 41(5):1816–1824, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2005:EPR

[LYY05] Gongbao Li, Shusen Yan, and Jianfu Yang. An elliptic problem related to planar vortex pairs. SIAM Journal on Mathematical Analysis, 36(5):1444–1460, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43055.

Lee:2007:CIF

[LYY07] Yeon Ju Lee, Gang Joon Yoon, and Jungho Yoon. Convergence of increasingly flat radial basis interpolants to polynomial interpolants. SIAM Journal on Mathematical Analysis, 39(2):537–553, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2003:TZN

[LZ03] Jiequan Li and Peng Zhang. The transition from Zeldovich– von Neumann–Doring to Chapman– Jouguet theories for a nonconvex scalar combustion model. SIAM Journal on Mathematical Analysis, 34(3):675–699, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38675.

Lei:2005:GEC

[LZ05] Zhen Lei and Yi Zhou. Global existence of classical solutions for the two-dimensional Oldroyd model via the incompressible limit. SIAM Journal on Mathematical Analysis, 37(3):797–814, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        61881.

Macia:2004:WMD

[Mac04] Fabricio Macià. Wigner measures in the discrete setting: High-frequency analysis of sampling and reconstruction operators. SIAM Journal on Mathematical Analysis, 36(2):347–383, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43152.

Mallik:2000:SLH

[Mal00] Ranjan K. Mallik. On the solution of a linear homogeneous difference equation with variable coefficients. SIAM Journal on Mathematical Analysis, 31(2):375–385, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32964.

Maris:2006:GBT

[Mar06a] Mihai Maris. Global branches of travelling-waves to a Gross– Pitaevskii–Schrödinger system in one dimension. SIAM Journal on Mathematical Analysis, 37(5):1535–1559, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Martel:2006:LPR

[Mar06b] Yvan Martel. Linear problems related to asymptotic stability of solitons of the generalized KdV equations. SIAM Journal on Mathematical Analysis, 38(3):759–781, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mari:2008:NST

[Mar08] Mihai Mari. Nonexistence of supersonic traveling waves for nonlinear Schrödinger equations with nonzero conditions at infinity. SIAM Journal on Mathematical Analysis, 40(3): 1076–1103, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Masselin:2001:RBR

[Mas01] Vincent Masselin. A result on the blow-up rate for the Zakharov system in dimension 3. SIAM Journal on Mathematical Analysis, 33(2):440– 447, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36368.

Mikelic:2008:AME

[MB08] Andro Mikelić and Hans Bruining. Analysis of model equations for stress-enhanced diffusion in coal layers. Part I: Existence of a weak solution. SIAM Journal on Mathematical Analysis, 40(4):1671–1691, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:2001:ESM

[MC01] P. D. Miller and S. R. Clarke.

An exactly solvable model for the interaction of linear waves with Korteweg–de Vries solitons. SIAM Journal on Mathematical Analysis, 33(2):261– 285, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36543.

Marciniak-Czochra:2008:DMR

[MCP08] Anna Marciniak-Czochra and Mariya Ptashnyk. Derivation of a macroscopic receptorbased model using homogenization techniques. SIAM Journal on Mathematical Analysis, 40(1):215–237, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mikeli:2006:RUR

[MDvD06] Andro Mikeli, Vincent Devigne, and C. J. van Duijn. Rigorous upscaling of the reactive flow through a pore, under dominant Peclet and Damkohler numbers. SIAM Journal on Mathematical Analysis, 38(4):1262–1287, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meirmanov:2008:DSA

[Mei08] Anvarbek Meirmanov. A description of seismic acoustic wave propagation in porous media via homogenization. SIAM Journal on Mathematical Analysis, 40(3):1272–1289, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Montesinos:2005:RSN

[MG05] María Teresa González Montesinos and Francisco Ortegón Gallego. Renormalized solutions to a nonlinear parabolicelliptic system. SIAM Journal on Mathematical Analysis, 36(6):1991–2003, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42304.

Menon:2001:IDG

[MH01] Govind Menon and György Haller. Infinite dimensional geometric singular perturbation theory for the Maxwell– Bloch equations. SIAM Journal on Mathematical Analysis, 33(2):315–346, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36045.

Mareno:2006:GCS

[MH06] Anita Mareno and Timothy J. Healey. Global continuation in second-gradient nonlinear elasticity. SIAM Journal on Mathematical Analysis, 38(1):103– 115, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mielke:2004:EMI

[Mie04] Alexander Mielke. Existence of minimizers in incremental elasto-plasticity with finite strains. SIAM Journal on Mathematical Analysis, 36 (2):384–404, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
42990.

Mikulevicius:2002:CPS

[Mik02]  R. Mikulevicius.  On the

Cauchy problem for stochastic Stokes equations. SIAM Journal on Mathematical Analysis, 34(1):121–141, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39031.

Mikulevicius:2009:SSS

[Mik09] R. Mikulevicius. On strong H_{2}^{1}-solutions of stochastic Navier–Stokes equation in a bounded domain. SIAM Journal on Mathematical Analysis, 41(3):1206–1230, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mizumachi:2001:LTA

[Miz01]  Tetsu Mizumachi.  Large

time asymptotics of solutions around solitary waves to the generalized Korteweg–de Vries equations. SIAM Journal on Mathematical Analysis, 32(5):1050–1080, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34682.

Mizumachi:2003:WIB

[Miz03] Tetsu Mizumachi. Weak interaction between solitary waves of the generalized KdV equations. SIAM Journal on Mathematical Analysis, 35(4): 1042–1080, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40871.

Moeser:2004:SPS

[MJZ04] Jamison T. Moeser, Christopher K. R. T. Jones, and Vadim Zharnitsky. Stable pulse solutions for the nonlinear Schrödinger equation with higher order dispersion management. SIAM Journal on Mathematical Analysis, 35(6):1486–1511, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41258.

Mingione:2005:IFG

[MM05] Giuseppe Mingione and Domenico Mucci. Integral functionals and the gap problem: Sharp bounds for relaxation and energy concentration. SIAM Journal on Mathematical Analysis, 36(5):1540–1579, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42411.

Maeght:2000:RDE

[MN00] Jean Maeght and Dominikus Noll. Resolution in dynamic emission tomography. SIAM Journal on Mathematical Analysis, 31(5):1100–1120, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34545.

Mercier:2005:EUR

[MN05] D. Mercier and S. Nicaise. Existence, uniqueness, and regularity results for piezoelectric systems. SIAM Journal on Mathematical Analysis, 37 (2):651–672, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
61772.

Macri:2005:ASV

[MNR05] Marta Macrì, Margherita Nolasco, and Tonia Ricciardi. Asymptotics for selfdual vortices on the torus and on the plane: A gluing technique. SIAM Journal on Mathematical Analysis, 37(1):1–16, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61984.

Migorski:2009:QSH

[MO09] Stanislaw Migórski and Anna Ochal. Quasi-static hemivariational inequality via vanishing acceleration approach. SIAM Journal on Mathematical Analysis, 41(4):1415–1435, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moore:2002:LTO

[Moo02] K. S. Moore. Large torsional oscillations in a suspension bridge: Multiple periodic solutions to a nonlinear wave equation. SIAM Journal on Mathematical Analysis, 33(6):1411–1429, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38809.

Mackey:2006:POB

[MOPMW06] Michael C. Mackey, Chunhua Ou, Laurent Pujo-Menjouet, and Jianhong Wu. Periodic oscillations of blood cell populations in chronic myelogenous leukemia. SIAM Journal on Mathematical Analysis, 38(1): 166–187, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morini:2003:SSP

[Mor03]  Massimiliano Morini.  Se-

quences of singularly perturbed functionals generating free-discontinuity problems. SIAM Journal on Mathematical Analysis, 35(3):759– 805, January 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moser:2005:HOA

[Mos05] Roger Moser. A higher order asymptotic problem related to phase transitions. SIAM Journal on Mathematical Analysis, 37(3):712–736, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61676.

Marcati:2001:DPS

[MP01] Pierangelo Marcati and Ronghua Pan. On the diffusive profiles for the system of compressible adiabatic flow through porous media. SIAM Journal on Mathematical Analysis, 33 (4):790–826, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36440.

Makridakis:2003:ORC

[MP03] Charalambos Makridakis and Benoît Perthame. Optimal rate of convergence for anisotropic vanishing viscosity limit of a scalar balance law. SIAM Journal on Mathematical Analysis, 34(6):1300–1307, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40799.

Menon:2005:DSS

[MP05] Govind Menon and Robert L. Pego. Dynamical scaling in Smoluchowski’s coagulation equations: Uniform convergence. SIAM Journal on Mathematical Analysis, 36(5): 1629–1651, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        43026.

Marigo:2008:FDM

[MP08] Alessia Marigo and Benedetto Piccoli. A fluid dynamic model for T -junctions. SIAM Journal on Mathematical Analysis, 39 (6):2016–2032, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mielke:2009:EAM

[MPP09] Alexander Mielke, Laetitia Paoli, and Adrien Petrov. On existence and approximation for a 3D model of thermally induced phase transformations in shape-memory alloys. SIAM Journal on Mathematical Analysis, 41(4):1388– 1414, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morales:2005:ELA

[MPS05] C. A. Morales, M. J. Pacifico, and B. San Martin. Expanding Lorenz attractors through resonant double homoclinic loops. SIAM Journal on Mathematical Analysis, 36(6):1836–1861, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41578.

Morales:2006:CLA

[MPS06] C. A. Morales, M. J. Pacifico, and B. San Martin. Contracting Lorenz attractors through resonant double homoclinic loops. SIAM Journal on Mathematical Analysis, 38(1):309– 332, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Molinet:2002:GCP

[MR02] Luc Molinet and Francis Ribaud. The global Cauchy problem in Bourgain’s-type spaces for a dispersive dissipative semilinear equation. SIAM Journal on Mathematical Analysis, 33(6):1269–1296, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37463.

Mikulevicius:2004:SNS

[MR04] R. Mikulevicius and B. L. Rozovskii. Stochastic Navier– Stokes equations for turbulent flows. SIAM Journal on Mathematical Analysis, 35(5):1250–1310, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40916.

Mucha:2008:NLE

[MR08] Piotr B. Mucha and Piotr Rybka. A new look at equilibria in Stefan-type problems in the plane. SIAM Journal on Mathematical Analysis, 39(4): 1120–1134, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Matthews:2004:WPF

[MS04] John V. Matthews and David G. Schaeffer. A well-posed free boundary value problem for a hyperbolic equation with Dirichlet boundary conditions. SIAM Journal on Mathematical Analysis, 36(1):256–271, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40870.

Metcalfe:2006:LTE

[MS06] Jason Metcalfe and Christopher D. Sogge. Long-time existence of quasilinear wave equations exterior to star-shaped obstacles via energy methods. SIAM Journal on Mathematical Analysis, 38(1):188– 209, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Molinet:2001:IPI

[MST01] L. Molinet, J. C. Saut, and N. Tzvetkov. Ill-posedness issues for the Benjamin–Ono and related equations. SIAM Journal on Mathematical Analysis, 33(4):982–988, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38530.

Molinet:2007:RMC

[MST07] L. Molinet, J. C. Saut, and N. Tzvetkov. Remarks on the mass constraint for KP-type equations. SIAM Journal on Mathematical Analysis, 39(2): 627–641, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Magnanini:2003:CVS

[MT03] Rolando Magnanini and Giorgio Talenti. On complex-valued solutions to a two-dimensional eikonal equation. II. existence theorems. SIAM Journal on Mathematical Analysis, 34 (4):805–835, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40087.

Masmoudi:2007:DLS

[MT07a] Nader Masmoudi and Mohamed Lazhar Tayeb. Diffusion limit of a semiconductor Boltzmann–Poisson system. SIAM Journal on Mathematical Analysis, 38(6):1788– 1807, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mielke:2007:TSH

[MT07b] Alexander Mielke and Aida M. Timofte. Two-scale homogenization for evolutionary variational inequalities via the energetic formulation. SIAM Journal on Mathematical Analysis, 39(2):642–668, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Molinet:2008:NLW

[MT08] Luc Molinet and Raafat Talhouk. Newtonian limit for weakly viscoelastic fluid flows of Oldroyd type. SIAM Journal on Mathematical Analysis, 39(5):1577–1594, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morita:2009:ESL

[MT09] Yoshihisa Morita and Koichi Tachibana. An entire solution to the Lotka–Volterra competition-diffusion equations. SIAM Journal on Mathematical Analysis, 40(6):2217– 2240, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mustafa:2005:EUS

[Mus05] Octavian G. Mustafa. Existence and uniqueness of solutions with low regularity for a class of nonlinear dispersive equations. SIAM Journal on Mathematical Analysis, 37(4):1117–1130, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61239.

Medville:2005:BBP

[MV05] Kai Medville and Michael Vogelius. Blow-up behavior of planar harmonic functions satisfying a certain exponential Neumann boundary condition. SIAM Journal on Mathematical Analysis, 36(6):1772–1806, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43609.

Madureira:2007:APP

[MV07] Alexandre L. Madureira and Frédéric Valentin. Asymptotics of the Poisson problem in domains with curved rough boundaries. SIAM Journal on Mathematical Analysis, 38(5): 1450–1473, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mellet:2008:EUG

[MV08] A. Mellet and A. Vasseur.

Existence and uniqueness of global strong solutions for one-dimensional compressible Navier–Stokes equations.

SIAM Journal on Mathematical Analysis, 39(4):1344–1365, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Munoz-Villarragut:2008:NFD

[MVNO08] Víctor Muñoz-Villarragut, Sylvia Novo, and Rafael Obaya. Neutral functional differential equations with applications to compartmental systems. SIAM Journal on Mathematical Analysis, 40(3):1003– 1028, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Martinez:2009:SPP

[MW09] Sandra Martínez and Noemi Wolanski. A singular perturbation problem for a quasi-linear operator satisfying the natural growth condition of Lieberman. SIAM Journal on Mathematical Analysis, 41(1):318– 359, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Magal:2005:GAS

[MZ05a] Pierre Magal and Xiao-Qiang Zhao. Global attractors and steady states for uniformly persistent dynamical systems. SIAM Journal on Mathematical Analysis, 37(1):251–275, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43917.

Mascia:2005:SLA

[MZ05b] Corrado Mascia and Kevin Zumbrun. Stability of largeamplitude shock profiles of general relaxation systems. SIAM Journal on Mathematical Analysis, 37(3):889–913, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43584.

Nakanishi:2001:AFS

[Nak01] Kenji Nakanishi. Asymptoticallyfree solutions for the shortrange nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 32(6): 1265–1271, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36908.

Nakamura:2003:ADT

[Nak03] Tohru Nakamura. Asymptotic decay toward the rarefaction waves of solutions for viscous conservation laws in a one-dimensional half space. SIAM Journal on Mathematical Analysis, 34(6):1308–1317, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40693.

Nara:2008:LTB

[Nar08] Mitsunori Nara. Large time behavior of radially symmetric surfaces in the mean curvature flow. SIAM Journal on Mathematical Analysis, 39(6): 1978–1995, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nesenenko:2007:HV

[Nes07] Sergiy Nesenenko. Homogenization in viscoplasticity. SIAM Journal on Mathematical Analysis, 39(1):236–262, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nikolov:2001:IDS

[Nik01]  Geno Nikolov.  Inequali-

ties of Duffin–Schaeffer type. SIAM Journal on Mathematical Analysis, 33(3):686–698, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37391.

Nishibata:2002:AST

[Nis02] Shinya Nishibata. Asymptotic stability of traveling waves to a certain discrete velocity model of the Boltzmann equation in the half space. SIAM Journal on Mathematical Analysis, 34(3):555–572, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39068.

Neff:2008:RBN

[NK08] Patrizio Neff and Dorothee Knees. Regularity up to the boundary for nonlinear elliptic systems arising in time-incremental infinitesimal elasto-plasticity. SIAM Journal on Mathematical Analysis, 40(1):21–43, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishihara:2001:ABS

[NN01] Kenji Nishihara and Masataka Nishikawa. Asymptotic behavior of solutions to the system of compressible adiabatic flow through porous media. SIAM Journal on Mathematical Analysis, 33(1):216– 239, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36467.

Nakamura:2009:CRT

[NN09] Tohru Nakamura and Shinya Nishibata. Convergence rate toward planar stationary waves for compressible viscous fluid in multidimensional half space. SIAM Journal on Mathematical Analysis, 41(5):1757–1791, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Noble:2009:PRW

[Nob09] Pascal Noble. Persistence of roll waves for the Saint Venant equations. SIAM Journal on Mathematical Analysis, 40(5): 1783–1814, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Novo:2009:EON

[NOV09] Sylvia Novo, Rafael Obaya, and Víctor M. Villarragut. Exponential ordering for nonautonomous neutral functional differential equations. SIAM Journal on Mathematical Analysis, 41(3):1025–1053, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Niethammer:2000:IVP

[NP00] Barbara Niethammer and Robert L. Pego. On the initialvalue problem in the Lifshitz– Slyozov–Wagner theory of Ostwald ripening. SIAM Journal on Mathematical Analysis, 31 (3):467–485, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
33821.

Narcowich:2006:LTF

[NPW06] F. J. Narcowich, P. Petrushev, and J. D. Ward. Localized tight frames on spheres. SIAM Journal on Mathematical Analysis, 38(2):574–594, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Neuss-Radu:2007:ETC

[NRJ07] Maria Neuss-Radu and Willi Jäger. Effective transmission conditions for reactiondiffusion processes in domains separated by an interface. SIAM Journal on Mathematical Analysis, 39(3):687–720, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishiura:2004:HDS

[NS04] Yasumasa Nishiura and Hiromasa Suzuki. Higher dimensional SLEP equation and applications to morphological stability in polymer problems. SIAM Journal on Mathematical Analysis, 36(3):916– 966, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42015.

Neumann:2005:CGE

[NS05] Lukas Neumann and Christian Schmeiser. Convergence to global equilibrium for a kinetic fermion model. SIAM Journal on Mathematical Analysis, 36(5):1652–1663, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43653.

Nakamura:2007:OBD

[NS07] Gen Nakamura and Mourad Sini. Obstacle and boundary determination from scattering data. SIAM Journal on Mathematical Analysis, 39 (3):819–837, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Natile:2009:WAO

[NS09] Luca Natile and Giuseppe Savaré. A Wasserstein approach to the one-dimensional sticky particle system. SIAM Journal on Mathematical Analysis, 41(4):1340–1365, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2008:PEE

[NT08] Truyen Nguyen and Adrian Tudorascu. Pressureless Euler/Euler–Poisson systems via adhesion dynamics and scalar conservation laws. SIAM Journal on Mathematical Analysis, 40(2):754–775, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Narcowich:2002:SDI

[NW02] Francis J. Narcowich and Joseph D. Ward. Scattered data interpolation on spheres: Error estimates and locally supported basis functions. SIAM Journal on Mathematical Analysis, 33(6):1393–1410, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39505.

Nakamura:2003:UCE

[NW03] Gen Nakamura and Jenn-Nan Wang. Unique continuation for an elasticity system with residual stress and its applications. SIAM Journal on Mathematical Analysis, 35(2):304– 317, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39974.

Narcowich:2004:SDI

[NW04] Francis J. Narcowich and Joseph D. Ward. Scattereddata interpolation on R^{n}: Error estimates for radial basis and band-limited functions. SIAM Journal on Mathematical Analysis, 36(1):284–300, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41357.

Nishihara:2004:NSS

[NYZ04] Kenji Nishihara, Tong Yang, and Huijiang Zhao. Nonlinear stability of strong rarefaction waves for compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 35(6):1561–1597, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42735.

ODell:2008:ISS

[O’D08] Stephen O’Dell. Inverse scattering for Schrödinger-type operators with interface conditions across smooth surfaces. SIAM Journal on Mathematical Analysis, 39(5):1595–1626, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ogawa:2003:SSI

[Oga03]  Takayoshi Ogawa.  Sharp

Sobolev inequality of logarithmic type and the limiting regularity condition to the harmonic heat flow. SIAM Journal on Mathematical Analysis, 34(6):1318–1330, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39586.

Otto:2006:CRD

[ORS06] Felix Otto, Tobias Rump, and Dejan Slepcev. Coarsening rates for a droplet model: Rigorous upper bounds. SIAM Journal on Mathematical Analysis, 38(2):503–529, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ortner:2006:GFS

[Ort06] Christoph Ortner. Gradient flows as a selection procedure for equilibria of nonconvex energies. SIAM Journal on Mathematical Analysis, 38(4):1214– 1234, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Oshita:2004:SSP

[Osh04]  Yoshihito Oshita.  Stable

stationary patterns and interfaces arising in reactiondiffusion systems. SIAM Journal on Mathematical Analysis, 36(2):479–497, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40672.

Oshita:2007:SLP

[Osh07] Yoshihito Oshita. Singular limit problem for some elliptic systems. SIAM Journal on Mathematical Analysis, 38(6): 1886–1911, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ohta:2007:SIS

[OT07] Masahito Ohta and Grozdena Todorova. Strong instability of standing waves for the nonlinear Klein–Gordon equation and the Klein–Gordon– Zakharov system. SIAM Journal on Mathematical Analysis, 38(6):1912–1931, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Otto:2005:ECC

[OW05] Felix Otto and Michael Westdickenberg. Eulerian calculus for the contraction in the Wasserstein distance. SIAM Journal on Mathematical Analysis, 37(4):1227–1255, January 2005. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ou:2007:TWD

[OW07] Chunhua Ou and Jianhong Wu. Traveling wavefronts in a delayed food-limited population model. SIAM Journal on Mathematical Analysis, 39 (1):103–125, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Palamodov:2009:RDF

[Pal09] Victor Palamodov. Reconstruction of a differential form from Doppler Transform. SIAM Journal on Mathematical Analysis, 41(4):1713–1720, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pan:2003:SFP

[Pan03]  Xing-Bin Pan.  Supercon-

ducting films in perpendicular fields and the effect of the de Gennes parameter. SIAM Journal on Mathematical Analysis, 34(4):957–991, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40673.

Pan:2006:LGM

[Pan06] Xing-Bin Pan. Landau–de Gennes model of liquid crystals with small Ginzburg–Landau parameter. SIAM Journal on Mathematical Analysis, 37(5): 1616–1648, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Panov:2009:WCS

[Pan09] Evgeniy Panov. On weak completeness of the set of entropy solutions to a scalar conservation law. SIAM Journal on Mathematical Analysis, 41(1): 26–36, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pedregal:2000:ODC

[Ped00] Pablo Pedregal. Optimal design and constrained quasiconvexity. SIAM Journal on Mathematical Analysis, 32 (4):854–869, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35650.

Pedregal:2004:CTY

[Ped04] Pablo Pedregal. Γ-Convergence through Young measures. SIAM Journal on Mathematical Analysis, 36(2):423–440, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42569.

Pedregal:2006:YMA

[Ped06] Pablo Pedregal. Young measures associated with homogenization. SIAM Journal on Mathematical Analysis, 37(5): 1454–1464, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peletier:2001:SBV

[Pel01] Mark A. Peletier. Sequential buckling: A variational analysis. SIAM Journal on Mathematical Analysis, 32(5): 1142–1168, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        35992.

Perepelitsa:2006:GEW

[Per06] Mikhail Perepelitsa. On the global existence of weak solutions for the Navier–Stokes equations of compressible fluid flows. SIAM Journal on Mathematical Analysis, 38(4):1126– 1153, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinezich:2000:PST

[Pin00] J. D. Pinezich. Periodic structure in two-dimensional Riemann problems for Hamilton– Jacobi equations. SIAM Journal on Mathematical Analysis, 32(3):688–706, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34214.

Pisante:2004:SCE

[Pis04]  Giovanni Pisante.  Suffi-

cient conditions for the existence of viscosity solutions for nonconvex Hamiltonians. SIAM Journal on Mathematical Analysis, 36(1):186–203, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42690.

Pivovarchik:2000:IPS

[Piv00] Vyacheslav Pivovarchik. Inverse problem for the Sturm– Liouville equation on a simple graph. SIAM Journal on Mathematical Analysis, 32(4):801– 819, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36824.

Plakhov:2009:BSR

[Pla09a] Alexander Plakhov. Billiard scattering on rough sets: Twodimensional case. SIAM Journal on Mathematical Analysis, 40(6):2155–2178, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Plakhov:2009:CFD

[Pla09b] Alexander Plakhov. Comment on “Functions and Domains Having Minimal Resistance Under a Single-Impact Assumption” [SIAM J. Math. Anal., 34 (2002), pp. 101–120]. SIAM Journal on Mathematical Analysis, 41(4):1721–1724, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [CLR02].

Pan:2002:ABO

[PLN02] Tao Pan, Hongxia Liu, and Kenji Nishihara. Asymptotic behavior of a one-dimensional compressible viscous gas with free boundary. SIAM Journal on Mathematical Analysis, 34 (2):273–291, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38574.

Pava:2008:PPF

[PN08] Jaime Angulo Pava and Fábio M. A. Natali. Positivity properties of the Fourier transform and the stability of periodic travelling-wave solutions. SIAM Journal on Mathematical Analysis, 40(3):1123–1151, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ponsiglione:2007:EES

[Pon07] Marcello Ponsiglione. Elastic energy stored in a crystal induced by screw dislocations: From discrete to continuous. SIAM Journal on Mathematical Analysis, 39(2): 449–469, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Potthast:2007:CNR

[Pot07] Roland Potthast. On the convergence of the no response test. SIAM Journal on Mathematical Analysis, 38(6):1808– 1824, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pagani:2000:SNW

[PP00a] Carlo D. Pagani and Dario Pierotti. On solvability of the nonlinear wave resistance problem for a surfacepiercing symmetric cylinder. SIAM Journal on Mathematical Analysis, 32(1):214–233, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34814.

Payne:2000:SDS

[PP00b] L. E. Payne and G. A.

Philippin. On the spatial decay of solutions to a quasi-linear parabolic initialboundary value problem and their derivatives. SIAM Journal on Mathematical Analysis, 32(2):291–303, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35651.

Pascucci:2003:CPN

[PP03] Andrea Pascucci and Sergio Polidoro. On the Cauchy problem for a nonlinear Kolmogorov equation. SIAM Journal on Mathematical Analysis, 35(3):579–595, January 2003. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pagani:2004:SMS

[PP04] Carlo D. Pagani and Dario Pierotti. The subcritical motion of a semisubmerged body: Solvability of the free boundary problem. SIAM Journal on Mathematical Analysis, 36(1):69–93, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
42598.

Piatnitski:2001:HED

[PR01] Andrey Piatnitski and Elisabeth Remy. Homogenization of elliptic difference operators. SIAM Journal on Mathematical Analysis, 33(1):53– 83, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33808.

Protasov:2000:CSC

[Pro00] Vladimir Protasov. A complete solution characterizing smooth refinable functions. SIAM Journal on Mathematical Analysis, 31(6):1332–1350, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34296.

Protasov:2001:SSO

[Pro01] Vladimir Protasov. The stability of subdivision operator at its fixed point. SIAM Journal on Mathematical Analysis, 33(2):448–460, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35628.

Promislow:2002:RMM

[Pro02] Keith Promislow. A renormalization method for modulational stability of quasi-steady patterns in dispersive systems. SIAM Journal on Mathematical Analysis, 33(6):1455–1482, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37754.

Plotnikov:2008:IBV

[PRS08] P. I. Plotnikov, E. V. Ruban, and J. Sokolowski. Inhomogeneous boundary value problems for compressible Navier– Stokes equations: Well-posedness and sensitivity analysis. SIAM Journal on Mathematical Analysis, 40(3):1152–1200, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peherstorfer:2000:SAO

[PS00] Franz Peherstorfer and Robert Steinbauer. Strong asymptotics of orthonormal polynomials with the aid of Green’s function. SIAM Journal on Mathematical Analysis, 32(2): 385–402, 2000. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34304.

Pucci:2006:DCB

[PS06] Patrizia Pucci and James Serrin. Dead cores and bursts for quasilinear singular elliptic equations. SIAM Journal on Mathematical Analysis, 38(1): 259–278, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Paivarinta:2008:TE

[PS08a] Lassi Päivärinta and John Sylvester. Transmission eigenvalues. SIAM Journal on Mathematical Analysis, 40(2): 738–753, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pruss:2008:SES

[PS08b] Jan Prüss and Gieri Simonett. Stability of equilibria for the Stefan problem with surface tension. SIAM Journal on Mathematical Analysis, 40 (2):675–698, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Parini:2009:PPP

[PS09a] Enea Parini and Athanasios Stylianou. On the positivity preserving property of hinged plates. SIAM Journal on Mathematical Analysis, 41(5): 2031–2037, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Petrov:2009:MRE

[PS09b] Adrien Petrov and Michelle Schatzman. Mathematical results on existence for viscoelastodynamic problems with unilateral constraints. SIAM Journal on Mathematical Analysis, 40(5):1882–1904, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pennacchio:2005:MMB

[PSF05] Micol Pennacchio, Giuseppe Savaré, and Piero Colli Franzone. Multiscale modeling for the bioelectric activity of the heart. SIAM Journal on Mathematical Analysis, 37(4):1333– 1370, January 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Petrosyan:2005:DED

[PV05] Arshak Petrosyan and Enrico Valdinoci. Density estimates for a degenerate/singular phase-transition model. SIAM Journal on Mathematical Analysis, 36(4): 1057–1079, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43767.

Peng:2008:RDI

[PW08] Yue-Jun Peng and Shu Wang. Rigorous derivation of incompressible e-MHD equations from compressible Euler– Maxwell equations. SIAM Journal on Mathematical Analysis, 40(2):540–565, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Qian:2005:PML

[QT05] Dingbian Qian and Pedro J. Torres. Periodic motions of linear impact oscillators via the successor map. SIAM Journal on Mathematical Analysis, 36(6):1707–1725, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43771.

Qiu:2000:UAF

[QW00] W.-Y. Qiu and R. Wong. Uniform asymptotic formula for orthogonal polynomials with exponential weight. SIAM Journal on Mathematical Analysis, 31(5):992–1029, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34467.

Qin:2009:SWP

[QW09] Xiaohong Qin and Yi Wang.

Stability of wave patterns to the inflow problem of full compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 41(5):2057– 2087, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Qin:2008:SSW

[QXM08] Wen-Xin Qin, Chun-Lan Xu, and Xin Ma. Stability of singlewave-form solutions in the underdamped Frenkel–Kontorova model. SIAM Journal on Mathematical Analysis, 40(3): 952–967, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reddien:2002:PTT

[Red02]  G. W. Reddien.  Path-

tracking through singularities. SIAM Journal on Mathematical Analysis, 33(6):1305–1319, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37387.

Rein:2001:RCC

[Rei01] Gerhard Rein. Reduction and a concentration-compactness principle for energy-Casimir functionals. SIAM Journal on Mathematical Analysis, 33 (4):896–912, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38927.

Renato:2000:NFS

[Ren00] Manfrin Renato. A note on the formation of singularities for quasi-linear hyperbolic systems. SIAM Journal on Mathematical Analysis, 32(2):261– 290, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34152.

Renard:2006:UCS

[Ren06]  Yves Renard.  A unique-

ness criterion for the Signorini problem with Coulomb friction. SIAM Journal on Mathematical Analysis, 38(2):452– 467, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Riaza:2004:ADD

[Ria04] Ricardo Riaza. Attraction domains of degenerate singular equilibria in quasi-linear ODEs. SIAM Journal on Mathematical Analysis, 36(2): 678–690, 2004. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
42500.

Rieger:2003:YMS

[Rie03] Marc Oliver Rieger. Young measure solutions for nonconvex elastodynamics. SIAM Journal on Mathematical Analysis, 34(6):1380–1398, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39214.

Rigot:2006:QPP

[Rig06] Séverine Rigot. Quasiminimal partitions with prescribed measure. SIAM Journal on Mathematical Analysis, 37(5): 1589–1615, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rybka:2005:EEM

[RL05] Piotr Rybka and Mitchell Luskin. Existence of energy minimizers for magnetostrictive materials. SIAM Journal on Mathematical Analysis, 36(6):2004–2019, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44202.

Rohwer:2006:DPT

[RL06] C. H. Rohwer and D. P. Laurie. The discrete pulse transform. SIAM Journal on Mathematical Analysis, 38(3):1012– 1034, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Robinson:2000:NLA

[Rob00]  Clark Robinson.  Non-

symmetric Lorenz attractors from a homoclinic bifurcation. SIAM Journal on Mathematical Analysis, 32(1):119–141, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34359.

Robinson:2004:SEN

[Rob04] Stephen B. Robinson. On the second eigenvalue for nonhomogeneous quasi-linear operators. SIAM Journal on Mathematical Analysis, 35(5): 1241–1249, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
42600.

Roger:2005:EWS

[Rög05] Matthias Röger. Existence of weak solutions for the Mullins– Sekerka flow. SIAM Journal on Mathematical Analysis, 37 (1):291–301, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43964.

Rohde:2005:SCL

[Roh05] Christian Rohde. Scalar conservation laws with mixed local and nonlocal diffusiondispersion terms. SIAM Journal on Mathematical Analysis, 37(1):103–129, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44330.

Rondi:2005:OSR

[Ron05] Luca Rondi. Optimal stability of reconstruction of plane Lipschitz cracks. SIAM Journal on Mathematical Analysis, 36(4):1282–1292, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43583.

Rousset:2003:VAS

[Rou03] Frederic Rousset. Viscous approximation of strong shocks of systems of conservation laws. SIAM Journal on Mathematical Analysis, 35(2):492–519, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40311.

Rousset:2006:CIP

[Rou06] Mathias Rousset. On the control of an interacting particle estimation of Schrödinger ground states. SIAM Journal on Mathematical Analysis, 38(3):824–844, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Roubicek:2007:IIN

[Rou07] Tomáš Roubíček. Incompressible ionized non-Newtonian fluid mixtures. SIAM Journal on Mathematical Analysis, 39 (3):863–890, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Resnikoff:2001:BWS

[RTW01] Howard L. Resnikoff, Jun Tian, and Raymond O. Wells, Jr. Biorthogonal wavelet space: Parametrization and factorization. SIAM Journal on Mathematical Analysis, 33(1):194– 215, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36036.

Ren:2000:MST

[RW00a] Xiaofeng Ren and Juncheng Wei. On the multiplicity of solutions of two nonlocal variational problems. SIAM Journal on Mathematical Analysis, 31(4):909–924, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34817.

Ruskai:2000:SCR

[RW00b] Mary Beth Ruskai and Elisabeth Werner. Study of a class of regularizations of 1/|x| using Gaussian integrals. SIAM Journal on Mathematical Analysis, 32(2):435–463, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35375.

Ren:2003:STD

[RW03] Xiaofeng Ren and Juncheng Wei. On the spectra of threedimensional lamellar solutions of the diblock copolymer problem. SIAM Journal on Mathematical Analysis, 35(1):1–32, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41334.

Ren:2005:WLS

[RW05] Xiaofeng Ren and Juncheng Wei. Wriggled lamellar solutions and their stability in the diblock copolymer problem. SIAM Journal on Mathematical Analysis, 37(2):455– 489, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43358.

Ren:2008:SSN

[RW08] Xiaofeng Ren and Juncheng Wei. Spherical solutions to a nonlocal free boundary problem from diblock copolymer morphology. SIAM Journal on Mathematical Analysis, 39(5): 1497–1535, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rybkin:2001:TAI

[Ryb01]  Alexei Rybkin.  On the

trace approach to the inverse scattering problem in dimension one. SIAM Journal on Mathematical Analysis, 32(6):1248–1264, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36562.

Sacchetti:2004:NTD

[Sac04]  Andrea Sacchetti.  Non-

linear time-dependent onedimensional Schrödinger equation with double-well potential. SIAM Journal on Mathematical Analysis, 35(5):1160– 1176, 2004. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41543.

Sangalli:2005:UAN

[San05] Giancarlo Sangalli. A uniform analysis of nonsymmetric and coercive linear operators. SIAM Journal on Mathematical Analysis, 36(6):2033–2048, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43499.

Schatzman:2000:SRP

[Sch00] Michelle Schatzman. Stability with respect to pseudodifferential perturbations of some nonlinear diffusive equations. SIAM Journal on Mathematical Analysis, 32(3):637–650, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35786.

Schreiber:2002:PWC

[Sch02] Sebastian J. Schreiber. Permanence of weakly coupled vector fields. SIAM Journal on Mathematical Analysis, 33(5):1048–1057, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38743.

Schweizer:2005:LCP

[Sch05] Ben Schweizer. Laws for the capillary pressure in a deterministic model for fronts in porous media. SIAM Journal on Mathematical Analysis, 36(5):1489–1521, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42305.

Schweizer:2008:HDT

[Sch08] Ben Schweizer. Homogenization of degenerate two-phase flow equations with oil trapping. SIAM Journal on Mathematical Analysis, 39(6):1740– 1763, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schonbek:2009:EDP

[Sch09] Maria E. Schonbek. Existence and decay of polymeric flows. SIAM Journal on Mathematical Analysis, 41(2):564–587, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sincich:2006:SDS

[Sin06] E. Sincich. Stable determination of the surface impedance of an obstacle by far field measurements. SIAM Journal on Mathematical Analysis, 38(2): 434–451, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Slepcev:2008:CNI

[Sle08] Dejan Slepčev. Coarsening in nonlocal interfacial systems. SIAM Journal on Mathematical Analysis, 40(3):1029–1048, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:2007:OTV

[SP07] Wen Shen and Mee Rea Park. Optimal tracing of viscous shocks in solutions of viscous conservation laws. SIAM Journal on Mathematical Analysis, 38(5):1474–1488, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Spirn:2003:VML

[Spi03] Daniel Spirn. Vortex motion law for the Schrödinger– Ginzburg–Landau equations. SIAM Journal on Mathematical Analysis, 34(6):1435–1476, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39666.

Saint-Raymond:2009:WCM

[SR09] Laure Saint-Raymond. Weak compactness methods for singular penalization problems with boundary layers. SIAM Journal on Mathematical Analysis, 41(1):153–177, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sandier:2003:DBS

[SS03a] Etienne Sandier and Sylvia Serfaty. The decrease of bulk-superconductivity close to the second critical field in the Ginzburg–Landau model. SIAM Journal on Mathematical Analysis, 34(4):939–956, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40608.

Schonbek:2003:ABD

[SS03b] Maria E. Schonbek and Tomas P.

Schonbek. Asymptotic behavior to dissipative quasigeostrophic flows. SIAM Journal on Mathematical Analysis, 35(2):357–375, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        40936.

Sandstede:2008:HBV

[SS08] Björn Sandstede and Arnd Scheel. Hopf bifurcation from viscous shock waves. SIAM Journal on Mathematical Analysis, 39(6):2033–2052, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Salem:2009:SAS

[SS09a] Walid K. Abou Salem and Catherine Sulem. Stochastic acceleration of solitons for the nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 41(1):117– 152, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simpson:2009:AEN

[SS09b] Henry C. Simpson and Scott J. Spector. Applications of estimates near the boundary to regularity of solutions in linearized elasticity. SIAM Journal on Mathematical Analysis, 41(3):923–935, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sideris:2001:GES

[ST01] Thomas C. Sideris and Shu-Yi Tu. Global existence for systems of nonlinear wave equations in 3D with multiple speeds. SIAM Journal on Mathematical Analysis, 33 (2):477–488, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37896.

Saxton:2008:GES

[ST08] Ralph Saxton and Feride Tiğlay. Global existence of some infinite energy solutions for a perfect incompressible fluid. SIAM Journal on Mathematical Analysis, 40(4):1499– 1515, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stevenson:2004:COW

[Ste04] Rob Stevenson. On the compressibility of operators in wavelet coordinates. SIAM Journal on Mathematical Analysis, 35(5):1110–1132, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41152.

Stefanelli:2005:ATM

[Ste05]  Ulisse Stefanelli.  Analy-

sis of a thermomechanical model for shape memory alloys. SIAM Journal on Mathematical Analysis, 37(1):130– 155, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44425.

Stefanelli:2008:VPH

[Ste08] Ulisse Stefanelli. A variational principle for hardening elastoplasticity. SIAM Journal on Mathematical Analysis, 40(2): 623–652, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stiefenhofer:2000:USP

[Sti00] Matthias Stiefenhofer. Unfolding singularly perturbed Bogdanov points. SIAM Journal on Mathematical Analysis, 32 (4):820–853, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        33423.

So:2002:SLD

[STZ02] Joseph W.-H. So, Xianhua Tang, and Xingfu Zou. Stability in a linear delay system without instantaneous negative feedback. SIAM Journal on Mathematical Analysis, 33(6):1297–1304, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38926.

Sugiyama:2009:RTA

[Sug09]  Yoshie Sugiyama.  On ε-

regularity theorem and asymptotic behaviors of solutions for Keller–Segel systems. SIAM Journal on Mathematical Analysis, 41(4):1664–1692, ????

2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sun:2001:ELA

[Sun01] S. M. Sun. Existence of large amplitude periodic waves in two-fluid flows of infinite depth. SIAM Journal on Mathematical Analysis, 32(5):1014–1031, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35272.

Sun:2007:NAS

[Sun07] Qiyu Sun. Nonuniform average sampling and reconstruction of signals with finite rate of innovation. SIAM Journal on Mathematical Analysis, 38 (5):1389–1422, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simpson:2008:ASA

[SW08] Gideon Simpson and Michael I.

Weinstein. Asymptotic stability of ascending solitary magma waves. SIAM Journal on Mathematical Analysis, 40 (4):1337–1391, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Santra:2009:HSF

[SW09] Sanjiban Santra and Juncheng Wei. Homoclinic solutions for fourth order traveling wave equations. SIAM Journal on Mathematical Analysis, 41(5): 2038–2056, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:2000:GAS

[SZ00] Hal L. Smith and Xiao-Qiang Zhao. Global asymptotic stability of traveling waves in delayed reaction-diffusion equations. SIAM Journal on Mathematical Analysis, 31(3):514– 534, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34678.

Swanson:2004:IWD

[SZ04] David Swanson and William P. Ziemer. The image of a weakly differentiable mapping. SIAM Journal on Mathematical Analysis, 35(5):1099–1109, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41206.

Stuart:2005:ATM

[SZ05] Charles A. Stuart and Huan-Song Zhou. Axisymmetric TEmodes in a self-focusing dielectric. SIAM Journal on Mathematical Analysis, 37(1):218– 237, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44175.

Seregin:2007:SCR

[SZ07] G. Seregin and W. Zajaczkowski. A sufficient condition of regularity for axially symmetric solutions to the Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 39(2):669–685, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schwetlick:2009:EDP

[SZ09] Hartmut Schwetlick and Johannes Zimmer. Existence of dynamic phase transitions in a one-dimensional lattice model with piecewise quadratic interaction potential. SIAM Journal on Mathematical Analysis, 41(3):1231–1271, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Szeftel:2005:MDS

[Sze05] Jérémie Szeftel. Microlocal dispersive smoothing for the nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 37(2):549– 597, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43210.

Takeuchi:2000:BSN

[Tak00] Shingo Takeuchi. Behavior of solutions near the flat hats of stationary solutions for a degenerate parabolic equation.

SIAM Journal on Mathematical Analysis, 31(3):678–692, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34257.

Taniguchi:2007:TFP

[Tan07] Masaharu Taniguchi. Traveling fronts of pyramidal shapes in the Allen–Cahn equations. SIAM Journal on Mathematical Analysis, 39(1):319–344, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Thevenot:2004:OFD

[The04] Laurent Thevenot. On the optimization of the fuel distribution in a nuclear reactor. SIAM Journal on Mathematical Analysis, 35(5):1133–1159, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41506.

Torres:2004:PLM

[Tor04]  Monica Torres.  Plane-like

minimal surfaces in periodic media with exclusions. SIAM Journal on Mathematical Analysis, 36(2):523–551, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39997.

Tovbis:2008:BSP

[Tov08] Alexander Tovbis. Breaking of symmetrical periodic solutions in a singularly perturbed KDV model. SIAM Journal on Mathematical Analysis, 40(4): 1516–1549, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trakhinin:2006:DSH

[Tra06]  Yuri Trakhinin.  Dissipa-

tive symmetrizers of hyperbolic problems and their applications to shock waves and characteristic discontinuities. SIAM Journal on Mathematical Analysis, 37(6):1988–2024, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trivisa:2008:DLV

[Tri08] Konstantina Trivisa. On the dynamics of liquid-vapor phase transition. SIAM Journal on Mathematical Analysis, 39(6): 1788–1820, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tsai:2007:AST

[Tsa07] Je-Chiang Tsai. Asymptotic stability of traveling wave fronts in the buffered bistable system. SIAM Journal on Mathematical Analysis, 39(1): 138–159, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tadmor:2000:PEE

[TT00] Eitan Tadmor and Tao Tang. Pointwise error estimates for relaxation approximations to conservation laws. SIAM Journal on Mathematical Analysis, 32(4):870–886, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        34949.

Tang:2003:FRC

[TTX03] Tao Tang, Zhen-Huan Teng, and Zhouping Xin. Fractional rate of convergence for viscous approximation to nonconvex conservation laws. SIAM Journal on Mathematical Analysis, 35(1):98–122, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38899.

Tao:2009:CCH

[TW09] Youshan Tao and Mingjun Wang. A combined chemotaxishaptotaxis system: The role of logistic source. SIAM Journal on Mathematical Analysis, 41 (4):1533–1558, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trouve:2005:LGD

[TY05] Alain Trouvé and Laurent Younes. Local geometry of deformable templates. SIAM Journal on Mathematical Analysis, 37(1):17–59, 2005. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40483.

Tzirakis:2006:MCP

[Tzi06] Nikolaos Tzirakis. Mass concentration phenomenon for the quintic nonlinear Schrödinger equation in one dimension. SIAM Journal on Mathematical Analysis, 37(6):1923–1946, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tzoukmanis:2006:NPL

[Tzo06]  Niko Tzoukmanis.  Near-

periodic local minimizers of singularly perturbed functionals with nonlocal term. SIAM Journal on Mathematical Analysis, 37(5):1396–1416, January 2006. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Uhlmann:2007:CSW

[UW07] Gunther Uhlmann and Jenn-Nan Wang. Complex spherical waves for the elasticity system and probing of inclusions. SIAM Journal on Mathematical Analysis, 38(6):1967–1980, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Villarroel:2003:DSS

[VA03] Javier Villarroel and Mark J. Ablowitz. On the discrete spectrum of systems in the plane and the Davey–Stewartson II equation. SIAM Journal on Mathematical Analysis, 34(6):1253–1278, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39162.

VanAssche:1985:WZD

[Van85] Walter Van Assche. Weighted zero distribution for polynomials orthogonal on an infinite interval. SIAM Journal on Mathematical Analysis, 16(6):1317– 1334, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Van01].

VanAssche:2001:EWZ

[Van01] Walter Van Assche. Erratum to “Weighted Zero Distribution For Polynomials Orthogonal on an Infinite Interval”. SIAM Journal on Mathematical Analysis, 32(5):1169–1170, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35987.  See

[Van85].

Varvaruca:2008:SPT

[Var08] Eugen Varvaruca. On some properties of traveling water waves with vorticity.

SIAM Journal on Mathematical Analysis, 39(5):1686–1692, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vazquez:2003:DLT

[Váz03] Juan L. Vázquez. Darcy’s law and the theory of shrinking solutions of fast diffusion equations. SIAM Journal on Mathematical Analysis, 35(4):1005–1028, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39654.

vanBaalen:2008:LTL

[vBPW08] Guillaume van Baalen, Nikola Popović, and C. Eugene Wayne. Long tails in the long-time asymptotics of quasilinear hyperbolic-parabolic systems of conservation laws. SIAM Journal on Mathematical Analysis, 39(6):1951–1977, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vandenBerg:2001:TWF

[vdBHV01] Jan Bouwe van den Berg, Josephus Hulshof, and Robertus C. Vandervorst. Travelling waves for fourth order parabolic equations. SIAM Journal on Mathematical Analysis, 32(6): 1342–1374, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.

                                        org/sam-bin/dbq/article/
                                        35830.

vanDuijn:2007:NCE

[vDPP07] C. J. van Duijn, L. A. Peletier, and I. S. Pop. A new class of entropy solutions of the Buckley–Leverett equation. SIAM Journal on Mathematical Analysis, 39(2):507– 536, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Visintin:2007:HDN

[Vis07] Augusto Visintin. Homogenization of a doubly nonlinear Stefan-type problem. SIAM Journal on Mathematical Analysis, 39(3):987–1017, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Visintin:2009:PTG

[Vis09] Augusto Visintin. Phase transitions and glass formation in binary alloys. SIAM Journal on Mathematical Analysis, 41 (5):1725–1756, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vuillaume:2003:SBT

[Vui03] Grégory Vuillaume. Study of the buckling of a tapered rod with the genus of a set. SIAM Journal on Mathematical Analysis, 34(5):1128–1151, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40432.

Visan:2007:BCF

[VZ07] Monica Visan and Xiaoyi Zhang. On the blowup for the L^{2}-critical focusing nonlinear Schrödinger equation in higher dimensions below the energy class. SIAM Journal on Mathematical Analysis, 39(1):34–56, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vancostenoble:2009:HIO

[VZ09] J. Vancostenoble and E. Zuazua.

Hardy inequalities, observability, and control for the wave and Schrödinger equations with singular potentials. SIAM Journal on Mathematical Analysis, 41(4):1508–1532, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wahlen:2006:SPC

[Wah06] Erik Wahlén. Steady periodic capillary-gravity waves with vorticity. SIAM Journal on Mathematical Analysis, 38(3): 921–943, January 2006. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walker:2009:PES

[Wal09a] Christoph Walker. Positive equilibrium solutions for ageand spatially-structured population models. SIAM Journal on Mathematical Analysis, 41 (4):1366–1387, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walsh:2009:SSP

[Wal09b]  Samuel Walsh.  Stratified

steady periodic water waves. SIAM Journal on Mathematical Analysis, 41(3):1054–1105, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2000:QBS

[Wan00]  Xuefeng Wang.  Qualita-

tive behavior of solutions of chemotactic diffusion systems: Effects of motility and chemotaxis and dynamics. SIAM Journal on Mathematical Analysis, 31(3):535– 560, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33989.

Wang:2008:AAP

[Wan08] Xiaoqiang Wang. Asymptotic analysis of phase field formulations of bending elasticity models. SIAM Journal on Mathematical Analysis, 39(5): 1367–1401, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2009:ASS

[WC09] Junde Wu and Shangbin Cui. Asymptotic stability of stationary solutions of a free boundary problem modeling the growth of tumors with fluid tissues. SIAM Journal on Mathematical Analysis, 41(1): 391–414, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2007:EMD

[WCD07] Wei Wang, Daomin Cao, and Jinqiao Duan. Effective macroscopic dynamics of stochastic partial differential equations in perforated domains. SIAM Journal on Mathematical Analysis, 38(5):1508–1527, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Westdickenberg:2002:SNV

[Wes02] Michael Westdickenberg. Some new velocity averaging results. SIAM Journal on Mathematical Analysis, 33(5):1007–1032, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38076.

Wilkening:2009:PEE

[Wil09] Jon Wilkening. Practical error estimates for Reynolds’ lubrication approximation and its higher order corrections. SIAM Journal on Mathematical Analysis, 41(2):588–630, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2009:ESD

[WLW09] Zhi-Cheng Wang, Wan-Tong Li, and Jianhong Wu. Entire solutions in delayed lattice differential equations with monostable nonlinearity. SIAM Journal on Mathematical Analysis, 40(6):2392–2420, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2007:EIP

[WNW07] Jianhua Wu, Hua Nie, and Gail S. K. Wolkowicz. The effect of inhibitor on the plasmid-bearing and plasmidfree model in the unstirred chemostat. SIAM Journal on Mathematical Analysis, 38(6): 1860–1885, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wright:2005:CKA

[Wri05] J. Douglas Wright. Corrections to the KdV approximation for water waves. SIAM Journal on Mathematical Analysis, 37(4): 1161–1206, January 2005. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2005:GSD

[Wu05] Jiahong Wu. Global solutions of the 2D dissipative quasi-geostrophic equation in Besov spaces. SIAM Journal on Mathematical Analysis, 36(3):1014–1030, 2005.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43557.

Wei:2002:CTS

[WW02] Juncheng Wei and Matthias Winter. Critical threshold and stability of cluster solutions for large reaction-diffusion systems in R^{1}. SIAM Journal on Mathematical Analysis, 33(5):1058–1089, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38170.

Walker:2007:GEC

[WW07] Christoph Walker and Glenn F.

Webb. Global existence of classical solutions for a haptotaxis model. SIAM Journal on Mathematical Analysis, 38(5): 1694–1713, ???? 2007. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2005:ZVL

[WX05] Ya-Guang Wang and Zhouping Xin. Zero-viscosity limit of the linearized compressible Navier–Stokes equations with highly oscillatory forces in the half-plane. SIAM Journal on Mathematical Analysis, 37(4): 1256–1298, January 2005. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2006:QNL

[WXM06] Shu Wang, Zhouping Xin, and Peter A. Markowich. Quasineutral limit of the drift diffusion models for semiconductors: The case of general sign-changing doping profile. SIAM Journal on Mathematical Analysis, 37(6):1854– 1889, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2008:BDS

[WY08] Shuyin Wu and Zhaoyang Yin. Blow-up and decay of the solution of the weakly dissipative Degasperis–Procesi equation. SIAM Journal on Mathematical Analysis, 40(2):475– 490, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinstein:2005:RDS

[WZ05] Michael I. Weinstein and Vadim Zharnitsky. A reaction dispersion system and Raman interactions. SIAM Journal on Mathematical Analysis, 36(6):1742–1771, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42817.

Wei:2008:GBS

[WZF08] Mingjun Wei, Ting Zhang, and Daoyuan Fang. Global behavior of spherically symmetric Navier–Stokes equations with degenerate viscosity coefficients. SIAM Journal on Mathematical Analysis, 40(3): 869–904, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2003:ETD

[Xu03] Xiangsheng Xu. On the effects of thermal degeneracy in the thermistor problem. SIAM Journal on Mathematical Analysis, 35(4):1081–1098, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41740.

Xu:2004:NFR

[Xu04] Junxiang Xu. Normal form of reversible systems and persistence of lower dimensional tori under weaker nonresonance conditions. SIAM Journal on Mathematical Analysis, 36(1):233–255, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
42192.

Xu:2009:RTL

[Xu09] Jiang Xu. Relaxation-time limit in the isothermal hydrodynamic model for semiconductors. SIAM Journal on Mathematical Analysis, 40(5): 1979–1991, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2001:EGS

[YCL01] Yin Yang, Hua Chen, and Weian Liu. On existence of global solutions and blow-up to a system of reaction-diffusion equations modelling chemotaxis. SIAM Journal on Mathematical Analysis, 33(4):763– 785, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33779.

Yin:2004:SSP

[Yin04] Zhaoyang Yin. On the structure of solutions to the periodic Hunter–Saxton equation. SIAM Journal on Mathematical Analysis, 36(1):272– 283, 2004. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42567.

Yoshikawa:2007:QLT

[Yoc07] Shuji Yoshikawa, Irena Paw ow, and Wojciech M. Zaj czkowski. Quasi-linear thermoelasticity system arising in shape memory materials. SIAM Journal on Mathematical Analysis, 38(6):1733–1759, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yoon:2001:SAO

[Yoo01] Jungho Yoon. Spectral approximation orders of radial basis function interpolation on the Sobolev space.

SIAM Journal on Mathematical Analysis, 33(4):946–958, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37381.

Yuan:2006:TST

[Yua06] Hairong Yuan. On transonic shocks in two-dimensional variable-area ducts for steady Euler system. SIAM Journal on Mathematical Analysis, 38 (4):1343–1370, January 2006. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yi:2008:VIA

[YYW08] Fahuai Yi, Zhou Yang, and Xiaohua Wang. A variational inequality arising from European installment call options pricing. SIAM Journal on Mathematical Analysis, 40(1):306– 326, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yordanov:2005:FTB

[YZ05] Borislav Yordanov and Qi S.

Zhang. Finite-time blowup for wave equations with a potential. SIAM Journal on Mathematical Analysis, 36(5): 1426–1433, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44019.

Yarahmadian:2009:PGF

[YZ09] Shantia Yarahmadian and Kevin Zumbrun. Pointwise Green function bounds and long-time stability of largeamplitude noncharacteristic boundary layers. SIAM Journal on Mathematical Analysis, 40(6):2328–2350, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zagatti:2005:MPN

[Zag05] Sandro Zagatti. On the minimum problem for nonconvex scalar functionals. SIAM Journal on Mathematical Analysis, 37(3):982–995, 2005. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        61250.

Zheng:2000:ECO

[ZD00] Yibin Zheng and Peter C. Doerschuk. Explicit computation of orthonormal symmetrized harmonics with application to the identity representation of the icosahedral group. SIAM Journal on Mathematical Analysis, 32(3):538– 554, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34177.

Zhang:2003:NPW

[Zha03a] Kewei Zhang. Neighborhoods of parallel wells in two dimensions that separate gradient Young measures. SIAM Journal on Mathematical Analysis, 34(5):1207–1225, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39277.

Zhang:2003:WMS

[Zha03b] Ping Zhang. Wigner measure and the semiclassical limit of Schrödinger–Poisson equations. SIAM Journal on Mathematical Analysis, 34(3):700– 718, 2003. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39340.

Zhang:2008:COE

[Zha08] Xu Zhang. Carleman and observability estimates for stochastic wave equations. SIAM Journal on Mathematical Analysis, 40(2):851–868, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhou:2001:SSL

[Zho01] Ding-Xuan Zhou. Self-similar lattice tilings and subdivision schemes. SIAM Journal on Mathematical Analysis, 33(1):1–15, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36797.

Zingano:2002:HLT

[ZS02] Paulo R. Zingano and Stanly L. Steinberg. On the Hardy– Littlewood theorem for functions of bounded variation. SIAM Journal on Mathematical Analysis, 33(5):1199–1210, 2002. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36930.

Zuidwijk:2000:DTS

[Zui00] Rob A. Zuidwijk. Directional and time-scale wavelet analysis. SIAM Journal on Mathematical Analysis, 31(2):416– 430, 2000. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33335.

Zworski:2001:RIP

[Zwo01]  Maciej Zworski.  A re-

mark on isopolar potentials. SIAM Journal on Mathematical Analysis, 32(6):1324–1326, 2001. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37558.

Zhang:2000:SND

[ZY00] Zhijun Zhang and Jianning Yu. On a singular nonlinear Dirichlet problem with a convection term. SIAM Journal on Mathematical Analysis, 32(4):916– 927, 2000. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33216.

Zhang:2008:NMR

[ZZ08] Hui Zhang and Pingwen Zhang. On the new multiscale rodlike model of polymeric fluids. SIAM Journal on Mathematical Analysis, 40(3): 1246–1271, ???? 2008. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).